### v2

# Dysca: A Dynamic and Scalable Benchmark for Evaluating Perception Ability of LVLMs

Abstract
Currently many benchmarks have been proposed to evaluate the perception ability of the Large Vision-Language Models (LVLMs). However, most benchmarks conduct questions by selecting images from existing datasets, resulting in the potential data leakage. Besides, these benchmarks merely focus on evaluating LVLMs on the realistic style images and clean scenarios, leaving the multi-stylized images and noisy scenarios unexplored. In response to these challenges, we propose a dynamic and scalable benchmark named Dysca for evaluating LVLMs by leveraging synthesis images. Specifically, we leverage Stable Diffusion and design a rule-based method to dynamically generate novel images, questions and the corresponding answers. We consider 51 kinds of image styles and evaluate the perception capability in 20 subtasks. Moreover, we conduct evaluations under 4 scenarios (i.e., Clean, Corruption, Print Attacking and Adversarial Attacking) and 3 question types (i.e., Multi-choices, True-or-false and Free-form). Thanks to the generative paradigm, Dysca serves as a scalable benchmark for easily adding new subtasks and scenarios. A total of 8 advanced open-source LVLMs with 10 checkpoints are evaluated on Dysca, revealing the drawbacks of current LVLMs. The benchmark is released in https://github.com/Benchmark-Dysca/Dysca.
footnotetext: Equal contribution.
Corresponding Author: zhangjie@ict.ac.cn
1 Introduction
Recent years have witnessed the great success of the Large Vision-Language Models (LVLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. These models leverage the powerful Large Language Models (LLMs) [11, 12, 13, 14, 15] as their brain and incorporate the state-of-the-art visual encoders [16, 17, 18] as their eyes. Thanks to the alignment of visual feature with textual space and the development of visual instruction tuning techniques [4], LVLMs showcase the impressive capability in terms of visual scene comprehension and multimodal instruction-following.
In order to comprehensively evaluate the capabilities of LVLMs, many benchmarks have been purposed [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], where we categorize the current benchmarks into three types [25]. The first type is the classical benchmarks, such as COCO Caption [30] and VQA [19, 31, 32]. Although these benchmarks provide high-quality evaluation data, they also have notable limitations. On the one hand, they are inadequate for measuring the fine-grained capabilities of current LVLMs, offering the limited insightful feedback for the future improvement. On the other hand, since these classical benchmarks have been available as the open-source test data for a long time, it is hard to prevent the data leakage problem. The second type of benchmarks evaluate the LVLMs through a subjective manner [28, 33]. Although the benchmarks reveal the insightful drawbacks of current models, their data scale is limited (i.e., less than 200 annotations) and they require manual evaluation by experts. The third type is built for objectively evaluating current LVLMs and the comparison between them are shown in Tab. 1. They provide an objective and automatic evaluation manner, giving the fine-grained evaluation for the LVLMs. However, these benchmarks conduct Vision-language QAs by selecting images from existing dataset. Although they claim that the questions are re-annotated, the previous work [29] has demonstrated that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs. Besides, most benchmarks focus on evaluating LVLMs in the realistic images and clean scenarios, leaving the multi-stylized images and noisy scenarios unexplored. While some works like MMCBench [34] and Typographic Dataset [35] have investigated the robustness of LVLMs with corrupted and print-attacked images, respectively, they have not explored the effect of these noisy images on various perceptual tasks.
Figure 1: Overview of the automatic pipeline for generating Vision-language QAs, cleaning Vision-language QAs and evaluating LVLMs. (a) We first constructs prompts in terms of content, style and background, leveraging the Text-to-Image (T2I) diffusion model (e.g., SDXL [36]) to synthesis images to be asked. Then based on the scenarios and the question type, we post-process the synthesis images and generate the specific textual questions, respectively. (b) We further filter out low quality Vision-language QAs by utilizing trained models to form the final Dysca. (c) Finally, we evaluate LVLMs on our Dysca and feedback the fine-grained evaluation results.
In this paper, aiming to address these challenges above, we propose Dysca which is a dynamic and scalable benchmark for evaluating the perception ability of LVLMs via various subtasks and scenarios. Inspired by the prior evaluation works for LLMs [37], we investigate on whether we could leverage the large-scale synthesized images for evaluating LVLMs. We display the overview of our pipeline in Fig. 1. Specifically, we leverage Stable Diffusion and design a rule-based method to dynamically generate novel images, questions and corresponding answers. We decouple the prompt into 4 part, i.e., attribute, foreground, style and background, and design pre-defined templates to dynamically generate prompts, as displayed in Fig. 3. Then we utilize the state-of-the-art text-to-image diffusion models (e.g., SDXL [36]) to generate the corresponding images. Since we already know the main information of the images through prompts, we easily generate question-answer textual pairs by the rule-based method. After that, in order to obtain the high quality Vision-language QAs, we employ CLIP [16] to perform data cleaning on the generated Vision-language QA pairs. Dysca focuses on assessing the fine-grained perception abilities, including recognizing human, animal, object, landmark, etc. Dysca evaluates LVLMs with 20 perceptual subtasks, containing a total number of 51 different artistic styles. Besides, to evaluate the robustness of the models across different scenarios and question types, we construct 4 testing scenarios (clean, corruption, print attacking and adversarial attacking) and 3 question types (multi-choices, true-or-false and free-form questions). In the end, Dysca consists of 617K Vision-language QA pairs ($\times$20 larger than MM-BigBench [38] and $\times$25 larger than Seed-Bench2 [24] as shown in Tab. 1). Thanks to the generative paradigm, Dysca achieves the scalable benchmark to new subtasks and scenarios and dynamically generate unlimited Vision-language QAs for evaluation.
In summary, our work makes the following key contributions:
  • Dynamic and Scalable Benchmark: We propose Dysca, a benchmark that is able to dynamically generate the test data that users need and is easily to scale up to to new subtasks and scenarios.
  • Multi-grained Perceptual Subtasks and Multi-scenarios: Dysca evaluates the 20 perceptual subtasks performance of 8 mainstream LVLMs with 10 checkpoints under 4 image scenarios (i.e., clean, corruption, print attacking and adversarial attacking) and 3 question types (i.e., multi-choices, true-or-false and free-form questions).
  • Analysis and Observations: We demonstrate for the first time that evaluating LVLMs using large-scale synthetic data is valid. Experiments show the strong correlation coefficient between our evaluation rankings and the rankings obtained from non-synthetic benchmarks. The evaluation results also reveal the weakness of current LVLMs when facing different question types, image styles and image scenarios.
Table 1: Comparisons between existing LVLM benchmarks. ’✓ ’ indicates that the benchmarks include both newly collected images / annotations and images / annotations gathered from existing datasets. ’*’ The scale of our released benchmark is 617K, however Dysca is able to generate unlimited data to be tested.
Benchmark #Evaluation Data Scale #Perceptual Tasks
Automatic
Annotation
Collecting from Existing Datasets Question Type
Automatic
Evaluation
LLaVA-Bench [4] 0.15K - Free-form
MME [25] 2.3K 10 True-or-false
LVLM-eHub [21] - 3 Free-form
tiny-LVLM-eHub [22] 2.1K 3 Free-form
SEED-Bench [23] 19K 8 Multi-choices
MMBench [39] 2.9K 12 Multi-choices
TouchStone [26] 0.9K 10 Free-form
REFORM-EVAL [40] 50K 7 Multi-choices
MM-BigBench [38] 30K 6 Multi-choices
MM-VET [27] 0.2K 4 Free-form
MLLM-Bench [41] 0.42K 7 Free-form
SEED-Bench2 [24] 24K 10 Multi-choices
BenchLMM [42] 2.4K 15 Free-form
JourneyDB [43] 5.4K 2
Free-form
Multi-choices
[…]
Free-form
Multi-choices
True-or-false
2 Related Works
2.1 Large Vision-Language Models
The landscape of Large Vision-Language Models (LVLMs) has been significantly shaped by the pioneering success of Large Language Models (LLMs) such as GPTs [44, 45, 46] and LLaMA [13], catalyzing advancements in multimodal content understanding and generation [47], including intricate tasks like image-text comprehension. At the forefront of these developments, BLIP-2 [1] introduces a lightweight Q-Former [1] that facilitates alignment between textual and visual representations through a cross-attention mechanism [1]. InstructBLIP [3] takes a step further by incorporating textual instructions into the Q-Former, which significantly improves performance. LLAVA [4] employs GPT-4 [14] to transform data into multimodal instruction-following data and uses CLIP [16] and LLAMA [13] for fine-tuning instructions, achieving advanced performance. LLAVA-1.5 [48] extends this paradigm by integrating MLP projection and introducing academic task-specific Vision-language QA data. Recently, models like Otter [5], MiniGPT-4 [2], Qwen-VL-Chat [49] and XComposer-VL [7] further unleash the cross-modal understanding capabilities of LVLMs.
2.2 Benchmarks for LVLMs
The great progress of LVLMs triggers the development of benchmarks for evaluating these models, where we divide previous benchmarks into three categories. The first type is the classical benchmarks which focuses on evaluating LVLMs abilities via image caption [50] and VQA [51, 19]. However, these benchmarks cannot provide the fine-grained feedback on how to improve the models. Besides, since these benchmarks have been the public resources for a long time, it is hardly to guarantee that the LVLMs have not use them for training. The second type subjectively evaluates LVLMs by experts [28, 33]. Although these benchmarks reveal the insightful feedback of the LVLMs, their scale is limited (i.e., less than 200 annotations). The subjective manner also makes the evaluation expensive and hardly to expand the scale of the benchmarks.
The third type [4, 25, 21, 22, 23, 24, 39, 26, 40, 38, 27, 41, 42, 29, 52] focuses on evaluating LVLMs in an objective and large-scaled manner, where we list the detailed information of them in the Tab. 1. Some of them have been adopted by the community [53] as the standard benchmarks for evaluating LVLMs [12, 5, 7], like MME [25] and MMBench [39]. These benchmarks evaluate models through the objective answer types and most of them leverage the automatic annotation and evaluation manner for revealing the fine-grained drawbacks of current LVLMs. However, the previous benchmarks primarily concentrate on evaluating LVLMs using realistic images and clean scenario, leaving multi-stylized images and noisy scenarios unexplored. Moreover, many of them conduct QA by selecting images from publicly available datasets (e.g., [50, 54]). While they state that the questions have been re-annotated, they cannot guarantee that the LVLMs have not seen the image during training stage. The previous work [29] has proved that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs. One possible way to solve data leakage is using novel but synthesis images, where JourneyDB [43] is the first work aiming to leverage synthesis images to evaluate current LVLMs. The prompts and the corresponding images are downloaded from Midjourney [55] and ChatGPT [12] is leveraged to label the images. However, JourneyDB is a top-down framework where the number of images is fixed. Besides, the ChatGPT labeling may cause hallucinate annotations, leading to the unreliable evaluation results. Although 40 annotators have involved to clean the data, the data cleaning cost are expensive and it limits the data scale. In contrast, our Dysca serves as the bottom-up framework, allowing for dynamic and scalable generation for both images and evaluation questions. The rule-based question generation method also makes the annotations more accuracy. Besides, Dysca contains 20 subtasks which is more comprehensive than JourneyDB.
3 Dysca
3.1 Overview of Our Pipeline
The overview of our pipeline is shown in Fig. 1, containing data generation, data cleaning and LVLMs evaluation. For the data generation, our Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, where $M$ means "Metadata", $P$ means "Prompt", $I$ means "Image" and $Q$ means "Question-answer pair". We further decouple the metadata $M$ into 4 parts, i.e., "style", "attribute", "foreground" and "background", and the combination of the four parts constitute the image prompts $P$. Then, given the prompt $P$ and the selected scenario, we leverage the Text-to-Image (T2I) diffusion model (e.g., SDXL [36]) to synthesis image $I$ and add the specific perturbation to the image $I$. After that, since the prompt already includes the question angle and the corresponding answer, we construct a rule-based approach to generate the $Q$. Three types of questions are considered, i.e., multi-choice, true-or-false and free-form. Multi-choice and true-or-false questions utilize a closed-ended manner to assess LVLMs, while free-form questions employ an open-ended manner through image captioning for evaluation. For the data cleaning, considering that the T2I diffusion model may generate unsuccessful outcomes, we then use CLIP [16] and PP-OCRv3 [56] to automatically clean the whole dataset to obtain the final Dysca. Finally, we evaluate 8 open-source LVLMs with 10 checkpoints on our proposed Dysca.
Statistic #Number
Total questions 617K
- Clean 156K (25.2%)
- Print attacking 149K (24.1%)
- Adversarial attacking 156K (25.2%)
- Corruption 156K (25.2%)
Question type
- Multi-choices 251K (40.6%)
- True-or-false 250K (40.5%)
- Free-form 116K (18.8%)
Image resolution 1024*1024
Unique number of images 289K
Unique number of questions 162K
Unique number of answers 31K
Average question length 37.8
Average answer length 2.7
Average choice number 3.0
Table 2: Key statistics of Dysca.
Figure 2: Overview of the dataset distribution of 20 perceptual tasks. The number in each subtask shows the corresponding amount of their annotation.
3.2 Perceptual Tasks
Evaluation dimensions. Perception is one of the most fundamental capabilities of LVLMs and previous works [25] have shown that the lack of perceptual ability may result in hallucination [57]. In order to comprehensively evaluate LVLMs’ perception capability, we design 20 perceptual subtasks where we show all subtasks and the corresponding amount of their annotation in the Fig. 2. We investigate on two types of perception dimensions, i.e., coarse-grained and fine-grained perception. Coarse-grained perception involves recognizing the style, background and color of images. Fine-grained perception involves recognizing the animal, object, plant, food, age, gender, expression, race, celebrity, action, text, clothes, movie, anime, landmark, profession and TV shows.
Data sources. For each perceptual subtask, we collect the textual data first to construct the metadata $M$. For the TV shows, Anime and Movie, we select the titles from the rating list of IMDb***https://www.imdb.com/ based on the number of user reviews. For the styles, we utilize the style lists collected from the communityhttps://stable-diffusion-art.com/sdxl-styles/ and remove those which have strong reflect on the image content like "architectural style" and "Pokemon style". Note that the style list does not include the style prompt associated with a particular artist’s name. Besides, for the remaining contents, we select them from the label of current dataset (e.g., ImageNet [54]). All the selected textual data above constitute the metadata $M$. We provide the detailed information of the metadata in the Appendix C.
3.3 Construction of Questions & Answers
Recall that the data generation for Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, denoting the metadata ($M$), prompt ($P$), image ($I$) and question-answer pairs ($Q$), respectively. The relationships between these parts and the process of constructing Dysca are shown in Fig. 3. The metadata M is the core of the whole Dysca, containing all the information for generating $P$, $I$ and $Q$. The metadata $M$ consists of foreground, attribute, background and style, and these information guide the generation of the prompt ($P$) through pre-designed templates. Then, we utilize the T2I diffusion model to generate the corresponding image using the prompt $P$. For generating the image with specific text on it for the OCR subtask, we leverage TextDiffusion2 [58], which is the state-of-the-art text rendering method. For the rest of images, we leverage Stable Diffusion XL [36]. Subsequently, based on the different question types we select, i.e., multi-choices, true-or-false and free-form, we generate the corresponding Vision-language QA pairs in Dysca.
Besides, in order to evaluate the model performance under various scenarios, we conduct experiments on 4 scenarios, i.e., clean, corruption, print attacking and adversarial attacking. For the print attacking, followed by [35], we add the deceptive text on the image, where the text is a wrong option. Besides, to comprehensively evaluate the performance of LVLMs under corruption scenario, we add more typographic factors to original settings (i.e., different font orientations and font positions). For the adversarial attacking, we leverage PGD [59] to generate the adversarial image. We use InstructBLIP [3] as the proxy model and regard others as the black box models. The reason why we choose InstructBLIP is that it has shown superior performance in clean scenario. Besides, the black-box setting better reflects the robustness of the models when they face the real-world adversarial attacks. For the corruption, we leverage the image corruption methods collected from [34]. We remove some hard corruptions as they significantly impact the quality of the image, leading to human failure in judging the style and content of the image. The detailed examples are shown in Appendix D.
Consider that the Text-to-image diffusion model may generate the failure cases that affect the quality of the proposed benchmark, we leverage the off-the-shelf models, i.e., PP-OCRv3 [56] and CLIP-L-14 [16], to clean the data. PP-OCRv3 [56] is leveraged as the filter to exclude the failure image that TextDiffusion2 [58] generates the wrong text on the image. For the other images, we use CLIP-L-14 [16] to filter out the images with low text-image consistency. In the end, we filter out nearly 15% of low quality samples. The final statistics of our released Dysca are shown in Tab. 2. Note that the OCR subtask does not involve print attacking scenario as misidentifying adversarial text does not indicate poor OCR robustness of the LVLMs. Therefore, there are 7K fewer questions in the print attacking scenario. Besides, for the free-form question type, since it allows to assess the model’s perception abilities across multiple subtasks at the same time, we reduce the number of free-form questions for achieving a balanced data distribution.
Figure 3: The process of generating the prompt (P), image (I) and question-answer pairs (Q) from the metadata (M).
3.4 Evaluation Strategy
[…]
Table 3: Evaluation results on the 20 perceptual subtasks. The top two results on each subtask are bolded and underlined, respectively. "MC" and "TF" indicate the accuracy (%) of "Multi-choices" and "True-or-false", respectively.
Movie Action TV Show Profession Landmark
Model Language Model MC TF MC TF MC TF MC TF MC TF
Blip2 [1] Flan-T5-XL 71.46 68.61 96.24 93.04 56.29 61.99 77.88 75.87 98.70 94.58
InstructBLIP [3] Flan-T5-XL 77.35 67.38 97.28 93.61 65.89 60.96 79.17 76.51 98.04 95.75
XComposer-VL [7] InternLM-7B 81.90 78.36 97.03 94.75 77.81 73.29 83.97 75.87 98.04 94.81
LLava-1.5 [48] Vicuna-7B 55.95 51.68 74.78 50.75 56.29 50.68 58.65 56.83 85.00 54.48
LLava-1.5 [48] Vicuna-13B 66.29 54.04 77.93 59.71 59.27 54.11 66.67 59.05 94.35 55.90
MiniGPT-4 [2] Vicuna-7B 32.68 52.35 42.96 51.00 30.46 50.00 33.97 50.16 53.26 52.36
Otter [5] LLaMA-7B 65.36 59.08 65.85 72.36 68.54 57.53 66.99 54.92 58.48 66.51
Qwen-VL-Chat [49] Qwen-7B 71.25 51.01 96.35 63.14 67.22 49.32 75.64 54.60 96.96 54.72
Shikra [6] Vicuna-7B 66.29 59.64 78.00 77.29 60.26 56.85 78.85 68.25 89.35 70.99
Shikra-VQA [6] Vicuna-7B 66.39 61.66 96.17 80.32 59.93 60.27 78.21 67.94 92.83 72.88
Anime Clothes Celebrity Food Plant
Model Language Model MC TF MC TF MC TF MC TF MC TF
Blip2 [1] Flan-T5-XL 57.07 61.94 82.38 73.79 81.98 76.48 91.52 88.04 92.50 91.19
InstructBLIP [3] Flan-T5-XL 61.28 64.04 86.51 81.08 83.69 76.14 91.71 88.85 93.31 91.92
XComposer-VL [7] InternLM-7B 75.41 74.58 86.18 85.73 88.53 87.28 92.33 90.21 93.25 93.14
LLava-1.5 [48] Vicuna-7B 47.15 48.46 47.19 50.41 57.36 54.72 54.63 51.10 49.08 48.59
LLava-1.5 [48] Vicuna-13B 57.61 50.14 65.20 57.20 60.14 57.94 78.87 57.45 79.72 55.51
MiniGPT-4 [2] Vicuna-7B 29.21 48.74 31.25 50.47 28.62 49.64 44.66 50.78 45.38 50.53
Otter [5] LLaMA-7B 61.82 57.30 47.13 69.52 42.41 63.52 46.81 79.62 66.41 81.51
Qwen-VL-Chat [49] Qwen-7B 71.60 54.63 78.44 60.72 87.37 50.05 89.37 53.37 92.32 59.91
Shikra [6] Vicuna-7B 47.96 57.72 75.21 59.65 63.76 59.86 83.30 68.37 88.55 66.60
Shikra-VQA [6] Vicuna-7B 49.05 57.58 76.57 63.10 63.73 62.13 89.25 71.22 88.34 71.76
Age Gender Expression Race Animal
Model Language Model MC TF MC TF MC TF MC TF MC TF
Blip2 [1] Flan-T5-XL 62.61 59.65 99.37 94.86 89.27 75.92 74.38 71.95 96.64 95.27
InstructBLIP [3] Flan-T5-XL 65.14 59.55 99.51 88.21 91.85 79.34 77.99 74.62 97.21 94.98
XComposer-VL [7] InternLM-7B 67.89 78.35 99.60 98.06 89.64 82.52 80.42 76.69 97.83 96.70
LLava-1.5 [48] Vicuna-7B 38.35 55.15 54.14 49.78 63.86 49.76 43.98 50.56 49.26 50.90
LLava-1.5 [48] Vicuna-13B 49.55 59.59 98.71 83.53 71.29 58.33 70.16 62.84 85.99 58.23
MiniGPT-4 [2] Vicuna-7B 31.75 51.71 56.49 49.22 42.01 50.91 27.54 50.41 45.97 49.57
Otter [5] LLaMA-7B 37.98 51.19 78.23 77.96 73.86 60.36 43.00 57.88 81.32 83.63
Qwen-VL-Chat [49] Qwen-7B 53.48 49.09 97.72 58.55 85.22 64.30 73.50 56.88 95.60 64.37
Shikra [6] Vicuna-7B 65.37 56.71 97.85 73.16 90.47 70.55 74.17 54.88 89.82 69.06
Shikra-VQA [6] Vicuna-7B 65.50 57.34 99.39 81.41 92.36 74.71 73.90 55.24 91.18 74.26
Object OCR Style Background Color
Model Language Model MC TF MC TF MC TF MC TF MC TF
Blip2 [1] Flan-T5-XL 89.32 87.75 71.89 62.07 99.86 88.62 63.97 68.37 88.01 85.55
InstructBLIP [3] Flan-T5-XL 89.87 90.13 72.61 60.93 99.96 81.23 66.50 69.19 90.93 85.55
XComposer-VL [7] InternLM-7B 90.23 91.72 69.22 78.66 95.83 81.32 71.26 75.27 87.97 89.02
LLava-1.5 [48] Vicuna-7B 58.63 49.99 47.08 51.02 46.35 50.49 44.33 50.95 41.91 51.90
LLava-1.5 [48] Vicuna-13B 77.86 57.93 58.19 53.17 85.80 51.56 62.67 55.33 63.81 52.90
MiniGPT-4 [2] Vicuna-7B 51.75 51.55 31.37 51.56 43.41 48.64 31.94 50.10 35.29 50.51
Otter [5] LLaMA-7B 47.31 82.51 66.23 61.41 98.04 61.58 63.06 63.40 48.28 57.97
Qwen-VL-Chat [49] Qwen-7B 88.32 63.14 71.74 51.86 98.21 52.43 69.42 51.59 86.39 53.05
Shikra [6] Vicuna-7B 70.91 69.22 60.79 53.63 95.13 58.95 71.24 62.66 84.91 61.83
Shikra-VQA [6] Vicuna-7B 89.43 76.00 59.63 54.14 99.51 61.48 70.98 65.96 83.79 64.49
[…]
4 Results and Analysis
In this section, we report the evaluation results and make insightful analysis. A total of 8 LVLMs with 10 checkpoints are evaluated on Dysca benchmark, including BLIP2 [1], InstructBLIP [3], LLavA [48], MiniGPT-4 [2], Otter [5], XComposer-VL [7], Qwen-VL-chat [49] and Shikra [6]. Each model is evaluated with all the 20 perception subtasks under 4 scenarios. The detailed rankings for each subtask can be found in the Appendix A.
4.1 Main Results
Clean Scenario. The evaluation results of various LVLMs in different perceptual subtasks under clean scenarios are presented in Tab. 3. Since the evaluation for free-form question type usually involves multiple subtasks, we can not calculate the results of free-form for each subtask individually. Instead, we display the overall score of free-form in the first row in Tab. 4. As can be seen, Xcomposer-VL [7] outperforms other LVLMs, achieving top-1 or top-2 results in most subtasks, but InstructBLIP [3], Qwen-VL-chat [49] and BLIP2 also take a lead in a few subtasks.
Noisy Scenarios The evaluation results of various LVLMs under noisy scenarios (i.e., corruption, print attacking and adversarial attacking) are presented in Tab. 4. As can be seen, For the multi-choices and true-or-false question type, Xcomposer-VL [7] still takes a lead on all 4 scenarios. For the free-form, LLava-1.5-7b [48] achieves the best.
Table 4: Evaluation results on 4 scenarios. "MC", "TF" and "FF" indicates "Multi-choices", "True-or-false" and "Free-form", respectively. "PrintAtt" and "AdverAtt" means "Print Attacking" and "Adversarial Attacking", respectively. "*": the model is under white-box setting.
Scenarios Blip2 [1] InstructBLIP [3] XComposer-VL [7] LLava-1.5-13b [48] LLava-1.5-7b [48]
MC TF FF MC TF FF MC TF FF MC TF FF MC TF FF
Clean 82.07 78.78 37.56 84.29 79.00 38.94 86.22 84.82 46.03 71.50 57.72 48.61 53.70 51.41 50.15
Corruption 81.17 77.98 37.32 83.77 78.57 38.72 85.20 84.16 45.52 70.34 57.57 48.28 53.21 51.40 49.54
PrintAtt 64.78 68.01 35.82 59.18 58.80 35.71 74.08 72.71 44.46 48.91 54.72 48.47 40.02 50.94 48.38
AdverAtt 23.28 48.98 25.70 21.49* 48.74* 29.15* 20.85 49.54 21.17 66.86 56.56 47.11 50.70 51.36 48.17
Scenarios MiniGPT-4 [2] Otter [5] Qwen-VL-Chat [49] Shikra [6] Shikra-VQA [6]
MC TF FF MC TF FF MC TF FF MC TF FF MC TF FF
Clean 38.50 50.51 38.90 61.36 65.99 39.12 82.31 55.84 49.90 76.61 63.79 47.17 79.31 66.69 29.93
Corruption 38.70 49.99 40.35 62.19 65.36 38.71 79.26 52.74 49.74 76.42 64.31 46.43 79.06 66.95 29.40
PrintAtt 36.16 50.00 42.45 38.11 36.27 36.93 59.68 46.40 48.65 60.02 40.18 47.03 62.20 41.05 27.91
AdverAtt 26.75 49.52 22.59 57.78 62.27 35.76 78.47 58.04 44.67 72.20 62.16 43.56 74.04 64.64 26.81
4.2 Analysis
4.2.1 Key Observations
(1) For individual models, their perceptual performance varies across different subtasks. For example, in the case of Qwen-VL-Chat [49], it achieves a score of 96.96% accuracy in landmark recognition task for multi-choices questions (2% below the first-place score), but obtains a score of 53.48% accuracy in age recognition task (12% below the first-place score). The results suggest that Qwen-VL-Chat [49] may require more fine-tuning in age perception data. Analyzing the performance of the models across various subtasks contribute to purposive improving.
Figure 4: Models exhibit different performance when facing the same image but different question types.
(2) Models exhibit performance inconsistency when facing multi-choices and true-or-false question types. As can be seen, in the object recognition subtask, Otter [5] achieves an accuracy of 47.31% in the multi-choices question type (22% higher than random guessing), but obtains an accuracy of 82.51% in the true-or-false question type (32% higher than random guessing). Interestingly, we observe the opposite results in the Qwen-VL-Chat [49]. In the object recognition subtask, it achieves an accuracy of 88.32% in the multi-choices but achieves an accuracy of 63.14% in the true-of-false. We also observe the same problem shown in other models and we display two examples in Fig. 4. We speculate that the inconsistency may be attributed to the bias in the training dataset towards particular question types, such as using more multi-choices or true-or-false questions.
(3) Each model exhibits robustness in the corruption scenario, but suffers degradation in the two attacking scenarios. As shown, all models exhibit minor score variation of less than 1% under the corruption scenario. However, they exhibit degradation when facing print attacking (e.g., 84.29% vs. 59.18% for InstructBlip [3] in multi-choices accuracy). XComposer-VL [7] shows the strongest robustness, maintaining over 70% accuracy for both multi-choices and true-or-false. Besides, since our adversarial algorithm specifically targets the image encoder, the LVLMs that share the same encoder architecture (i.e., Blip2, InstructBLip and XComposer-VL all using EVA-CLIP [62] as the image encoder) exhibit significant performance degradation, with accuracy even falling below random selection. Models utilizing alternative image encoders also experience a performance decrease of approximately 5% to 10%. A more detailed result can be found in Appendix D.
4.2.2 The Validity of Dysca
Table 5: The correlation results on three benchmarks, where $\rho\in[-1,1]$ and $\tau\in[-1,1]$.
Style Method MMBench OCRBench SeedBench-2
All $\rho$ 0.70 0.90 0.46
$\tau$ 0.60 0.80 0.43
Realistic $\rho$ 0.70 1.00 0.64
$\tau$ 0.60 1.00 0.62
In this section, we investigate on the evaluation gap between Dysca and non-synthesis benchmarks. We calculate the Spearman’s rank correlation coefficient [63] $\rho$ and the Kendall rank correlation coefficient [64] $\tau$ between the evaluation ranking of Dysca under clean scenario with the non-synthesis benchmark’s evaluation ranking, i.e., MMBench [39], OCRBench [52] and SeedBench-2 [24]. Both coefficient generate a score in the range of [-1,1], where 1 represents a perfect positive correlation, -1 represents a perfect negative correlation, and 0 represents no correlation. Specifically, we intersect our Dysca with current benchmarks based on the perceptual subtasks, evaluation models and evaluation question types. We then calculate the correlation of model evaluation rankings within this intersection. The results are shown in the first row of Tab. 5. For the MMbench [39] and OCRBench [52], both metrics show the high correlation, with $\rho$ and $\tau$ higher than 0.6. However, the correlation for SeedBench-2 [24] is not as strong. Considering that SeedBench-2 only contains realistic images, we conduct additional experiments using the evaluation ranks on our realistic style images only. As shown in the second row of Tab. 5, the correlation results of SeedBench-2 significantly improve (i.e., 0.46 vs. 0.64 for $\rho$ and 0.43 vs. 0.62 for $\tau$). The correlation with OCRBench also improves to 1, demonstrating the validity of using synthetic datasets for evaluation LVLMs.
To further explore the the impact of image styles on evaluation results, we present the average scores across all subtasks for each of the 51 styles in Fig. 5. We observe slight score differences across all styles. In the case of realistic styles such as "iPhone photo", all LVLMs perform better compared to other image styles. The LVLMs also exhibit better performance on unrealistic but common styles like "expressionist". However, for unrealistic and less common styles such as "gothic", all models show relatively poor performance. The results reveal that the gap between Dysca and non-synthesis benchmarks primarily stems from the more diverse range of image styles, making Dysca a more comprehensive benchmark for assessing the perception ability compared to previous benchmarks.
Figure 5: Illustration of each model’s performance across 51 image styles, where the darker colors represent higher scores. The representative styles are colored with non-black font. Realistic styles are shown in red font, unrealistic but common styles are displayed in yellow font, and unrealistic and less common styles are represented in blue font.
5 Conclusion
In this paper, we purpose Dysca, a dynamic and scalable benchmark for evaluating perception ability of Large Vision Language Models (LVLMs). Dysca consists of 617K Vision-language QA pairs, covering 20 perceptual subtasks, 4 image scenarios and 3 question types. We conduct the experiment on 8 advanced open-source LVLMs with 10 checkpoints, revealing the insightful weakness of current LVLMs when facing different question types, image styles and image conditions. Experiments demonstrate the validity on evaluating LVLMs by using synthesis images.
Acknowledgement
This work is partially supported by National Key R&\&&D Program of China (No. 2021YFC3310100), Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680000), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No. SYG202325), and Youth Innovation Promotion Association CAS.
Appendix
Appendix A The Leaderboards
Rank Model Score
  1 XComposer-VL 80.13
2 InstructBlip 72.36
3 Blip2 70.03
4 Shikra-VQA 64.03
5 Shikra 62.97
6 Otter 62.22
7 Qwen-VL-Chat 61.13
8 LLAVA-1.5-13B 60.17
9 LLAVA-1.5-7B 53.81
10 MiniGPT-4 42.52
(a) Clean-Movie
Rank Model Score
  1 XComposer-VL 95.89
2 InstructBlip 95.44
3 Blip2 94.64
4 Shikra-VQA 88.25
5 Qwen-VL-Chat 79.75
6 Shikra 77.65
7 Otter 69.10
8 LLAVA-1.5-13B 68.82
9 LLAVA-1.5-7B 62.77
10 MiniGPT-4 46.98
(b) Clean-Action
Rank Model Score
  1 XComposer-VL 75.55
2 InstructBlip 63.42
3 Otter 63.04
4 Shikra-VQA 60.10
5 Blip2 59.14
6 Shikra 58.55
7 Qwen-VL-Chat 58.27
8 LLAVA-1.5-13B 56.69
9 LLAVA-1.5-7B 53.48
10 MiniGPT-4 40.23
(c) Clean-TV Show
Rank Model Score
  1 XComposer-VL 79.92
2 InstructBlip 77.84
3 Blip2 76.88
4 Shikra 73.55
5 Shikra-VQA 73.07
6 Qwen-VL-Chat 65.12
7 LLAVA-1.5-13B 62.86
8 Otter 60.95
9 LLAVA-1.5-7B 57.74
10 MiniGPT-4 42.06
(d) Clean-Profession
Rank Model Score
  1 InstructBlip 96.90
2 Blip2 96.64
3 XComposer-VL 96.43
4 Shikra-VQA 82.85
5 Shikra 80.17
6 Qwen-VL-Chat 75.84
7 LLAVA-1.5-13B 75.12
8 LLAVA-1.5-7B 69.74
9 Otter 62.50
10 MiniGPT-4 52.81
(e) Clean-Landmark
Rank Model Score
  1 XComposer-VL 75.00
2 Qwen-VL-Chat 63.11
3 InstructBlip 62.66
4 Otter 59.56
5 Blip2 59.50
6 LLAVA-1.5-13B 53.88
7 Shikra-VQA 53.31
8 Shikra 52.84
9 LLAVA-1.5-7B 47.80
10 MiniGPT-4 38.98
(f) Clean-Anime
Rank Model Score
  1 XComposer-VL 85.96
2 InstructBlip 83.80
3 Blip2 78.09
4 Shikra-VQA 69.83
5 Qwen-VL-Chat 69.58
6 Shikra 67.43
7 LLAVA-1.5-13B 61.20
8 Otter 58.33
9 LLAVA-1.5-7B 48.80
10 MiniGPT-4 40.86
(g) Clean-Clothes
Rank Model Score
  1 XComposer-VL 87.91
2 InstructBlip 79.91
3 Blip2 79.23
4 Qwen-VL-Chat 68.71
5 Shikra-VQA 62.93
6 Shikra 61.81
7 LLAVA-1.5-13B 59.04
8 LLAVA-1.5-7B 56.04
9 Otter 52.97
10 MiniGPT-4 39.13
(h) Clean-Celebrity
Rank Model Score
  1 XComposer-VL 91.27
2 InstructBlip 90.28
3 Blip2 89.78
4 Shikra-VQA 80.23
5 Shikra 75.84
6 Qwen-VL-Chat 71.37
7 LLAVA-1.5-13B 68.16
8 Otter 63.22
9 LLAVA-1.5-7B 52.87
10 MiniGPT-4 47.72
(i) Clean-Food
Rank Model Score
  1 XComposer-VL 93.19
2 InstructBlip 92.62
3 Blip2 91.84
4 Shikra-VQA 80.05
5 Shikra 77.57
6 Qwen-VL-Chat 76.11
7 Otter 73.96
8 LLAVA-1.5-13B 67.61
9 LLAVA-1.5-7B 48.84
10 MiniGPT-4 47.95
(j) Clean-Plant
Rank Model Score
  1 XComposer-VL 73.12
2 InstructBlip 62.34
3 Shikra-VQA 61.42
4 Blip2 61.13
5 Shikra 61.04
6 LLAVA-1.5-13B 54.57
7 Qwen-VL-Chat 51.28
8 LLAVA-1.5-7B 46.75
9 Otter 44.58
10 MiniGPT-4 41.73
(k) Clean-Age
Rank Model Score
  1 XComposer-VL 98.83
2 Blip2 97.12
3 InstructBlip 93.86
4 LLAVA-1.5-13B 91.12
5 Shikra-VQA 90.40
6 Shikra 85.50
7 Qwen-VL-Chat 78.13
8 Otter 78.09
9 MiniGPT-4 52.86
10 LLAVA-1.5-7B 51.96
(l) Clean-Gender
Rank Model Score
  1 XComposer-VL 86.08
2 InstructBlip 85.59
3 Shikra-VQA 83.53
4 Blip2 82.59
5 Shikra 80.51
6 Qwen-VL-Chat 74.76
7 Otter 67.11
8 LLAVA-1.5-13B 64.81
9 LLAVA-1.5-7B 56.81
10 MiniGPT-4 46.46
(m) Clean-Expression
Rank Model Score
  1 XComposer-VL 78.56
2 InstructBlip 76.31
3 Blip2 73.16
4 LLAVA-1.5-13B 66.50
5 Qwen-VL-Chat 65.19
6 Shikra-VQA 64.56
7 Shikra 64.53
8 Otter 50.44
9 LLAVA-1.5-7B 47.27
10 MiniGPT-4 38.98
(n) Clean-Race
Rank Model Score
  1 XComposer-VL 97.27
2 InstructBlip 96.09
3 Blip2 95.95
4 Shikra-VQA 82.72
5 Otter 82.47
6 Qwen-VL-Chat 79.98
7 Shikra 79.44
8 LLAVA-1.5-13B 72.11
9 LLAVA-1.5-7B 50.08
10 MiniGPT-4 47.77
(o) Clean-Animal
Rank Model Score
  1 XComposer-VL 90.97
2 InstructBlip 90.00
3 Blip2 88.53
4 Shikra-VQA 82.72
5 Qwen-VL-Chat 75.73
6 Shikra 70.06
7 LLAVA-1.5-13B 67.89
8 Otter 64.91
9 LLAVA-1.5-7B 54.31
10 MiniGPT-4 51.65
(p) Clean-Object
Rank Model Score
  1 Blip2 68.38
2 InstructBlip 67.50
3 XComposer-VL 66.06
4 Shikra-VQA 64.86
5 Otter 64.37
6 Qwen-VL-Chat 63.28
7 Shikra 63.02
8 LLAVA-1.5-13B 58.84
9 LLAVA-1.5-7B 44.96
10 MiniGPT-4 43.70
(q) Clean-OCR
Rank Model Score
  1 XComposer-VL 73.94
2 Blip2 66.98
3 InstructBlip 66.77
4 Otter 63.82
5 Qwen-VL-Chat 61.80
6 Shikra 57.21
7 Shikra-VQA 56.89
8 LLAVA-1.5-13B 55.68
9 LLAVA-1.5-7B 49.05
10 MiniGPT-4 41.47
(r) Clean-Style
Rank Model Score
  1 XComposer-VL 73.27
2 Shikra-VQA 68.47
3 InstructBlip 67.84
4 Shikra 66.95
5 Blip2 66.17
6 Otter 63.23
7 Qwen-VL-Chat 60.51
8 LLAVA-1.5-13B 59.00
9 LLAVA-1.5-7B 47.64
10 MiniGPT-4 41.02
(s) Clean-Background
Rank Model Score
  1 XComposer-VL 88.50
2 InstructBlip 88.24
3 Blip2 86.78
4 Shikra-VQA 74.14
5 Shikra 73.37
6 Qwen-VL-Chat 69.72
7 LLAVA-1.5-13B 58.36
8 Otter 53.12
9 LLAVA-1.5-7B 46.91
10 MiniGPT-4 42.90
(t) Clean-Color
Table 6: Leaderboards of all Perceptual tasks under clean scenario.
Rank Model Score
  1 XComposer-VL 78.20
2 InstructBlip 71.30
3 Blip2 68.69
4 Shikra-VQA 64.19
5 Shikra 63.47
6 Otter 61.39
7 Qwen-VL 59.27
8 LLAVA-1.5-13B 58.67
9 LLAVA-1.5-7B 52.36
10 MiniGPT-4 43.02
(a) Corru.-Movie
Rank Model Score
  1 XComposer-VL 95.86
2 InstructBlip 95.09
3 Blip2 93.87
4 Shikra-VQA 88.72
5 Shikra 77.77
6 Qwen-VL 75.09
7 Otter 70.09
8 LLAVA-1.5-13B 68.34
9 LLAVA-1.5-7B 63.00
10 MiniGPT-4 47.22
(b) Corru.-Action
Rank Model Score
  1 XComposer-VL 72.86
2 InstructBlip 62.42
3 Otter 60.71
4 Shikra 60.09
5 Shikra-VQA 59.58
6 Blip2 58.16
7 Qwen-VL 57.95
8 LLAVA-1.5-13B 55.18
9 LLAVA-1.5-7B 52.34
10 MiniGPT-4 38.70
(c) Corru.-TV Show
Rank Model Score
  1 XComposer-VL 78.48
2 InstructBlip 77.84
3 Blip2 75.77
4 Shikra 73.72
5 Shikra-VQA 73.38
6 Qwen-VL 64.97
7 Otter 62.71
8 LLAVA-1.5-13B 62.23
9 LLAVA-1.5-7B 56.95
10 MiniGPT-4 43.22
(d) Corru.-Profession
Rank Model Score
  1 InstructBlip 96.33
2 Blip2 96.28
3 XComposer-VL 95.97
4 Shikra-VQA 83.01
5 Shikra 82.38
6 LLAVA-1.5-13B 74.63
7 Qwen-VL 73.55
8 LLAVA-1.5-7B 68.78
9 Otter 61.65
10 MiniGPT-4 49.55
(e) Corru.-Landmark
Rank Model Score
  1 XComposer-VL 72.28
2 InstructBlip 62.79
3 Qwen-VL 60.44
4 Blip2 58.67
5 Otter 57.24
6 LLAVA-1.5-13B 53.67
7 Shikra-VQA 52.28
8 Shikra 52.00
9 LLAVA-1.5-7B 47.87
10 MiniGPT-4 40.91
(f) Corru.-Anime
Rank Model Score
  1 XComposer-VL 84.87
2 InstructBlip 82.84
3 Blip2 76.69
4 Shikra-VQA 70.59
5 Shikra 67.05
6 Qwen-VL 64.26
7 LLAVA-1.5-13B 60.88
8 Otter 59.87
9 LLAVA-1.5-7B 48.12
10 MiniGPT-4 41.38
(g) Corru.-Clothes
Rank Model Score
  1 XComposer-VL 87.25
2 InstructBlip 79.23
3 Blip2 78.36
4 Qwen-VL 64.89
5 Shikra-VQA 62.55
6 Shikra 61.68
7 LLAVA-1.5-13B 58.64
8 LLAVA-1.5-7B 56.36
9 Otter 53.59
10 MiniGPT-4 38.80
(h) Corru.-Celebrity
Rank Model Score
  1 XComposer-VL 90.61
2 InstructBlip 90.08
3 Blip2 89.19
4 Shikra-VQA 80.55
5 Shikra 75.63
6 Qwen-VL 70.71
7 LLAVA-1.5-13B 67.46
8 Otter 65.06
9 LLAVA-1.5-7B 52.94
10 MiniGPT-4 47.76
(i) Corru.-Food
Rank Model Score
  1 XComposer-VL 92.45
2 InstructBlip 92.23
3 Blip2 91.15
4 Shikra-VQA 80.80
5 Shikra 77.69
6 Otter 73.42
7 Qwen-VL 72.01
8 LLAVA-1.5-13B 67.22
9 LLAVA-1.5-7B 48.84
10 MiniGPT-4 48.04
(j) Corru.-Plant
Rank Model Score
  1 XComposer-VL 73.04
2 InstructBlip 61.59
3 Shikra-VQA 61.09
4 Shikra 60.40
5 Blip2 60.03
6 Qwen-VL 55.26
7 LLAVA-1.5-13B 53.67
8 LLAVA-1.5-7B 46.76
9 Otter 45.09
10 MiniGPT-4 40.84
(k) Corru.-Age
Rank Model Score
  1 XComposer-VL 98.53
2 Blip2 97.76
3 InstructBlip 93.75
4 LLAVA-1.5-13B 91.09
5 Shikra-VQA 90.17
6 Shikra 85.91
7 Otter 80.22
8 Qwen-VL 67.00
9 MiniGPT-4 52.75
10 LLAVA-1.5-7B 52.27
(l) Corru.-Gender
Rank Model Score
  1 XComposer-VL 86.24
2 InstructBlip 84.95
3 Shikra-VQA 83.56
4 Blip2 81.56
5 Shikra 80.56
6 Qwen-VL 70.55
7 Otter 66.64
8 LLAVA-1.5-13B 64.94
9 LLAVA-1.5-7B 56.16
10 MiniGPT-4 46.55
(m) Corru.-Expression
Rank Model Score
  1 XComposer-VL 78.80
2 InstructBlip 76.19
3 Blip2 71.72
4 LLAVA-1.5-13B 65.33
5 Shikra-VQA 64.73
6 Shikra 63.94
7 Qwen-VL 63.55
8 Otter 49.73
9 LLAVA-1.5-7B 47.92
10 MiniGPT-4 39.65
(n) Corru.-Race
Rank Model Score
  1 XComposer-VL 97.06
2 InstructBlip 95.86
3 Blip2 95.69
4 Shikra-VQA 83.03
5 Otter 82.84
6 Shikra 79.44
7 LLAVA-1.5-13B 71.47
8 Qwen-VL 71.43
9 LLAVA-1.5-7B 50.67
10 MiniGPT-4 48.66
(o) Corru.-Animal
Rank Model Score
  1 XComposer-VL 90.81
2 InstructBlip 89.76
3 Blip2 88.05
4 Shikra-VQA 82.76
5 Shikra 71.49
6 Qwen-VL 67.78
7 LLAVA-1.5-13B 67.23
8 Otter 66.18
9 LLAVA-1.5-7B 54.09
10 MiniGPT-4 51.02
(p) Corru.-Object
Rank Model Score
  1 Blip2 94.06
2 InstructBlip 90.54
3 XComposer-VL 87.19
4 Shikra-VQA 80.80
5 Otter 78.70
6 Shikra 77.17
7 Qwen-VL 69.66
8 LLAVA-1.5-13B 68.58
9 LLAVA-1.5-7B 49.10
10 MiniGPT-4 44.63
(q) Corru.-OCR
Rank Model Score
  1 XComposer-VL 72.34
2 InstructBlip 65.06
3 Blip2 64.86
4 Otter 61.14
5 Qwen-VL 60.34
6 Shikra 56.61
7 Shikra-VQA 55.81
8 LLAVA-1.5-13B 53.65
9 LLAVA-1.5-7B 47.67
10 MiniGPT-4 40.81
(r) Courruption-Style
Rank Model Score
  1 XComposer-VL 72.53
2 Shikra-VQA 68.25
3 InstructBlip 67.48
4 Shikra 67.19
5 Blip2 64.69
6 Otter 62.99
7 Qwen-VL 60.50
8 LLAVA-1.5-13B 58.50
9 LLAVA-1.5-7B 47.12
10 MiniGPT-4 40.53
(s) Corru.-Background
Rank Model Score
  1 XComposer-VL 88.24
2 InstructBlip 88.09
3 Blip2 86.25
4 Shikra-VQA 74.14
5 Shikra 73.07
6 Qwen-VL 70.78
7 LLAVA-1.5-13B 57.70
8 Otter 56.14
9 LLAVA-1.5-7B 46.76
10 MiniGPT-4 42.86
(t) Corru.-Color
Table 7: Leaderboards of all Perceptual tasks under corruption scenario.
Rank Model Score
  1 XComposer-VL 67.09
2 Blip2 52.55
3 LLAVA-1.5-13B 46.74
4 LLAVA-1.5-7B 45.24
5 Shikra 44.84
6 Shikra-VQA 44.36
7 Qwen-VL 44.30
8 InstructBlip 43.06
9 MiniGPT-4 40.26
10 Otter 37.97
(a) Pr.Att.-Movie
Rank Model Score
  1 XComposer-VL 86.24
2 Blip2 79.13
3 InstructBlip 75.94
4 Shikra-VQA 71.36
5 Qwen-VL 65.26
6 Shikra 58.82
7 LLAVA-1.5-13B 57.66
8 LLAVA-1.5-7B 51.62
9 MiniGPT-4 47.07
10 Otter 39.77
(b) Pr.Att.-Action
Rank Model Score
  1 XComposer-VL 55.08
2 LLAVA-1.5-7B 43.55
3 Qwen-VL 40.72
4 LLAVA-1.5-13B 40.09
5 Blip2 39.71
6 Shikra 39.01
7 MiniGPT-4 37.40
8 Shikra-VQA 36.50
9 InstructBlip 32.17
10 Otter 31.48
(c) Pr.Att.-TV Show
Rank Model Score
  1 XComposer-VL 67.14
2 Blip2 59.62
3 Shikra 55.70
4 InstructBlip 54.52
5 Shikra-VQA 54.25
6 Qwen-VL 48.66
7 LLAVA-1.5-13B 47.49
8 LLAVA-1.5-7B 42.86
9 MiniGPT-4 42.22
10 Otter 38.12
(d) Pr.Att.-Profession
Rank Model Score
  1 XComposer-VL 89.81
2 Blip2 84.85
3 InstructBlip 76.78
4 Shikra-VQA 68.28
5 LLAVA-1.5-13B 65.98
6 Shikra 64.47
7 Qwen-VL 64.29
8 LLAVA-1.5-7B 56.39
9 MiniGPT-4 49.58
10 Otter 43.89
(e) Pr.Att.-Landmark
Rank Model Score
  1 XComposer-VL 59.72
2 Blip2 45.05
3 Qwen-VL 44.95
4 LLAVA-1.5-7B 42.23
5 LLAVA-1.5-13B 40.95
6 InstructBlip 38.26
7 Shikra-VQA 38.25
8 MiniGPT-4 37.27
9 Shikra 37.02
10 Otter 36.16
(f) Pr.Att.-Anime
Rank Model Score
  1 XComposer-VL 76.43
2 Blip2 66.53
3 InstructBlip 62.14
4 Qwen-VL 51.56
5 LLAVA-1.5-13B 49.63
6 Shikra-VQA 49.02
7 Shikra 47.23
8 LLAVA-1.5-7B 42.22
9 MiniGPT-4 39.66
10 Otter 34.67
(g) Pr.Att.-Clothes
Rank Model Score
  1 XComposer-VL 73.48
2 Blip2 57.30
3 Qwen-VL 52.34
4 LLAVA-1.5-7B 45.11
5 InstructBlip 43.36
6 Shikra 43.13
7 Shikra-VQA 42.01
8 LLAVA-1.5-13B 41.59
9 MiniGPT-4 38.64
10 Otter 33.95
(h) Pr.Att.-Celebrity
Rank Model Score
  1 XComposer-VL 83.60
2 Blip2 81.67
3 InstructBlip 75.94
4 Qwen-VL 60.43
5 Shikra-VQA 58.61
6 LLAVA-1.5-13B 56.20
7 Shikra 56.08
8 MiniGPT-4 46.70
9 LLAVA-1.5-7B 45.20
10 Otter 38.39
(i) Pr.Att.-Food
Rank Model Score
  1 XComposer-VL 82.18
2 Blip2 82.16
3 InstructBlip 74.20
4 Qwen-VL 60.29
5 Shikra 54.03
6 LLAVA-1.5-13B 53.77
7 Shikra-VQA 52.86
8 MiniGPT-4 46.45
9 LLAVA-1.5-7B 42.07
10 Otter 39.52
(j) Pr.Att.-Plant
Rank Model Score
  1 XComposer-VL 60.37
2 LLAVA-1.5-7B 42.53
3 Blip2 42.20
4 LLAVA-1.5-13B 41.30
5 MiniGPT-4 39.18
6 Shikra 38.98
7 Qwen-VL 37.56
8 Shikra-VQA 37.47
9 InstructBlip 35.75
10 Otter 32.18
(k) Pr.Att.-Age
Rank Model Score
  1 Blip2 93.97
2 InstructBlip 80.44
3 XComposer-VL 79.56
4 LLAVA-1.5-13B 74.03
5 Shikra-VQA 63.62
6 Shikra 62.06
7 Qwen-VL 56.05
8 MiniGPT-4 51.87
9 LLAVA-1.5-7B 51.33
10 Otter 39.31
(l) Pr.Att.-Gender
Rank Model Score
  1 XComposer-VL 78.59
2 Blip2 67.94
3 InstructBlip 67.72
4 Shikra-VQA 65.00
5 Shikra 61.51
6 Qwen-VL 59.16
7 LLAVA-1.5-13B 58.33
8 LLAVA-1.5-7B 53.00
9 MiniGPT-4 45.37
10 Otter 39.91
(m) Pr.Att.-Expression
Rank Model Score
  1 XComposer-VL 62.21
2 Blip2 52.98
3 InstructBlip 48.42
4 LLAVA-1.5-13B 47.80
5 Qwen-VL 44.81
6 Shikra-VQA 42.09
7 LLAVA-1.5-7B 41.75
8 Shikra 41.07
9 MiniGPT-4 38.03
10 Otter 28.77
(n) Pr.Att.-Race
Rank Model Score
  1 Blip2 90.07
2 XComposer-VL 87.80
3 InstructBlip 82.43
4 Qwen-VL 65.20
5 LLAVA-1.5-13B 60.13
6 Shikra-VQA 55.53
7 Shikra 53.98
8 MiniGPT-4 48.02
9 Otter 46.92
10 LLAVA-1.5-7B 44.73
(o) Pr.Att.-Animal
Rank Model Score
  1 Blip2 83.55
2 XComposer-VL 81.03
3 InstructBlip 76.06
4 Qwen-VL 61.06
5 LLAVA-1.5-13B 56.84
6 Shikra-VQA 55.02
7 MiniGPT-4 49.61
8 Shikra 46.96
9 LLAVA-1.5-7B 44.26
10 Otter 38.32
(p) Pr.Att.-Object
Rank Model Score
  1 XComposer-VL 60.60
2 Blip2 52.84
3 LLAVA-1.5-13B 45.26
4 Qwen-VL 43.55
5 LLAVA-1.5-7B 42.23
6 Shikra 41.06
7 MiniGPT-4 40.32
8 Shikra-VQA 40.12
9 InstructBlip 40.03
10 Otter 33.86
(q) Pr.Att.-Style
Rank Model Score
  1 XComposer-VL 64.99
2 Shikra-VQA 55.50
3 Blip2 54.94
4 Shikra 54.16
5 LLAVA-1.5-13B 51.16
6 InstructBlip 49.98
7 Qwen-VL 49.17
8 LLAVA-1.5-7B 44.34
9 MiniGPT-4 39.23
10 Otter 37.94
(r) Pr.Att.-Background
Rank Model Score
  1 XComposer-VL 78.66
2 Blip2 74.47
3 InstructBlip 63.59
4 Qwen-VL 58.40
5 Shikra 51.75
6 Shikra-VQA 51.10
7 LLAVA-1.5-13B 49.48
8 LLAVA-1.5-7B 43.40
9 MiniGPT-4 41.70
10 Otter 35.56
(s) Pr.Att.-Color
Table 8: Leaderboards of all Perceptual tasks under print attacking scenario.
Rank Model Score
  1 Qwen-VL 58.17
2 Shikra 57.41
3 Shikra-VQA 57.26
4 LLAVA-1.5-13B 57.17
5 Otter 56.61
6 LLAVA-1.5-7B 51.91
7 MiniGPT-4 36.95
8 Blip2 33.33
9 InstructBlip 31.97
10 XComposer-VL 31.75
(a) Ad.Att.-Movie
Rank Model Score
  1 Shikra-VQA 85.84
2 Qwen-VL 79.11
3 Shikra 77.17
4 LLAVA-1.5-13B 66.37
5 Otter 65.81
6 LLAVA-1.5-7B 60.84
7 MiniGPT-4 37.59
8 Blip2 34.64
9 XComposer-VL 33.88
10 InstructBlip 32.85
(b) Ad.Att.-Action
Rank Model Score
  1 Qwen-VL 55.79
2 Shikra-VQA 53.57
3 Otter 53.05
4 LLAVA-1.5-13B 53.03
5 Shikra 52.61
6 LLAVA-1.5-7B 48.36
7 MiniGPT-4 38.22
8 XComposer-VL 34.45
9 Blip2 32.04
10 InstructBlip 30.93
(c) Ad.Att.-TV Show
Rank Model Score
  1 Shikra-VQA 71.47
2 Shikra 67.66
3 Qwen-VL 64.80
4 Otter 61.44
5 LLAVA-1.5-13B 59.83
6 LLAVA-1.5-7B 56.15
7 MiniGPT-4 36.95
8 Blip2 36.14
9 InstructBlip 35.02
10 XComposer-VL 30.86
(d) Ad.Att.-Profession
Rank Model Score
  1 Shikra-VQA 78.40
2 Shikra 77.85
3 Qwen-VL 76.34
4 LLAVA-1.5-13B 71.92
5 LLAVA-1.5-7B 68.62
6 Otter 56.87
7 MiniGPT-4 35.73
8 XComposer-VL 35.33
9 Blip2 34.53
10 InstructBlip 32.03
(e) Ad.Att.-Landmark
Rank Model Score
  1 Qwen-VL 60.07
2 Otter 51.86
3 LLAVA-1.5-13B 51.81
4 Shikra-VQA 48.45
5 Shikra 47.75
6 LLAVA-1.5-7B 47.32
7 MiniGPT-4 38.38
8 Blip2 37.95
9 InstructBlip 34.29
10 XComposer-VL 30.89
(f) Ad.Att.-Anime
Rank Model Score
  1 Qwen-VL 69.92
2 Shikra-VQA 68.32
3 Shikra 65.73
4 LLAVA-1.5-13B 58.83
5 Otter 57.58
6 LLAVA-1.5-7B 47.05
7 MiniGPT-4 38.05
8 InstructBlip 36.65
9 XComposer-VL 36.41
10 Blip2 36.18
(g) Ad.Att.-Clothes
Rank Model Score
  1 Qwen-VL 64.77
2 Shikra 60.02
3 Shikra-VQA 59.70
4 LLAVA-1.5-13B 55.53
5 LLAVA-1.5-7B 54.95
6 Otter 52.26
7 MiniGPT-4 37.48
8 Blip2 34.72
9 InstructBlip 33.52
10 XComposer-VL 33.09
(h) Ad.Att.-Celebrity
Rank Model Score
  1 Shikra-VQA 75.26
2 Qwen-VL 73.72
3 Shikra 71.07
4 LLAVA-1.5-13B 64.95
5 Otter 64.39
6 LLAVA-1.5-7B 52.30
7 MiniGPT-4 36.94
8 XComposer-VL 33.16
9 Blip2 32.48
10 InstructBlip 32.13
(i) Ad.Att.-Food
Rank Model Score
  1 Qwen-VL 79.59
2 Shikra-VQA 76.06
3 Shikra 73.53
4 Otter 67.16
5 LLAVA-1.5-13B 64.44
6 LLAVA-1.5-7B 47.41
7 MiniGPT-4 37.74
8 Blip2 35.84
9 InstructBlip 34.02
10 XComposer-VL 34.00
(j) Ad.Att.-Plant
Rank Model Score
  1 Shikra 56.95
2 Shikra-VQA 56.80
3 LLAVA-1.5-13B 51.43
4 Qwen-VL 48.11
5 LLAVA-1.5-7B 45.89
6 Otter 42.82
7 MiniGPT-4 37.14
8 Blip2 34.16
9 InstructBlip 33.14
10 XComposer-VL 30.78
(k) Ad.Att.-Age
Rank Model Score
  1 LLAVA-1.5-13B 89.35
2 Shikra-VQA 89.06
3 Shikra 85.34
4 Otter 82.25
5 Qwen-VL 79.18
6 LLAVA-1.5-7B 51.95
7 MiniGPT-4 48.38
8 Blip2 39.48
9 InstructBlip 38.51
10 XComposer-VL 38.09
(l) Ad.Att.-Gender
Rank Model Score
  1 Shikra-VQA 82.04
2 Shikra 79.86
3 Qwen-VL 73.91
4 LLAVA-1.5-13B 63.13
5 Otter 60.44
6 LLAVA-1.5-7B 55.70
7 MiniGPT-4 41.07
8 Blip2 35.95
9 InstructBlip 33.52
10 XComposer-VL 32.80
(m) Ad.Att.-Expression
Rank Model Score
  1 LLAVA-1.5-13B 63.06
2 Shikra-VQA 61.19
3 Shikra 60.24
4 Qwen-VL 59.50
5 LLAVA-1.5-7B 45.97
6 Otter 45.50
7 MiniGPT-4 37.47
8 Blip2 36.81
9 InstructBlip 36.36
10 XComposer-VL 34.77
(n) Ad.Att.-Race
Rank Model Score
  1 Qwen-VL 81.78
2 Shikra-VQA 80.44
3 Otter 78.00
4 Shikra 77.66
5 LLAVA-1.5-13B 70.94
6 LLAVA-1.5-7B 49.30
7 MiniGPT-4 37.61
8 Blip2 35.23
9 XComposer-VL 34.38
10 InstructBlip 34.01
(o) Ad.Att.-Animal
Rank Model Score
  1 Shikra-VQA 79.05
2 Qwen-VL 77.26
3 Shikra 69.98
4 LLAVA-1.5-13B 65.33
5 Otter 62.69
6 LLAVA-1.5-7B 52.62
7 MiniGPT-4 37.84
8 Blip2 37.62
9 XComposer-VL 36.97
10 InstructBlip 36.88
(p) Ad.Att.-Object
Rank Model Score
  1 Shikra-VQA 79.12
2 Shikra 77.62
3 Qwen-VL 76.50
4 Otter 74.91
5 LLAVA-1.5-13B 66.43
6 XComposer-VL 53.21
7 InstructBlip 50.03
8 Blip2 47.80
9 LLAVA-1.5-7B 46.78
10 MiniGPT-4 36.78
(q) Ad.Att.-OCR
Rank Model Score
  1 Qwen-VL 56.80
2 Otter 53.47
3 Shikra 51.05
4 Shikra-VQA 50.30
5 LLAVA-1.5-13B 50.01
6 LLAVA-1.5-7B 44.85
7 MiniGPT-4 37.98
8 XComposer-VL 37.20
9 InstructBlip 36.20
10 Blip2 35.87
(r) Ad.Att.-Style
Rank Model Score
  1 Shikra-VQA 65.41
2 Shikra 64.50
3 Otter 60.30
4 Qwen-VL 60.09
5 LLAVA-1.5-13B 56.28
6 LLAVA-1.5-7B 46.53
7 MiniGPT-4 36.91
8 Blip2 35.05
9 XComposer-VL 34.71
10 InstructBlip 34.45
(s) Ad.Att.-Background
Rank Model Score
  1 Qwen-VL 69.63
2 Shikra 69.57
3 Shikra-VQA 69.05
4 LLAVA-1.5-13B 54.28
5 Otter 53.12
6 LLAVA-1.5-7B 46.08
7 MiniGPT-4 37.54
8 XComposer-VL 37.23
9 Blip2 36.83
10 InstructBlip 35.81
(t) Ad.Att.-Color
Table 9: Leaderboards of all Perceptual tasks under adversarial attacking scenario.
Rank Model Score
  1 XComposer-VL 72.37
2 InstructBlip 67.41
3 Blip2 66.13
4 Qwen-VL-Chat 62.68
5 Shikra 62.52
6 LLAVA-1.5-13B 59.28
7 Shikra-VQA 58.64
8 Otter 55.49
9 LLAVA-1.5-7B 51.75
10 MiniGPT-4 42.64
(a) Clean
Rank Model Score
  1 XComposer-VL 71.63
2 InstructBlip 67.62
3 Blip2 65.49
4 Shikra 62.37
5 Qwen-VL-Chat 60.58
6 LLAVA-1.5-13B 58.73
7 Shikra-VQA 58.47
8 Otter 55.42
9 LLAVA-1.5-7B 46.44
10 MiniGPT-4 42.87
(b) Corru.
Rank Model Score
  1 XComposer-VL 63.75
2 Blip2 56.20
3 Qwen-VL-Chat 51.57
4 InstructBlip 51.23
5 LLAVA-1.5-13B 50.70
6 Shikra 49.08
7 LLAVA-1.5-7B 46.47
8 Shikra-VQA 43.72
9 MiniGPT-4 42.87
10 Otter 47.10
(c) Print Attack
Rank Model Score
  1 Qwen-VL-Chat 60.39
2 Shikra 59.31
3 LLAVA-1.5-13B 56.84
4 Shikra-VQA 55.16
5 Otter 51.94
6 LLAVA-1.5-7B 50.07
7 InstructBlip 33.13
8 MiniGPT-4 32.95
9 Blip2 32.65
10 XComposer-VL 30.52
(d) Adversarial Attack
Table 10: Leaderboards of comprehensive performance in clean, corruption, print attack and adversarial attack scenario.
The model performance leaderboards for each subtask under each scenario are shown in Tab. 6, Tab. 7, Tab. 8, and Tab. 9. We compute the average of the multi-choice and true-or-false evaluation results under each subtask. Since the free-form question allows to assess the model’s perception abilities across multiple subtasks at the same time, the results of free-form are not taken into account.
The overall performances on each scenario are displayed in Tab. 10. We calculate the average of the scores of the three questioning types (i.e., multi-choice, true-or-false and free-form) as the final score.
Appendix B Discussion
B.1 General Discussion
Limitation. Dysca is the dynamic and scalable benchmark, offering evaluation for 20 perceptual subtasks under 51 image styles and 4 scenarios. However, generating data for evaluating cognition abilities (e.g., commonsense reasoning) presents challenge within the existing framework. This limitation arises from the reliance on predefined rules for prompt and question generation, which may not adequately capture the complexity of cognitive-level questions.
Synthesis Data for Training / Fine-tuning. The use of synthetic data for model training / fine-tuning has been adopted in the field of Natural Language Processing (NLP) [65]. In this work, we do not explore the possibility of utilizing our benchmark for model training. Our primary goal in this paper is to provide a large-scale evaluation benchmark that addresses the issue of data leakage in current multimodal evaluation benchmarks and offers evaluation results across multiple subtasks, scenarios, question types and styles. Nevertheless, considering that Dysca has the capability to synthesize high-resolution and unlimited amounts of annotated multimodal data, we believe that Dysca also holds potential as a training data synthesis tool for LVLMs.
Reproducibility and Licence. All the experiments are built on 8 * RTX 4090. All the data and the code for generation and evaluation are released in https://github.com/Benchmark-Dysca/Dysca. The licence of Dysca is "CreativeML Open RAIL++-M", which follows the licence set by the Stable Diffusion XL.
Ethical Concerns. Our Dysca leverages the Stable Diffusion XL [36] to generate images. In order to prevent the model generating unsafe images, e.g., NSFW and offensive images, lots of efforts have been made. First, we use the safety checker [66] to post filter the unsafe images. With the unsafe image that is recognized by the safety checker, the model’s output will be a blank image. Besides, we manually exclude the specific styles or the word that may trigger the unsafe images generation from the Metadata $M$. We believe that our Dysca involves fewer ethical concerns.
B.2 The Stability of Dysca
Figure 6: The tendency of model’s overall performance under clean scenario with different scale of evaluation data.
In this section, we focus on examining the stability of Dysca. We partition Dysca into 11 different scales: 1%, 10%, 20%, 30%, 40%, 50%, 60%, 70%, 80%, 90% and 100%. We compute the evaluation scores using each of these data scales. The score is calculated as the sum of scores obtained from multiple-choice, true-or-false and free-form questions. As can be seen in Fig. 6, when the evaluation data scale is less than 30% of Dysca (i.e., less than 46.8K samples), the evaluation score show significant fluctuations. When the data scale exceeds 40%, we obtain the stable results, reflecting current scale of Dysca achieves the stable and reliable evaluation results. Although 40% evaluation scale of Dysca has achieved stable scores, Dysca aims to provide more than just stable rankings, but also draws on massive amounts of data to provide in-depth feedback across different image styles and perceptual subtasks.
Appendix C The Metadata ($M$) of Dysca
Metadata ($M$) is the core of Dysca, which is randomly assembled from our collected source material and contains all the information needed to generate prompt ($P$), image ($I$), and question-answer pairs($Q$). Specifically, the metadata is a data container that contains information in multiple dimensions about the foreground, the attributes corresponding to the foreground, the background, and the artistic style required to generate an image. Therefore, each instance of M is mapped one-to-one to a prompt, an image, and a set of question-answer pairs, respectively.
In order to ensure the quality and stability of the generated images, we carefully select the source material. First, for each perceptual subtask, we collect rich annotation material as described in Section 3.2. However, the metadata composed of these raw annotations is not always usable. On the one hand, some of the content is polysemous, which can easily be misinterpreted by the model’s when generating images. On the other hand, there are backgrounds or artistic styles (e.g., “Pokemon Style", “architectural style", etc.) that negatively affect the quality of the image and do not accurately generate the desired content. In order to test the usability of these source materials, we went through several small-scale pre-generations covering all the source materials. After careful selection, we retain the clips that consistently produced high-quality images. The detailed information of the source materials are shown in Tab. 11.
Table 11: Detailed information of the source materials.
Category Data Description #Numbers
Style We collected artistic styles from the community, which can be well rendered by Stable Diffusion model. We removed those which have strong reflect on the image content or may generate unsafe image. 51
Background We have selected 20 rich backgrounds and they can be accurately generated by Stable Diffusion Model. 20
Age We chose four well-characterized age nodes: 14, 25, 40, and 80. 4
Expression We chose three characteristic expressions: smiling happily, shouting furiously, calm and placid. 3
Gender Male and Female. 2
Race We identified these five races based on the ethnicity that can be generated by Stable Diffusion: Caucasian, Asian, African, Indian, and Middle Eastern races 5
Profession After pre-generation and careful selection, we chose 20 occupations with distinctive characteristics that are easy to generate. 20
Action After pre-generation and careful selection, we chose 20 occupations with distinctive characteristics that are easy to generate. 20
Celebrity After pre-generation and careful selection, we chose 50 well-known celebrities. 50
Animal We selected a rich variety of animals, including mammals, birds, reptiles, insects, and aquatic animals, and can they be generated by the Stable Diffusion model accurately. 67
Plant We selected a rich variety of plants, including flowers, trees, fruits, and vegetables, and they can be generated by the Stable Diffusion model accurately. 37
Clothes We selected 16 common types of clothing that are highly distinguishable from each other. 16
Object We took the annotations from the MSCOCO [50] dataset after removing people, animals, and plants, and added some additional common objects. 80
Landmark We chose 23 characteristic landmarks from around the globe and they can be generated by the Stable Diffusion model accurately. 23
Food We collected 29 special dishes from around the globe and they can be generated by the Stable Diffusion model accurately. 29
Movie We selected 106 movies titles from the rating list of IMDb based on the number of user reviews. 106
Anime We selected 44 anime titles from the rating list of IMDb based on the number of user reviews. 44
TV shows We selected 20 TV show titles from the rating list of IMDb based on the number of user reviews. 20
OCR We randomly selected 5000 words from the IELTS vocabulary for the text material. Among them, words with length less than 3 were removed. 5000
Color We selected 8 easily distinguishable colors: red, orange, yellow, green, blue, purple, white, black. 8
Appendix D Scenarios Details
D.1 Print Attack Scenario
Followed by the settings in [35], we add the attack text on the images. Consider that the image resolution in our Dysca is much higher than the one in [35], we increase more font form in terms of font position and font orientation. Fig. 7 to Fig. 11 shows the detailed examples.
Figure 7: Images with different font size. ’112px’ means that the typos are 112 pixels in size.
Figure 8: Images with different font color. ’Blue’ means that the color of typos are in blue.
Figure 9: Images with different font opacity. ’40%’ means that the transparency of typos are 40% and ’100%’ implies full opacity of typos.
Figure 10: Images with different font orientation. ’$15^{\circ}$’ means that the orientation of typos are $15^{\circ}$ and $0^{\circ}$ implies the typos are horizontal.
Figure 11: Images with different font position. An image is divided into a grid of 5 rows and 5 columns, leading to 25 sections. ’R1C1’ means the typo is located in row 1, column 1, which is the top left corner of the image.
D.2 Corrupted Scenario
Examples of the 11 image corruptions are shown in Fig 12.
Figure 12: Examples of 11 image corruptions applied to a single clean image.
Appendix E More Examples of Dysca
For each subject we collected in Metadata ($M$), we display one example of their prompt ($P$), generated image ($I$) and corresponding question-answer pairs ($Q$).
Figure 13: Plant
Figure 14: Profession
Figure 15: Celebrity
Figure 16: Action
Figure 17: Landmark
Figure 18: Face
Figure 19: Animal
Figure 20: Object
Figure 21: Clothes
Figure 22: Food
Figure 23: Movie
Figure 24: TV
Figure 25: Anime
Figure 26: OCR
Appendix F Data Sheet
We follow the documentation frameworks provided by Gebru et al. [67].
F.1 Motivation
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
  • The proposed dataset is used for evaluating current LVLMs perception ability. We use the synthesis images to prevent the potential data leakage problem in current benchmarks. The dataset test LVLMs in 20 subtasks under 4 scenarios and 3 question type, revealing the existing drawbacks of current LVLMs.
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
  • The dataset is created by the AI Safety and Trustworthiness Group on behalf of Key Laboratory of AI Safety of CAS, Institute of Computing Technology.
Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grant or and the grant name and number.
  • It is funded by the Key Laboratory of AI Safety of CAS, Institute of Computing Technology.
F.2 Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
How many instances are there in total (of each type, if appropriate)?
  • There are a total of 20 subtasks in our Dysca. For details of each subtasks please see refer Fig. 2.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
  • No. The images in Dysca are completely generated from scratch.
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
  • Each instance consists of the prompt, the image generated by stable diffusion, the question and corresponding answer.
Is there a label or target associated with each instance? If so, please provide a description.
  • Yes, Dysca provides the ground truth for each instance.
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
  • No.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
  • There are no relationships between individual instances.
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
  • Following our motivation, the entire proposed dataset is used for testing purposes.
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
  • Errors in image generation resulting from stable diffusion are unavoidable. However, we have performed dataset cleaning to minimize these errors. Furthermore, the stability experiment in Appendix B demonstrates that these errors do not affect the overall evaluation results of the dataset.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a dataset consumer? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
  • The proposed Dysca dose not rely on any external resources.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)? If so, please provide a description.
  • No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
  • No. To ensure that the generated images do not contain offensive, insulting, threatening, or anxiety-inducing content, we manually filter out words from the metadata $M$ that could potentially trigger the diffusion model to generate such images. Safety checker also used to further avoid unsafe image generation.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
  • Yes.
Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
  • Yes. There are the age, gender and race recognition subtasks in Dysca. Each of them are divided to several subpopulations and the selection of these subpopulations is based on the ability of stable diffusion to generate the representative subpopulations.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
  • Yes. There is the celebrity recognition task in our dataset, where 50 well-know celebrity are chosen. We choose the celebrity who can be generated well by stable diffusion XL.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals race or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
  • No, our benchmark does not contain any sensitive data.
F.3 Collection Process
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
  • We display the detailed explanation in Tab. 11.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
  • We collect the data by manual human curation.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
  • No.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
  • We collect the metadata of Tab. 11 by authors. The images are generated by stable diffusion and labels of each image are also automatically generated.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
  • Our dataset was conducted in April of 2024, but the results do not depend on the date of data collection.
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
  • No.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
  • No.
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
  • N/A. Our Dysca does not involve the collection from the individuals.
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
  • N/A. Our Dysca does not involve the collection from the individuals.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
  • N/A. Our Dysca does not involve the collection from the individuals.
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
  • No.
F.4 Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remaining questions in this section.
  • Yes. We leverage the off-the-shelf models, i.e., PP-OCRv3 [56] and CLIP-L-14 [16], to clean the data. PP-OCRv3 [56] is leveraged as the filter to exclude the failure image that TextDiffusion2 [58] generates the wrong text on the image. For the other images, we use CLIP-L-14 [16] to filter out the images with low text-image consistency.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
  • Yes. We have saved all the data. However, most of these images are filtered and considered to be useless.
Is the software that was used to preprocess/clean/label the data available? If so, please provide a link or other access point.
F.5 Uses
Has the dataset been used for any tasks already? If so, please provide a description.
  • No. The proposed dataset is the novel one which is used for evaluation current LVLMs perception ability.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
  • Yes. We plan to create a section on the project homepage to keep track of LVLMs papers for researchers to analyze and compare.
What (other) tasks could the dataset be used for?
  • In this work, we do not explore the possibility of utilizing our benchmark for model training / fine-tuning. Our primary goal in this paper is to provide a large-scale evaluation benchmark that addresses the issue of data leakage in current multimodal evaluation benchmarks and offers evaluation results across multiple subtasks, scenarios, question types and styles. Nevertheless, considering that Dysca has the capability to synthesize high-resolution and unlimited amounts of annotated multimodal data, we believe that Dysca also holds potential as a training data synthesis tool for LVLMs.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a dataset consumer might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other risks or harms (e.g., legal risks, financial harms)? If so, please provide a description. Is there anything a dataset consumer could do to mitigate these risks or harms?
  • Yes.
Are there tasks for which the dataset should not be used? If so, please provide a description.
  • The proposed dataset should not be used to generate offensive data.
F.6 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
  • Yes.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
  • We will open-source our dataset on our GitHub project homepage. At the moment, we do not have a DOI number.
When will the dataset be distributed?
  • The dataset can be downloaded right now.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
  • No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
  • Not yet.
F.7 Maintenance
Who will be supporting/hosting/maintaining the dataset?
  • Jie Zhang and Zhongqi Wang are hosting and maintaining the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
  • Email: zhangjie@ict.ac.cn
Is there an erratum? If so, please provide a link or other access point.
  • No.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to dataset consumers (e.g., mailing list, GitHub)?
  • There are no plans at the moment, but if there are updates, they will be announced, and the download source will be updated on the project homepage.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were the individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
  • No.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to dataset consumers.
  • Yes. If there are any updates, the previous version of the dataset will also be shared on website for download.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to dataset consumers? If so, please provide a description.
  • Yes. We welcome and encourage researchers to extend/augment/build on/contribute to our dataset for non-profit purposes without the need for prior notification.

### v3

@@
 Thanks to the generative paradigm, Dysca serves as a scalable benchmark for easily adding new subtasks and scenarios.
-A total of 8 advanced open-source LVLMs with 10 checkpoints are evaluated on Dysca, revealing the drawbacks of current LVLMs.
-The benchmark is released in https://github.com/Benchmark-Dysca/Dysca.
-footnotetext: Equal contribution.
-Corresponding Author: zhangjie@ict.ac.cn
+A total of 24 advanced open-source LVLMs and 2 close-source LVLMs are evaluated on Dysca, revealing the drawbacks of current LVLMs.
+The benchmark is released at https://github.com/Robin-WZQ/Dysca.
 1 Introduction
-Recent years have witnessed the great success of the Large Vision-Language Models (LVLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
-These models leverage the powerful Large Language Models (LLMs) [11, 12, 13, 14, 15] as their brain and incorporate the state-of-the-art visual encoders [16, 17, 18] as their eyes.
-Thanks to the alignment of visual feature with textual space and the development of visual instruction tuning techniques [4], LVLMs showcase the impressive capability in terms of visual scene comprehension and multimodal instruction-following.
-In order to comprehensively evaluate the capabilities of LVLMs, many benchmarks have been purposed [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], where we categorize the current benchmarks into three types [25].
-The first type is the classical benchmarks, such as COCO Caption [30] and VQA [19, 31, 32].
+Recent years have witnessed the great success of the Large Vision-Language Models (LVLMs) (Li et al., 2023d; Zhu et al., 2023; Dai et al., 2023; Liu et al., 2023b; Li et al., 2023a; Chen et al., 2023b; Zhang et al., 2023; Su et al., 2023; Gong et al., 2023; Sun et al., 2023b).
+These models leverage the powerful Large Language Models (LLMs) (Chung et al., 2022; OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023; FastChat, 2023) as their brain and incorporate the state-of-the-art visual encoders (Radford et al., 2021; Fang et al., 2023; Dosovitskiy et al., 2020) as their eyes.
+Thanks to the alignment of visual feature with textual space and the development of visual instruction tuning techniques (Liu et al., 2023b), LVLMs showcase the impressive capability in terms of visual scene comprehension and multimodal instruction-following.
+In order to comprehensively evaluate the capabilities of LVLMs, many benchmarks have been purposed (Antol et al., 2015a; Singh et al., 2019; Xu et al., 2023; Shao et al., 2023; Li et al., 2023c; b; Fu et al., 2023; Bai et al., 2023b; Yu et al., 2023; Yang et al., 2023b; Chen et al., 2024), where we categorize the current benchmarks into three types (Fu et al., 2023).
+The first type is the classical benchmarks, such as COCO Caption (Chen et al., 2015) and VQA (Antol et al., 2015a; Goyal et al., 2017; Marino et al., 2019).
 Although these benchmarks provide high-quality evaluation data, they also have notable limitations.
@@
 On the other hand, since these classical benchmarks have been available as the open-source test data for a long time, it is hard to prevent the data leakage problem.
-The second type of benchmarks evaluate the LVLMs through a subjective manner [28, 33].
+The second type of benchmarks evaluate the LVLMs through a subjective manner (Yang et al., 2023b; Wu et al., 2023).
 Although the benchmarks reveal the insightful drawbacks of current models, their data scale is limited (i.e., less than 200 annotations) and they require manual evaluation by experts.
@@
 They provide an objective and automatic evaluation manner, giving the fine-grained evaluation for the LVLMs.
-However, these benchmarks conduct Vision-language QAs by selecting images from existing dataset.
-Although they claim that the questions are re-annotated, the previous work [29] has demonstrated that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs.
+However, these benchmarks conduct Vision-language QAs by selecting images from existing dataset and annotate the textual questions.
+Although they claim that the questions are re-annotated, the previous work (Chen et al., 2024) has demonstrated that these benchmarks have Models unintentionally leaked into the training data of LLMs and LVLMs.
 Besides, most benchmarks focus on evaluating LVLMs in the realistic images and clean scenarios, leaving the multi-stylized images and noisy scenarios unexplored.
-While some works like MMCBench [34] and Typographic Dataset [35] have investigated the robustness of LVLMs with corrupted and print-attacked images, respectively, they have not explored the effect of these noisy images on various perceptual tasks.
+While some works like MMCBench (Zhang et al., 2024b) and Typographic Dataset (Cheng et al., 2024) have investigated the robustness of LVLMs with corrupted and print-attacked images, respectively, they have not explored the effect of these noisy images on various perceptual tasks.
 Figure 1: Overview of the automatic pipeline for generating Vision-language QAs, cleaning Vision-language QAs and evaluating LVLMs.
-(a) We first constructs prompts in terms of content, style and background, leveraging the Text-to-Image (T2I) diffusion model (e.g., SDXL [36]) to synthesis images to be asked.
+(a) We first constructs prompts in terms of content, style and background, leveraging the Text-to-Image (T2I) diffusion model (e.g., SDXL (Podell et al., 2023)) to synthesis images to be asked.
 Then based on the scenarios and the question type, we post-process the synthesis images and generate the specific textual questions, respectively.
@@
 In this paper, aiming to address these challenges above, we propose Dysca which is a dynamic and scalable benchmark for evaluating the perception ability of LVLMs via various subtasks and scenarios.
-Inspired by the prior evaluation works for LLMs [37], we investigate on whether we could leverage the large-scale synthesized images for evaluating LVLMs.
+Inspired by the prior evaluation works for LLMs (Liang et al., 2023), we investigate on whether we could leverage the large-scale synthesized images for evaluating LVLMs.
 We display the overview of our pipeline in Fig.
@@
 3.
-Then we utilize the state-of-the-art text-to-image diffusion models (e.g., SDXL [36]) to generate the corresponding images.
+Then we utilize the state-of-the-art text-to-image diffusion models (e.g., SDXL (Podell et al., 2023)) to generate the corresponding images.
 Since we already know the main information of the images through prompts, we easily generate question-answer textual pairs by the rule-based method.
-After that, in order to obtain the high quality Vision-language QAs, we employ CLIP [16] to perform data cleaning on the generated Vision-language QA pairs.
+After that, in order to obtain the high quality Vision-language QAs, we employ CLIP (Radford et al., 2021) to perform data cleaning on the generated Vision-language QA pairs.
 Dysca focuses on assessing the fine-grained perception abilities, including recognizing human, animal, object, landmark, etc.
@@
 Besides, to evaluate the robustness of the models across different scenarios and question types, we construct 4 testing scenarios (clean, corruption, print attacking and adversarial attacking) and 3 question types (multi-choices, true-or-false and free-form questions).
-In the end, Dysca consists of 617K Vision-language QA pairs ($\times$20 larger than MM-BigBench [38] and $\times$25 larger than Seed-Bench2 [24] as shown in Tab.
-1).
-Thanks to the generative paradigm, Dysca achieves the scalable benchmark to new subtasks and scenarios and dynamically generate unlimited Vision-language QAs for evaluation.
+Compared to previous works in Tab.
+1, we provide an end-to-end process from image to Vision-QA generation.
+The approach significantly reduces annotation costs compared to manually labeling images (e.g., MME (Fu et al., 2023)) while achieving the correctness for evaluating LVLMs.
+It also avoids the risk of hallucinate annotations that may occur when using ChatGPT for labeling based on image prompts (e.g., JourneyDB (Sun et al., 2023a)).
+This novel pipeline enables us to create a benchmark that is easily scalable and adaptable for incorporating new subtasks and scenarios.
+Thanks to the generative paradigm, Dysca can be customized to meet the specific requirements of the evaluator for testing purposes.
 In summary, our work makes the following key contributions:
   • Dynamic and Scalable Benchmark: We propose Dysca, a benchmark that is able to dynamically generate the test data that users need and is easily to scale up to to new subtasks and scenarios.
-  • Multi-grained Perceptual Subtasks and Multi-scenarios: Dysca evaluates the 20 perceptual subtasks performance of 8 mainstream LVLMs with 10 checkpoints under 4 image scenarios (i.e., clean, corruption, print attacking and adversarial attacking) and 3 question types (i.e., multi-choices, true-or-false and free-form questions).
+  • Multi-grained Perceptual Subtasks and Multi-scenarios: Dysca aims to testing LVLMs’ performance on diverse styles, 4 image scenarios (i.e., clean, corruption, print attacking and adversarial attacking) and 3 question types (i.e., multi-choices, true-or-false and free-form questions), reporting the 20 perceptual subtasks performance of 26 mainstream LVLMs, including GPT-4o (Ope, ) and Gemini-1.5-Pro (Team et al., 2024).
   • Analysis and Observations: We demonstrate for the first time that evaluating LVLMs using large-scale synthetic data is valid.
@@
 Annotation
-Collecting from Existing Datasets Question Type
+Novel Images &\&& Novel Questions Question Type
 Automatic
 Evaluation
-LLaVA-Bench [4] 0.15K - Free-form
-MME [25] 2.3K 10 True-or-false
-LVLM-eHub [21] - 3 Free-form
-tiny-LVLM-eHub [22] 2.1K 3 Free-form
-SEED-Bench [23] 19K 8 Multi-choices
-MMBench [39] 2.9K 12 Multi-choices
-TouchStone [26] 0.9K 10 Free-form
-REFORM-EVAL [40] 50K 7 Multi-choices
-MM-BigBench [38] 30K 6 Multi-choices
-MM-VET [27] 0.2K 4 Free-form
-MLLM-Bench [41] 0.42K 7 Free-form
-SEED-Bench2 [24] 24K 10 Multi-choices
-BenchLMM [42] 2.4K 15 Free-form
-JourneyDB [43] 5.4K 2
+LLaVA-Bench 0.15K - Free-form
+MME 2.3K 10 True-or-false
+LVLM-eHub - 3 Free-form
+tiny-LVLM-eHub 2.1K 3 Free-form
+SEED-Bench 19K 8 Multi-choices
+MMBench 2.9K 12 Multi-choices
+TouchStone 0.9K 10 Free-form
+REFORM-EVAL 50K 7 Multi-choices
+MM-BigBench 30K 6 Multi-choices
+MM-VET 0.2K 4 Free-form
+MLLM-Bench 0.42K 7 Free-form
+SEED-Bench2 24K 10 Multi-choices
+BenchLMM 2.4K 15 Free-form
+JourneyDB 5.4K 2
 Free-form
@@
 2.1 Large Vision-Language Models
-The landscape of Large Vision-Language Models (LVLMs) has been significantly shaped by the pioneering success of Large Language Models (LLMs) such as GPTs [44, 45, 46] and LLaMA [13], catalyzing advancements in multimodal content understanding and generation [47], including intricate tasks like image-text comprehension.
-At the forefront of these developments, BLIP-2 [1] introduces a lightweight Q-Former [1] that facilitates alignment between textual and visual representations through a cross-attention mechanism [1].
-InstructBLIP [3] takes a step further by incorporating textual instructions into the Q-Former, which significantly improves performance.
-LLAVA [4] employs GPT-4 [14] to transform data into multimodal instruction-following data and uses CLIP [16] and LLAMA [13] for fine-tuning instructions, achieving advanced performance.
-LLAVA-1.5 [48] extends this paradigm by integrating MLP projection and introducing academic task-specific Vision-language QA data.
-Recently, models like Otter [5], MiniGPT-4 [2], Qwen-VL-Chat [49] and XComposer-VL [7] further unleash the cross-modal understanding capabilities of LVLMs.
+The landscape of Large Vision-Language Models (LVLMs) has been significantly shaped by the pioneering success of Large Language Models (LLMs) such as GPTs (Radford et al., 2019; Brown et al., 2020; Ouyang et al., 2022) and LLaMA (Touvron et al., 2023), catalyzing advancements in multimodal content understanding and generation (Zhang et al., 2024a), including intricate tasks like image-text comprehension.
+At the forefront of these developments, BLIP-2 (Li et al., 2023d) introduces a lightweight Q-Former (Li et al., 2023d) that facilitates alignment between textual and visual representations through a cross-attention mechanism (Li et al., 2023d).
+InstructBLIP (Dai et al., 2023) takes a step further by incorporating textual instructions into the Q-Former, which significantly improves performance.
+LLAVA (Liu et al., 2023b) employs GPT-4 (OpenAI, 2023) to transform data into multimodal instruction-following data and uses CLIP (Radford et al., 2021) and LLAMA (Touvron et al., 2023) for fine-tuning instructions, achieving advanced performance.
+LLAVA-1.5 (Liu et al., 2023a) extends this paradigm by integrating MLP projection and introducing academic task-specific Vision-language QA data.
+Recently, models like Otter (Li et al., 2023a), MiniGPT-4 (Zhu et al., 2023), Qwen-VL-Chat (Bai et al., 2023a) and XComposer-VL (Zhang et al., 2023) further unleash the cross-modal understanding capabilities of LVLMs.
+Besides, many powerful closed-source LVLMs, including Gemini-1.5-Pro (Team et al., 2024) and GPT-4o (Ope, ), have publicly released their APIs, promoting the development of downstream applications.
 2.2 Benchmarks for LVLMs
 The great progress of LVLMs triggers the development of benchmarks for evaluating these models, where we divide previous benchmarks into three categories.
-The first type is the classical benchmarks which focuses on evaluating LVLMs abilities via image caption [50] and VQA [51, 19].
+The first type is the classical benchmarks which focuses on evaluating LVLMs abilities via image caption (Lin et al., 2014) and VQA (Antol et al., 2015b; a).
 However, these benchmarks cannot provide the fine-grained feedback on how to improve the models.
 Besides, since these benchmarks have been the public resources for a long time, it is hardly to guarantee that the LVLMs have not use them for training.
-The second type subjectively evaluates LVLMs by experts [28, 33].
+The second type subjectively evaluates LVLMs by experts (Yang et al., 2023b; Wu et al., 2023).
 Although these benchmarks reveal the insightful feedback of the LVLMs, their scale is limited (i.e., less than 200 annotations).
-The subjective manner also makes the evaluation expensive and hardly to expand the scale of the benchmarks.
-The third type [4, 25, 21, 22, 23, 24, 39, 26, 40, 38, 27, 41, 42, 29, 52] focuses on evaluating LVLMs in an objective and large-scaled manner, where we list the detailed information of them in the Tab.
+The subjective manner also makes the evaluation expensive and hardly to expand the scale.
+The third type (Liu et al., 2023b; Fu et al., 2023; Xu et al., 2023; Shao et al., 2023; Li et al., 2023c; b; Liu et al., 2023c; Bai et al., 2023b; Li et al., 2023f; Yang et al., 2023a; Yu et al., 2023; Ge et al., 2023; Cai et al., 2023; Chen et al., 2024; Liu et al., 2024) focuses on evaluating LVLMs in an objective and large-scaled manner, where we list the detailed information of them in the Tab.
 1.
-Some of them have been adopted by the community [53] as the standard benchmarks for evaluating LVLMs [12, 5, 7], like MME [25] and MMBench [39].
+Some of them have been adopted by the community (Contributors, 2023) as the standard benchmarks for evaluating LVLMs (OpenAI, 2022; Zhang et al., 2023), like MME (Fu et al., 2023) and MMBench (Liu et al., 2023c).
 These benchmarks evaluate models through the objective answer types and most of them leverage the automatic annotation and evaluation manner for revealing the fine-grained drawbacks of current LVLMs.
 However, the previous benchmarks primarily concentrate on evaluating LVLMs using realistic images and clean scenario, leaving multi-stylized images and noisy scenarios unexplored.
-Moreover, many of them conduct QA by selecting images from publicly available datasets (e.g., [50, 54]).
+Moreover, many of them conduct QA by selecting images from publicly available datasets (e.g., (Lin et al., 2014; Russakovsky et al., 2014)).
 While they state that the questions have been re-annotated, they cannot guarantee that the LVLMs have not seen the image during training stage.
-The previous work [29] has proved that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs.
-One possible way to solve data leakage is using novel but synthesis images, where JourneyDB [43] is the first work aiming to leverage synthesis images to evaluate current LVLMs.
-The prompts and the corresponding images are downloaded from Midjourney [55] and ChatGPT [12] is leveraged to label the images.
+The previous work (Chen et al., 2024) has proved that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs.
+One possible way to solve data leakage is using novel but synthesis images, where JourneyDB (Sun et al., 2023a) is the first work aiming to leverage synthesis images to evaluate current LVLMs.
+The prompts and the corresponding images are downloaded from Midjourney (mid, ) and ChatGPT (OpenAI, 2022) is leveraged to label the images.
 However, JourneyDB is a top-down framework where the number of images is fixed.
 Besides, the ChatGPT labeling may cause hallucinate annotations, leading to the unreliable evaluation results.
-Although 40 annotators have involved to clean the data, the data cleaning cost are expensive and it limits the data scale.
-In contrast, our Dysca serves as the bottom-up framework, allowing for dynamic and scalable generation for both images and evaluation questions.
+In contrast, our Dysca serves as the bottom-up framework, allowing for dynamic and scalable generation for both images and evaluation questions and involving more evaluation tasks.
 The rule-based question generation method also makes the annotations more accuracy.
-Besides, Dysca contains 20 subtasks which is more comprehensive than JourneyDB.
 3 Dysca
@@
 1, containing data generation, data cleaning and LVLMs evaluation.
-For the data generation, our Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, where $M$ means "Metadata", $P$ means "Prompt", $I$ means "Image" and $Q$ means "Question-answer pair".
-We further decouple the metadata $M$ into 4 parts, i.e., "style", "attribute", "foreground" and "background", and the combination of the four parts constitute the image prompts $P$.
-Then, given the prompt $P$ and the selected scenario, we leverage the Text-to-Image (T2I) diffusion model (e.g., SDXL [36]) to synthesis image $I$ and add the specific perturbation to the image $I$.
+For the data generation, our Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, where $M$ means “Metadata”, $P$ means “Prompt”, $I$ means “Image” and $Q$ means “Question-answer pair”.
+We further decouple the metadata $M$ into 4 parts, i.e., “style”, “attribute”, “foreground” and “background”, and the combination of the four parts constitute the image prompts $P$.
+Then, given the prompt $P$ and the selected scenario, we leverage the Text-to-Image (T2I) diffusion model (e.g., SDXL (Podell et al., 2023)) to synthesis image $I$ and add the specific perturbation to the image $I$.
 After that, since the prompt already includes the question angle and the corresponding answer, we construct a rule-based approach to generate the $Q$.
@@
 Multi-choice and true-or-false questions utilize a closed-ended manner to assess LVLMs, while free-form questions employ an open-ended manner through image captioning for evaluation.
-For the data cleaning, considering that the T2I diffusion model may generate unsuccessful outcomes, we then use CLIP [16] and PP-OCRv3 [56] to automatically clean the whole dataset to obtain the final Dysca.
-Finally, we evaluate 8 open-source LVLMs with 10 checkpoints on our proposed Dysca.
+For the data cleaning, considering that the T2I diffusion model may generate unsuccessful outcomes, we then use CLIP (Radford et al., 2021) and PP-OCRv3 (Li et al., 2022) to automatically clean the whole dataset to obtain the final Dysca.
+Finally, we evaluate 14 open-sourced LVLMs and 2 closed-source LVLMs on our proposed Dysca.
+Table 2: Key statistics of Dysca.
 Statistic #Number
@@
 Average choice number 3.0
-Table 2: Key statistics of Dysca.
 Figure 2: Overview of the dataset distribution of 20 perceptual tasks. The number in each subtask shows the corresponding amount of their annotation.
@@
 Evaluation dimensions.
-Perception is one of the most fundamental capabilities of LVLMs and previous works [25] have shown that the lack of perceptual ability may result in hallucination [57].
-In order to comprehensively evaluate LVLMs’ perception capability, we design 20 perceptual subtasks where we show all subtasks and the corresponding amount of their annotation in the Fig.
+Perception is one of the most fundamental capabilities of LVLMs and previous works (Fu et al., 2023) have shown that the lack of perceptual ability may result in hallucination (Li et al., 2023e).
+In order to comprehensively evaluate LVLMs’ perception capability, we collect and organize existing sub-dimensions from current evaluation datasets, resulting in a total of 20 assessment dimensions where we show all the subtasks and the corresponding amount of their annotation in the Fig.
 2.
@@
 For each perceptual subtask, we collect the textual data first to construct the metadata $M$.
-For the TV shows, Anime and Movie, we select the titles from the rating list of IMDb***https://www.imdb.com/ based on the number of user reviews.
-For the styles, we utilize the style lists collected from the communityhttps://stable-diffusion-art.com/sdxl-styles/ and remove those which have strong reflect on the image content like "architectural style" and "Pokemon style".
+For the TV shows, Anime and Movie, we select the titles from the rating list of IMDb [^1] https://www.imdb.com/ based on the number of user reviews.
+For the styles, we utilize the style lists collected from the community [^2] https://stable-diffusion-art.com/sdxl-styles/ and remove those which have strong reflect on the image content like “architectural style” and “Pokemon style”.
 Note that the style list does not include the style prompt associated with a particular artist’s name.
-Besides, for the remaining contents, we select them from the label of current dataset (e.g., ImageNet [54]).
+Besides, for the remaining contents, we select them from the label of current dataset (e.g., ImageNet (Russakovsky et al., 2014)).
 All the selected textual data above constitute the metadata $M$.
-We provide the detailed information of the metadata in the Appendix C.
+We provide the detailed information of the metadata in the Appendix E.
 3.3 Construction of Questions & Answers
@@
 Then, we utilize the T2I diffusion model to generate the corresponding image using the prompt $P$.
-For generating the image with specific text on it for the OCR subtask, we leverage TextDiffusion2 [58], which is the state-of-the-art text rendering method.
-For the rest of images, we leverage Stable Diffusion XL [36].
-Subsequently, based on the different question types we select, i.e., multi-choices, true-or-false and free-form, we generate the corresponding Vision-language QA pairs in Dysca.
+For generating the image with specific text on it for the OCR subtask, we leverage TextDiffusion2 (Chen et al., 2023a), which is the state-of-the-art text rendering method.
+For the rest of images, we leverage Stable Diffusion XL (Podell et al., 2023).
+Subsequently, based on the different question types we select, i.e., multi-choices, true-or-false and free-form, we generate the corresponding VQA pairs in Dysca.
 Besides, in order to evaluate the model performance under various scenarios, we conduct experiments on 4 scenarios, i.e., clean, corruption, print attacking and adversarial attacking.
-For the print attacking, followed by [35], we add the deceptive text on the image, where the text is a wrong option.
+For the print attacking, followed by (Cheng et al., 2024), we add the deceptive text on the image, where the text is a wrong option.
 Besides, to comprehensively evaluate the performance of LVLMs under corruption scenario, we add more typographic factors to original settings (i.e., different font orientations and font positions).
-For the adversarial attacking, we leverage PGD [59] to generate the adversarial image.
-We use InstructBLIP [3] as the proxy model and regard others as the black box models.
+For the adversarial attacking, we leverage PGD (Madry et al., 2017) to generate the adversarial image.
+We use InstructBLIP (Dai et al., 2023) as the proxy model and regard others as the black box models.
 The reason why we choose InstructBLIP is that it has shown superior performance in clean scenario.
 Besides, the black-box setting better reflects the robustness of the models when they face the real-world adversarial attacks.
-For the corruption, we leverage the image corruption methods collected from [34].
+For the corruption, we leverage the image corruption methods collected from (Zhang et al., 2024b).
 We remove some hard corruptions as they significantly impact the quality of the image, leading to human failure in judging the style and content of the image.
-The detailed examples are shown in Appendix D.
-Consider that the Text-to-image diffusion model may generate the failure cases that affect the quality of the proposed benchmark, we leverage the off-the-shelf models, i.e., PP-OCRv3 [56] and CLIP-L-14 [16], to clean the data.
-PP-OCRv3 [56] is leveraged as the filter to exclude the failure image that TextDiffusion2 [58] generates the wrong text on the image.
-For the other images, we use CLIP-L-14 [16] to filter out the images with low text-image consistency.
-In the end, we filter out nearly 15% of low quality samples.
+The detailed examples are shown in Appendix F.
+Data Clean.
+To ensure the quality of Dysca, four steps are adopt: 1) First, we manually remove difficult-to-generate foregrounds and attributes, along with backgrounds and styles that could heavily affect image content.
+We believe this process can serve as a coarse-grained method to eliminate samples that are highly likely to be generated incorrectly.
+2) Then, we leverage the off-the-shelf models, i.e., PP-OCRv3 (Li et al., 2022) and CLIP-L-14 (Radford et al., 2021), to clean the data.
+PP-OCRv3 (Li et al., 2022) is leveraged as the filter to exclude the failure image that TextDiffusion2 (Chen et al., 2023a) generates the wrong text on the image.
+For the other images, we use CLIP-L-14 (Radford et al., 2021) with a threshold of 0.75 to filter out the images with low text-image consistency.
+We find that using 0.75 as the threshold achieves a good balance between image correctness and data scale.
+3) After that, We select the top six performing models and eliminate any question-answer pairs where the models either answer incorrectly or indicate that the answer was not included among the options.
+We observe that nearly 100% of the samples filtered out by these models are incorrect.
+4) Finally, we analyze the patterns in these incorrect samples, removing the associated vocabulary from our metadata and discarding all related samples.
+By meticulously refining the metadata manually and utilizing automated tools to assist in question filtering, Dysca ensures high-quality data synthesis.
+In the end, we filter out nearly 40% of low quality samples.
 The final statistics of our released Dysca are shown in Tab.
@@
 […]
-Table 3: Evaluation results on the 20 perceptual subtasks.
-The top two results on each subtask are bolded and underlined, respectively.
-"MC" and "TF" indicate the accuracy (%) of "Multi-choices" and "True-or-false", respectively.
-Movie Action TV Show Profession Landmark
-Model Language Model MC TF MC TF MC TF MC TF MC TF
-Blip2 [1] Flan-T5-XL 71.46 68.61 96.24 93.04 56.29 61.99 77.88 75.87 98.70 94.58
-InstructBLIP [3] Flan-T5-XL 77.35 67.38 97.28 93.61 65.89 60.96 79.17 76.51 98.04 95.75
-XComposer-VL [7] InternLM-7B 81.90 78.36 97.03 94.75 77.81 73.29 83.97 75.87 98.04 94.81
-LLava-1.5 [48] Vicuna-7B 55.95 51.68 74.78 50.75 56.29 50.68 58.65 56.83 85.00 54.48
-LLava-1.5 [48] Vicuna-13B 66.29 54.04 77.93 59.71 59.27 54.11 66.67 59.05 94.35 55.90
-MiniGPT-4 [2] Vicuna-7B 32.68 52.35 42.96 51.00 30.46 50.00 33.97 50.16 53.26 52.36
-Otter [5] LLaMA-7B 65.36 59.08 65.85 72.36 68.54 57.53 66.99 54.92 58.48 66.51
-Qwen-VL-Chat [49] Qwen-7B 71.25 51.01 96.35 63.14 67.22 49.32 75.64 54.60 96.96 54.72
-Shikra [6] Vicuna-7B 66.29 59.64 78.00 77.29 60.26 56.85 78.85 68.25 89.35 70.99
-Shikra-VQA [6] Vicuna-7B 66.39 61.66 96.17 80.32 59.93 60.27 78.21 67.94 92.83 72.88
-Anime Clothes Celebrity Food Plant
-Model Language Model MC TF MC TF MC TF MC TF MC TF
-Blip2 [1] Flan-T5-XL 57.07 61.94 82.38 73.79 81.98 76.48 91.52 88.04 92.50 91.19
-InstructBLIP [3] Flan-T5-XL 61.28 64.04 86.51 81.08 83.69 76.14 91.71 88.85 93.31 91.92
-XComposer-VL [7] InternLM-7B 75.41 74.58 86.18 85.73 88.53 87.28 92.33 90.21 93.25 93.14
-LLava-1.5 [48] Vicuna-7B 47.15 48.46 47.19 50.41 57.36 54.72 54.63 51.10 49.08 48.59
-LLava-1.5 [48] Vicuna-13B 57.61 50.14 65.20 57.20 60.14 57.94 78.87 57.45 79.72 55.51
-MiniGPT-4 [2] Vicuna-7B 29.21 48.74 31.25 50.47 28.62 49.64 44.66 50.78 45.38 50.53
-Otter [5] LLaMA-7B 61.82 57.30 47.13 69.52 42.41 63.52 46.81 79.62 66.41 81.51
-Qwen-VL-Chat [49] Qwen-7B 71.60 54.63 78.44 60.72 87.37 50.05 89.37 53.37 92.32 59.91
-Shikra [6] Vicuna-7B 47.96 57.72 75.21 59.65 63.76 59.86 83.30 68.37 88.55 66.60
-Shikra-VQA [6] Vicuna-7B 49.05 57.58 76.57 63.10 63.73 62.13 89.25 71.22 88.34 71.76
-Age Gender Expression Race Animal
-Model Language Model MC TF MC TF MC TF MC TF MC TF
-Blip2 [1] Flan-T5-XL 62.61 59.65 99.37 94.86 89.27 75.92 74.38 71.95 96.64 95.27
-InstructBLIP [3] Flan-T5-XL 65.14 59.55 99.51 88.21 91.85 79.34 77.99 74.62 97.21 94.98
-XComposer-VL [7] InternLM-7B 67.89 78.35 99.60 98.06 89.64 82.52 80.42 76.69 97.83 96.70
-LLava-1.5 [48] Vicuna-7B 38.35 55.15 54.14 49.78 63.86 49.76 43.98 50.56 49.26 50.90
-LLava-1.5 [48] Vicuna-13B 49.55 59.59 98.71 83.53 71.29 58.33 70.16 62.84 85.99 58.23
-MiniGPT-4 [2] Vicuna-7B 31.75 51.71 56.49 49.22 42.01 50.91 27.54 50.41 45.97 49.57
-Otter [5] LLaMA-7B 37.98 51.19 78.23 77.96 73.86 60.36 43.00 57.88 81.32 83.63
-Qwen-VL-Chat [49] Qwen-7B 53.48 49.09 97.72 58.55 85.22 64.30 73.50 56.88 95.60 64.37
-Shikra [6] Vicuna-7B 65.37 56.71 97.85 73.16 90.47 70.55 74.17 54.88 89.82 69.06
-Shikra-VQA [6] Vicuna-7B 65.50 57.34 99.39 81.41 92.36 74.71 73.90 55.24 91.18 74.26
-Object OCR Style Background Color
-Model Language Model MC TF MC TF MC TF MC TF MC TF
-Blip2 [1] Flan-T5-XL 89.32 87.75 71.89 62.07 99.86 88.62 63.97 68.37 88.01 85.55
-InstructBLIP [3] Flan-T5-XL 89.87 90.13 72.61 60.93 99.96 81.23 66.50 69.19 90.93 85.55
-XComposer-VL [7] InternLM-7B 90.23 91.72 69.22 78.66 95.83 81.32 71.26 75.27 87.97 89.02
-LLava-1.5 [48] Vicuna-7B 58.63 49.99 47.08 51.02 46.35 50.49 44.33 50.95 41.91 51.90
-LLava-1.5 [48] Vicuna-13B 77.86 57.93 58.19 53.17 85.80 51.56 62.67 55.33 63.81 52.90
-MiniGPT-4 [2] Vicuna-7B 51.75 51.55 31.37 51.56 43.41 48.64 31.94 50.10 35.29 50.51
-Otter [5] LLaMA-7B 47.31 82.51 66.23 61.41 98.04 61.58 63.06 63.40 48.28 57.97
-Qwen-VL-Chat [49] Qwen-7B 88.32 63.14 71.74 51.86 98.21 52.43 69.42 51.59 86.39 53.05
-Shikra [6] Vicuna-7B 70.91 69.22 60.79 53.63 95.13 58.95 71.24 62.66 84.91 61.83
-Shikra-VQA [6] Vicuna-7B 89.43 76.00 59.63 54.14 99.51 61.48 70.98 65.96 83.79 64.49
+Figure 4: The failure cases for the noisy scenarios. From left to right are: corruption scenario, adversarial attacking scenario, and print attacking scenario.
 […]
+Besides, thanks to generative evaluation framework of Dysca, we are able to effectively control variables and conduct a detailed analysis of the model’s fine-grained capabilities.
+Specifically, we introduce two novel metrics to measure the sensitivity of LVLMs on question types and covariate shift.
+The sensitivity to question type aims to evaluate whether LVLMs exhibit inconsistent performance when facing different types of question types (i.e., multi-choice vs.
+true-or-false).
+We first normalize the score of true-or-false by $TFSQ=\frac{S-50}{100-50}*100\%$ and multi-choices by $MCSQ=\frac{S-25}{100-25}*100\%$, where $S$ denotes the score of the models in each question type.
+The sensitivity of LVLMs to question types is defined by:
+$SQ=\frac{(TFSQ-SQ_{Avg})^{2}+(MCSQ-SQ_{Avg})^{2}}{2},$ (1)
+where $SQ_{Avg}=\frac{TFSQ+MCSQ}{2}$.
+The sensitivity to covariate shift aims to evaluate whether LVLMs exhibit inconsistent performance when facing the same content and question format, but with variations in image covariates (i.e., image style).
+It is defined by:
+$SC=\frac{\sum_{i=1}^{N}(S_{i}-SC_{Avg})^{2}}{N},$ (2)
+where $SC_{Avg}=\frac{\sum_{i=1}^{N}S}{N}$ and $S_{i}$ denotes the score of the models in each style. Since there are 51 image styles in Dysca, we set $N=51$.
 4 Results and Analysis
 In this section, we report the evaluation results and make insightful analysis.
-A total of 8 LVLMs with 10 checkpoints are evaluated on Dysca benchmark, including BLIP2 [1], InstructBLIP [3], LLavA [48], MiniGPT-4 [2], Otter [5], XComposer-VL [7], Qwen-VL-chat [49] and Shikra [6].
+A total of 26 LVLMs are evaluated on Dysca, including BLIP2 (Li et al., 2023d), InstructBLIP (Dai et al., 2023), LLavA (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2023), Otter (Li et al., 2023a), XComposer-VL (Zhang et al., 2023), Qwen-VL-chat (Bai et al., 2023a), Shikra (Chen et al., 2023b), Emu2-Chat (Sun et al., 2024), GLM-4V (GLM et al., 2024), MiniCPM-v2.5 (Yao et al., 2024), Yi-VL (AI et al., 2024), mPLUG-Owl-2 (Ye et al., 2023), Phi-3-Vision (Abdin et al., 2024), GPT-4o (OpenAI, 2023), Gemini-1.5-pro (Team et al., 2024).
 Each model is evaluated with all the 20 perception subtasks under 4 scenarios.
@@
 4.1 Main Results
+Blind Setting.
+We first evaluate LVLMs when only textual questions are provided.
+As shown in the “Blind” column of Tab.
+3, all LVLMs yield consistent results on the Dysca and perform comparable to random guessing.
+This outcome demonstrates that the generated paradigm employed by Dysca effectively mitigates the potential impact of data leakage  (Chen et al., 2024), thereby enhancing the fairness of the evaluation results.
+Additional comparisons can be found in Appendix D.
 Clean Scenario.
-The evaluation results of various LVLMs in different perceptual subtasks under clean scenarios are presented in Tab.
+The evaluation results of various LVLMs in different perceptual subtasks under clean scenarios are presented in the “clean” column of Tab.
 3.
-Since the evaluation for free-form question type usually involves multiple subtasks, we can not calculate the results of free-form for each subtask individually.
-Instead, we display the overall score of free-form in the first row in Tab.
+We calculate the average score of 3 question types.
+As can be seen, GLM-4V (GLM et al., 2024) outperforms other LVLMs, achieving top-1 performance.
+MiniCPM-v2.5 (Yao et al., 2024), Xcomposer2-VL (Dong et al., 2024) and Gemini-1.5-pro (Team et al., 2024) also perform well.
+It is evident that for the latest large models, their scores remain below 85.
+The results highlight that all existing LVLMs still struggle to provide accurate responses to questions formulated by Dysca.
+Noisy Scenarios.
+The evaluation results of various LVLMs under noisy scenarios (i.e., corruption, print attacking and adversarial attacking) are presented in last 3 columns in Tab.
+3.
+The value in the brackets shows the relative values with respect to the ones in the clean scenario.
+As can be seen, GLM-4V (GLM et al., 2024) still takes a lead on corruption and adversarial attacking scenarios.
+For the print attacking scenario, mPLUG-Owl-2 (Ye et al., 2023) performs the best.
+Here, we present a failure case sample for each of the three different scenarios on Fig.
 4.
-As can be seen, Xcomposer-VL [7] outperforms other LVLMs, achieving top-1 or top-2 results in most subtasks, but InstructBLIP [3], Qwen-VL-chat [49] and BLIP2 also take a lead in a few subtasks.
-Noisy Scenarios The evaluation results of various LVLMs under noisy scenarios (i.e., corruption, print attacking and adversarial attacking) are presented in Tab.
-4.
-As can be seen, For the multi-choices and true-or-false question type, Xcomposer-VL [7] still takes a lead on all 4 scenarios.
-For the free-form, LLava-1.5-7b [48] achieves the best.
-Table 4: Evaluation results on 4 scenarios.
-"MC", "TF" and "FF" indicates "Multi-choices", "True-or-false" and "Free-form", respectively.
-"PrintAtt" and "AdverAtt" means "Print Attacking" and "Adversarial Attacking", respectively.
-"*": the model is under white-box setting.
-Scenarios Blip2 [1] InstructBLIP [3] XComposer-VL [7] LLava-1.5-13b [48] LLava-1.5-7b [48]
-MC TF FF MC TF FF MC TF FF MC TF FF MC TF FF
-Clean 82.07 78.78 37.56 84.29 79.00 38.94 86.22 84.82 46.03 71.50 57.72 48.61 53.70 51.41 50.15
-Corruption 81.17 77.98 37.32 83.77 78.57 38.72 85.20 84.16 45.52 70.34 57.57 48.28 53.21 51.40 49.54
-PrintAtt 64.78 68.01 35.82 59.18 58.80 35.71 74.08 72.71 44.46 48.91 54.72 48.47 40.02 50.94 48.38
-AdverAtt 23.28 48.98 25.70 21.49* 48.74* 29.15* 20.85 49.54 21.17 66.86 56.56 47.11 50.70 51.36 48.17
-Scenarios MiniGPT-4 [2] Otter [5] Qwen-VL-Chat [49] Shikra [6] Shikra-VQA [6]
-MC TF FF MC TF FF MC TF FF MC TF FF MC TF FF
-Clean 38.50 50.51 38.90 61.36 65.99 39.12 82.31 55.84 49.90 76.61 63.79 47.17 79.31 66.69 29.93
-Corruption 38.70 49.99 40.35 62.19 65.36 38.71 79.26 52.74 49.74 76.42 64.31 46.43 79.06 66.95 29.40
-PrintAtt 36.16 50.00 42.45 38.11 36.27 36.93 59.68 46.40 48.65 60.02 40.18 47.03 62.20 41.05 27.91
-AdverAtt 26.75 49.52 22.59 57.78 62.27 35.76 78.47 58.04 44.67 72.20 62.16 43.56 74.04 64.64 26.81
-4.2 Analysis
-4.2.1 Key Observations
-(1) For individual models, their perceptual performance varies across different subtasks.
-For example, in the case of Qwen-VL-Chat [49], it achieves a score of 96.96% accuracy in landmark recognition task for multi-choices questions (2% below the first-place score), but obtains a score of 53.48% accuracy in age recognition task (12% below the first-place score).
-The results suggest that Qwen-VL-Chat [49] may require more fine-tuning in age perception data.
-Analyzing the performance of the models across various subtasks contribute to purposive improving.
-Figure 4: Models exhibit different performance when facing the same image but different question types.
-(2) Models exhibit performance inconsistency when facing multi-choices and true-or-false question types.
-As can be seen, in the object recognition subtask, Otter [5] achieves an accuracy of 47.31% in the multi-choices question type (22% higher than random guessing), but obtains an accuracy of 82.51% in the true-or-false question type (32% higher than random guessing).
-Interestingly, we observe the opposite results in the Qwen-VL-Chat [49].
-In the object recognition subtask, it achieves an accuracy of 88.32% in the multi-choices but achieves an accuracy of 63.14% in the true-of-false.
-We also observe the same problem shown in other models and we display two examples in Fig.
-4.
-We speculate that the inconsistency may be attributed to the bias in the training dataset towards particular question types, such as using more multi-choices or true-or-false questions.
-(3) Each model exhibits robustness in the corruption scenario, but suffers degradation in the two attacking scenarios.
-As shown, all models exhibit minor score variation of less than 1% under the corruption scenario.
-However, they exhibit degradation when facing print attacking (e.g., 84.29% vs.
-59.18% for InstructBlip [3] in multi-choices accuracy).
-XComposer-VL [7] shows the strongest robustness, maintaining over 70% accuracy for both multi-choices and true-or-false.
-Besides, since our adversarial algorithm specifically targets the image encoder, the LVLMs that share the same encoder architecture (i.e., Blip2, InstructBLip and XComposer-VL all using EVA-CLIP [62] as the image encoder) exhibit significant performance degradation, with accuracy even falling below random selection.
-Models utilizing alternative image encoders also experience a performance decrease of approximately 5% to 10%.
-A more detailed result can be found in Appendix D.
-4.2.2 The Validity of Dysca
+Table 3: Evaluation results of 26 LVLMs, where the darker colors represent better performance.
+The top 1 result on each column are bolded and the value in brackets is the relative values with respect to the ones in the clean scenario.
+“PrintAtt” and “AdverAtt” means “Print Attacking” and “Adversarial Attacking”, respectively.
+“*”: the model is under white-box setting.
+Scenarios
+Model LLM Visual Encoder Blind Clean $\uparrow$ Corruption $\uparrow$ AdverAtt $\uparrow$ PrintAtt $\uparrow$ SQ $\downarrow$ SC $\downarrow$
+MiniGPT-4 Vicuna-7B EVA-CLIP ViT-G 35.37 41.38 42.30 (+0.92) 34.42 (-6.96) 42.71 (+1.33) 193.6 1.7
+MiniGPT-4 Vicuna-13B EVA-CLIP ViT-G 35.21 50.17 49.63 (-0.54) 31.77 (-18.40) 47.55 (-2.62) 758.0 2.5
+MiniGPT-4 LLaMA2-7B EVA-CLIP ViT-G 34.77 56.61 55.70 (-0.91) 33.55 (-23.06) 49.78 (-6.83) 1344.5 3.0
+MiniGPT-2 LLaMA2-7B EVA-CLIP ViT-G 35.28 58.46 58.06 (-0.40) 56.62 (-1.84) 52.96 (-5.50) 1512.1 4.3
+BLIP2 Flan-T5-XL EVA-CLIP ViT-G 35.35 65.30 66.09 (+0.79) 32.55 (-32.75) 57.01 (-8.29) 165.5 8.5
+BLIP2 OPT-3B EVA-CLIP ViT-G 34.99 39.54 40.29 (+0.75) 30.62 (-8.92) 37.26 (-2.28) 110.7 2.3
+BLIP2 OPT-7B EVA-CLIP ViT-G 35.21 39.55 41.12 (+1.57) 31.76 (-7.79) 38.82 (-0.73) 50.2 1.8
+InstructBLIP Vicuna-7B EVA-CLIP ViT-G 35.14 67.54 67.01 (-0.53) 34.42 (-33.12) 52.58 (-14.96) 1287.1 2.8
+InstructBLIP Vicuna-13B EVA-CLIP ViT-G 34.37 64.89 64.68 (-0.21) 31.77 (-33.12) 53.53 (-11.36) 1252.4 2.5
+InstructBLIP Flan-T5-XL EVA-CLIP ViT-G 34.51 66.54 67.58 (+1.04) 32.95 (-33.59)* 52.09 (-14.45) 271.3 6.3
+InstructBLIP Flan-T5-XXL EVA-CLIP ViT-G 34.82 68.65 69.79 (+1.14) 32.95 (-35.70) 57.73 (-10.92) 215.0 6.8
+LLava-1.5 Vicuna-7B CLIP ViT-L 34.63 51.27 51.70 (+0.43) 49.62 (-1.65) 47.27 (-4.00) 788.1 3.2
+LLava-1.5 Vicuna-13B CLIP ViT-L 35.21 59.23 59.58 (+0.35) 56.87 (-2.36) 51.69 (-7.54) 912.5 7.7
+Otter LLaMA-7B CLIP ViT-L 35.19 54.90 56.02 (+1.12) 51.42 (-3.48) 37.78 (-17.12) 427.4 6.0
+Shikra LLaMA-7B CLIP ViT-L 34.96 62.24 63.06 (+0.82) 58.78 (-3.46) 49.56 (-12.68) 440.6 5.0
+Xcomposer-VL InternLM-7B EVA-CLIP ViT-G 32.33 71.40 72.08 (+0.68) 30.28 (-41.12) 64.71 (-6.69) 147.9 6.1
+Xcomposer2-VL InternLM2-7B CLIP ViT-L 32.76 79.13 78.64 (-0.49) 76.60 (-2.53) 66.34 (-12.79) 20.9 1.9
+Qwen-VL-Chat Qwen-7B OpenClip ViT-bigG 33.06 62.18 61.05 (-1.13) 59.85 (-2.33) 51.94 (-10.24) 1885.4 6.7
+Emu2-Chat LLaMA-33B EVA2-CLIP-E 35.14 63.64 62.81 (-0.83) 61.90 (-1.74) 54.82 (-8.82) 2497.9 1.6
+GLM-4V GLM-4-9B-Chat EVA2-CLIP-E 35.08 82.09 81.95 (-0.14) 80.72 (-1.37) 52.09 (-30.00) 25.5 0.8
+MiniCPM-V2.5 Llama3-Instruct 8B SigLIP SoViT-400m 34.99 78.75 77.41 (-1.34) 75.44 (-3.31) 60.77 (-17.98) 38.3 1.7
+Yi-VL Yi-6B-Chat OpenClip ViT-H 35.01 75.71 74.94 (-0.77) 72.53 (-3.18) 64.97 (-10.74) 233.1 1.9
+mPLUG-Owl-2 LLaMA2-7B CLIP ViT-L 35.03 74.09 72.85 (-1.24) 69.76 (-4.33) 72.85 (-1.24) 180.9 2.2
+Phi-3-Vision Phi-3 CLIP ViT-L 34.74 73.23 72.11 (-1.12) 69.66 (-3.57) 57.78 (-15.45) 292.7 1.5
+GPT-4o / / 35.02 75.69 75.52 (-0.17) 73.47 (-2.22) 56.34 (-19.35) 67.4 1.9
+Gemini-1.5-Pro / / 34.55 77.79 77.12 (-0.67) 75.89 (-1.90) 61.05 (-16.74) 439.6 1.6
+4.2 Key Observations
+(1) For LVLMs, the capacity of the language model plays a crucial role.
+When using the same visual encoder, models that adopt a language model with larger parameter sizes (e.g., Vicuna-13B models generally outperform Vicuna-7B models by 8%) or models with a different but stronger architecture (e.g., GLM-4V (GLM et al., 2024) with GLM-4-9B-Chat shows a 20% performance increase compared to Emu2 (Sun et al., 2024) with LLaMA-33B) tend to achieve better performance.
+(2) Each model shows robustness in the corruption scenario, but experiences significant degradation in both attack scenarios.
+In the image corruption scenario, all models demonstrate minimal score variations (less than 1%), indicating consistent performance under non-targeted disruptions.
+However, under print attacks, performance drops are notable.
+For example, two closed-source models experience substantial declines: Gemini-1.5-pro (Team et al., 2024) drops by 21.5% (from 77.79 to 61.05), and GPT-4o (OpenAI, 2023) suffers a 25.8% decrease (from 75.69 to 56.10).
+Among the advanced open-source models, GLM-4V’s (GLM et al., 2024) performance decreases sharply by 36.5% (from 82.09 to 52.09), and Phi-3-Vision (Abdin et al., 2024) records a 21.39% drop (from 73.23 to 57.78).
+Notably, mPLUG-Owl-2 (Ye et al., 2023) shows the highest robustness, with only a 1.7% reduction (from 74.09 to 72.84), while the XComposer-VL (Zhang et al., 2023; Dong et al., 2024) series also exhibits strong resistance to print attacks.
+In the adversarial attack scenario, where the attack algorithm directly targets the image encoder, LVLMs with shared encoder architecture (e.g., Blip2 (Li et al., 2023d), InstructBLIP (Dai et al., 2023), and XComposer-VL (Zhang et al., 2023), all using EVA-CLIP (Fang et al., 2022) as their image encoder) suffer significant performance drops, with some models even performing worse than random chance.
+For instance, XComposer-VL (Zhang et al., 2023) experiences a 57.6% drop (from 71.40 to 30.28).
+Models with alternative image encoders also experience degradation, ranging from 1% to 5%.
+The greater impact of adversarial noise compared to corruption noise suggests that adversarial attacks exhibit a certain degree of transferability across different model architectures.
+More detailed results are available in Appendix F.
+Figure 5: The Visualization of hierarchical clustering results of 20 subtasks.
+Figure 6: The Visualization of hierarchical clustering results of 51 animal categories.
+(3) Models exhibit varying sensitivity to different question types and covariate shifts.
+For the Sensitivity to Question Types, Xcomposer2-VL (Dong et al., 2024) achieves the best result with a score of 20.9.
+However, we observe that the perception ability of LVLM does not show a positive correlation with sensitivity to question types.
+For instance, while the GLM-4V (GLM et al., 2024) achieves the highest performance in evaluation tasks, it exhibits higher sensitivity to question types than Xcomposer2-VL (Dong et al., 2024).
+One of the factors influencing the sensitivity to question types may be the inherent biases of the language model.
+Using the same base language model may result in similar outcomes for this metric.
+it is also noted that Gemini (Team et al., 2024) performs bad in this metric, revealing its preference for certain question types.
+For the Sensitivity to Covariate Shifts, GLM-4V (Dong et al., 2024) achieves the best result with a score of 0.8.
+However, we also observe that the perception ability of LVLM does not show a positive correlation with covariate shifts.
+For example, InstructBLIP-Flan-T5-XXL (Dai et al., 2023) outperforms InstructBLIP-Flan-T5-XL in terms of performance but shows higher sensitivity to covariate shifts.
+4.3 Analysis on inter-task and intra-task
+Analysis on Inter-task.
+Here, we perform hierarchical clustering based on the Euclidean distance of the scores across 20 evaluation dimensions for 26 models.
+As shown in the Fig.
+6, based on the consistency of model performance across the 20 evaluation dimensions, the models can be categorized into four tiers.
+We also observe that LVLMs tend to perform better in dimensions that involve well-defined image perception, such as landmark recognition and object recognition.
+However, dimensions likes style recognition and movie recognition exhibit greater challenges for LVLMs, which may be attributed to the limited training resources in these specific domains.
+Besides, commercial models exhibit poor performance in tasks related to “people” (i.e., “race”, “age”, and “gender”).
+This is likely due to the additional safety training incorporated into closed-source models, which results in more conservative responses on related questions.
+Detailed results are provided in Appendix A.
+Analysis on Intra-task.
+Thanks to the diverse metadata design, Dysca enables highly granular analysis within a single task.
+Taking animal categories as an example, we analyze the performance of LVLMs on different animals.
+Here, we perform same hierarchical clustering methods across 51 animal categories.
+As shown in the Fig.
+6, LVLMs performance vary significantly across these categories.
+Models tend to perform poorly when applied to marine life.
+This could be caused by the challenges in collecting ocean-related data or possibly to the fact that human understanding of land environments is better than that of the ocean.
+This observation highlights the need to direct the model’s focus to oceanic domains.
+4.4 The Validity of Dysca
 Table 4: The correlation results on three benchmarks, where $\rho\in[-1,1]$ and $\tau\in[-1,1]$.
@@
 In this section, we investigate on the evaluation gap between Dysca and non-synthesis benchmarks.
-We calculate the Spearman’s rank correlation coefficient [63] $\rho$ and the Kendall rank correlation coefficient [64] $\tau$ between the evaluation ranking of Dysca under clean scenario with the non-synthesis benchmark’s evaluation ranking, i.e., MMBench [39], OCRBench [52] and SeedBench-2 [24].
+We calculate the Spearman’s rank correlation coefficient (Spearman, 1904) $\rho$ and the Kendall rank correlation coefficient (KENDALL, 1938) $\tau$ between the evaluation ranking of Dysca under clean scenario with the non-synthesis benchmark’s evaluation ranking, i.e., MMBench (Liu et al., 2023c), OCRBench (Liu et al., 2024) and SeedBench-2 (Li et al., 2023b).
 Both coefficient generate a score in the range of [-1,1], where 1 represents a perfect positive correlation, -1 represents a perfect negative correlation, and 0 represents no correlation.
+These coefficients are typical tools for measuring the correlation between variables in statistics.
+When the absolute value of either coefficient exceeds 0.6, it is considered to indicate a significant correlation  (Akoglu, 2018).
 Specifically, we intersect our Dysca with current benchmarks based on the perceptual subtasks, evaluation models and evaluation question types.
@@
 The results are shown in the first row of Tab.
-5.
-For the MMbench [39] and OCRBench [52], both metrics show the high correlation, with $\rho$ and $\tau$ higher than 0.6.
-However, the correlation for SeedBench-2 [24] is not as strong.
+4.
+For the MMbench (Liu et al., 2023c) and OCRBench (Liu et al., 2024), both metrics show the high correlation, with $\rho$ and $\tau$ higher than 0.6.
+However, the correlation for SeedBench-2 (Li et al., 2023b) is not as strong.
 Considering that SeedBench-2 only contains realistic images, we conduct additional experiments using the evaluation ranks on our realistic style images only.
 As shown in the second row of Tab.
-5, the correlation results of SeedBench-2 significantly improve (i.e., 0.46 vs.
+4, the correlation results of SeedBench-2 significantly improve (i.e., 0.46 vs.
 0.64 for $\rho$ and 0.43 vs.
@@
 The correlation with OCRBench also improves to 1, demonstrating the validity of using synthetic datasets for evaluation LVLMs.
-To further explore the the impact of image styles on evaluation results, we present the average scores across all subtasks for each of the 51 styles in Fig.
-5.
-We observe slight score differences across all styles.
-In the case of realistic styles such as "iPhone photo", all LVLMs perform better compared to other image styles.
-The LVLMs also exhibit better performance on unrealistic but common styles like "expressionist".
-However, for unrealistic and less common styles such as "gothic", all models show relatively poor performance.
-The results reveal that the gap between Dysca and non-synthesis benchmarks primarily stems from the more diverse range of image styles, making Dysca a more comprehensive benchmark for assessing the perception ability compared to previous benchmarks.
-Figure 5: Illustration of each model’s performance across 51 image styles, where the darker colors represent higher scores.
-The representative styles are colored with non-black font.
-Realistic styles are shown in red font, unrealistic but common styles are displayed in yellow font, and unrealistic and less common styles are represented in blue font.
+Figure 7: The KMMD distance between each benchmarks, with darker colors indicating larger distances.
+Besides, we calculate the data distribution distance between each benchmark to prove the low distance distribution between Dysca and non-synthesis benchmarks.
+We select CCBench (Liu et al., 2023c), COCO-Val  (Lin et al., 2014), MMVet (Yu et al., 2023), MMBench (Liu et al., 2023c), MME (Fu et al., 2023), MMStar (Chen et al., 2024), OCRBench (Liu et al., 2024) and ScienceQA (Lu et al., 2022).
+The reason why we choose these benchmark is that they have been widely used in evaluating LVLMs.
+We use Kernel Maximum Mean Discrepancy (KMMD) (Schölkopf et al., 2007) to measure the distribution distance.
+Specifically, we randomly sample 3,000 images from each benchmark (if the scale of the benchmark less than 3000, we use all that data) and utilize CLIP  (Radford et al., 2021) to encode these images.
+Then, we calculate the KMMD value using an RBF kernel between each pair.
+The results are shown in Fig.
+7, with darker colors indicating larger distances.
+Each row represents the value of KMMD between two benchmarks.
+The value of last row denotes the average value of KMMD.
+As can be seen, the distribution distance between Dysca and real-image benchmarks ranks in the middle compared to all other benchmarks.
+This demonstrates that the distribution between our benchmark and real benchmarks is similar, indicating that the evaluation results can effectively reflect the model’s performance in real-world scenarios.
 5 Conclusion
@@
 Dysca consists of 617K Vision-language QA pairs, covering 20 perceptual subtasks, 4 image scenarios and 3 question types.
-We conduct the experiment on 8 advanced open-source LVLMs with 10 checkpoints, revealing the insightful weakness of current LVLMs when facing different question types, image styles and image conditions.
+We conduct the experiment on 24 advanced open-source LVLMs and 2 closed-source LVLMs, revealing the insightful weakness of current LVLMs when facing different question types, image styles and image conditions.
 Experiments demonstrate the validity on evaluating LVLMs by using synthesis images.
 Acknowledgement
-This work is partially supported by National Key R&\&&D Program of China (No.
-2021YFC3310100), Strategic Priority Research Program of the Chinese Academy of Sciences (No.
-XDB0680000), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No.
+This work is partially supported by Strategic Priority Research Program of the Chinese Academy of Sciences (No.
+XDB0680202), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No.
 SYG202325), and Youth Innovation Promotion Association CAS.
@@
 Appendix A The Leaderboards
-Rank Model Score
-  1 XComposer-VL 80.13
-2 InstructBlip 72.36
-3 Blip2 70.03
-4 Shikra-VQA 64.03
-5 Shikra 62.97
-6 Otter 62.22
-7 Qwen-VL-Chat 61.13
-8 LLAVA-1.5-13B 60.17
-9 LLAVA-1.5-7B 53.81
-10 MiniGPT-4 42.52
-(a) Clean-Movie
-Rank Model Score
-  1 XComposer-VL 95.89
-2 InstructBlip 95.44
-3 Blip2 94.64
-4 Shikra-VQA 88.25
-5 Qwen-VL-Chat 79.75
-6 Shikra 77.65
-7 Otter 69.10
-8 LLAVA-1.5-13B 68.82
-9 LLAVA-1.5-7B 62.77
-10 MiniGPT-4 46.98
-(b) Clean-Action
-Rank Model Score
-  1 XComposer-VL 75.55
-2 InstructBlip 63.42
-3 Otter 63.04
-4 Shikra-VQA 60.10
-5 Blip2 59.14
-6 Shikra 58.55
-7 Qwen-VL-Chat 58.27
-8 LLAVA-1.5-13B 56.69
-9 LLAVA-1.5-7B 53.48
-10 MiniGPT-4 40.23
-(c) Clean-TV Show
-Rank Model Score
-  1 XComposer-VL 79.92
-2 InstructBlip 77.84
-3 Blip2 76.88
-4 Shikra 73.55
-5 Shikra-VQA 73.07
-6 Qwen-VL-Chat 65.12
-7 LLAVA-1.5-13B 62.86
-8 Otter 60.95
-9 LLAVA-1.5-7B 57.74
-10 MiniGPT-4 42.06
-(d) Clean-Profession
-Rank Model Score
-  1 InstructBlip 96.90
-2 Blip2 96.64
-3 XComposer-VL 96.43
-4 Shikra-VQA 82.85
-5 Shikra 80.17
-6 Qwen-VL-Chat 75.84
-7 LLAVA-1.5-13B 75.12
-8 LLAVA-1.5-7B 69.74
-9 Otter 62.50
-10 MiniGPT-4 52.81
-(e) Clean-Landmark
-Rank Model Score
-  1 XComposer-VL 75.00
-2 Qwen-VL-Chat 63.11
-3 InstructBlip 62.66
-4 Otter 59.56
-5 Blip2 59.50
-6 LLAVA-1.5-13B 53.88
-7 Shikra-VQA 53.31
-8 Shikra 52.84
-9 LLAVA-1.5-7B 47.80
-10 MiniGPT-4 38.98
-(f) Clean-Anime
-Rank Model Score
-  1 XComposer-VL 85.96
-2 InstructBlip 83.80
-3 Blip2 78.09
-4 Shikra-VQA 69.83
-5 Qwen-VL-Chat 69.58
-6 Shikra 67.43
-7 LLAVA-1.5-13B 61.20
-8 Otter 58.33
-9 LLAVA-1.5-7B 48.80
-10 MiniGPT-4 40.86
-(g) Clean-Clothes
-Rank Model Score
-  1 XComposer-VL 87.91
-2 InstructBlip 79.91
-3 Blip2 79.23
-4 Qwen-VL-Chat 68.71
-5 Shikra-VQA 62.93
-6 Shikra 61.81
-7 LLAVA-1.5-13B 59.04
-8 LLAVA-1.5-7B 56.04
-9 Otter 52.97
-10 MiniGPT-4 39.13
-(h) Clean-Celebrity
-Rank Model Score
-  1 XComposer-VL 91.27
-2 InstructBlip 90.28
-3 Blip2 89.78
-4 Shikra-VQA 80.23
-5 Shikra 75.84
-6 Qwen-VL-Chat 71.37
-7 LLAVA-1.5-13B 68.16
-8 Otter 63.22
-9 LLAVA-1.5-7B 52.87
-10 MiniGPT-4 47.72
-(i) Clean-Food
-Rank Model Score
-  1 XComposer-VL 93.19
-2 InstructBlip 92.62
-3 Blip2 91.84
-4 Shikra-VQA 80.05
-5 Shikra 77.57
-6 Qwen-VL-Chat 76.11
-7 Otter 73.96
-8 LLAVA-1.5-13B 67.61
-9 LLAVA-1.5-7B 48.84
-10 MiniGPT-4 47.95
-(j) Clean-Plant
-Rank Model Score
-  1 XComposer-VL 73.12
-2 InstructBlip 62.34
-3 Shikra-VQA 61.42
-4 Blip2 61.13
-5 Shikra 61.04
-6 LLAVA-1.5-13B 54.57
-7 Qwen-VL-Chat 51.28
-8 LLAVA-1.5-7B 46.75
-9 Otter 44.58
-10 MiniGPT-4 41.73
-(k) Clean-Age
-Rank Model Score
-  1 XComposer-VL 98.83
-2 Blip2 97.12
-3 InstructBlip 93.86
-4 LLAVA-1.5-13B 91.12
-5 Shikra-VQA 90.40
-6 Shikra 85.50
-7 Qwen-VL-Chat 78.13
-8 Otter 78.09
-9 MiniGPT-4 52.86
-10 LLAVA-1.5-7B 51.96
-(l) Clean-Gender
-Rank Model Score
-  1 XComposer-VL 86.08
-2 InstructBlip 85.59
-3 Shikra-VQA 83.53
-4 Blip2 82.59
-5 Shikra 80.51
-6 Qwen-VL-Chat 74.76
-7 Otter 67.11
-8 LLAVA-1.5-13B 64.81
-9 LLAVA-1.5-7B 56.81
-10 MiniGPT-4 46.46
-(m) Clean-Expression
-Rank Model Score
-  1 XComposer-VL 78.56
-2 InstructBlip 76.31
-3 Blip2 73.16
-4 LLAVA-1.5-13B 66.50
-5 Qwen-VL-Chat 65.19
-6 Shikra-VQA 64.56
-7 Shikra 64.53
-8 Otter 50.44
-9 LLAVA-1.5-7B 47.27
-10 MiniGPT-4 38.98
-(n) Clean-Race
-Rank Model Score
-  1 XComposer-VL 97.27
-2 InstructBlip 96.09
-3 Blip2 95.95
-4 Shikra-VQA 82.72
-5 Otter 82.47
-6 Qwen-VL-Chat 79.98
-7 Shikra 79.44
-8 LLAVA-1.5-13B 72.11
-9 LLAVA-1.5-7B 50.08
-10 MiniGPT-4 47.77
-(o) Clean-Animal
-Rank Model Score
-  1 XComposer-VL 90.97
-2 InstructBlip 90.00
-3 Blip2 88.53
-4 Shikra-VQA 82.72
-5 Qwen-VL-Chat 75.73
-6 Shikra 70.06
-7 LLAVA-1.5-13B 67.89
-8 Otter 64.91
-9 LLAVA-1.5-7B 54.31
-10 MiniGPT-4 51.65
-(p) Clean-Object
-Rank Model Score
-  1 Blip2 68.38
-2 InstructBlip 67.50
-3 XComposer-VL 66.06
-4 Shikra-VQA 64.86
-5 Otter 64.37
-6 Qwen-VL-Chat 63.28
-7 Shikra 63.02
-8 LLAVA-1.5-13B 58.84
-9 LLAVA-1.5-7B 44.96
-10 MiniGPT-4 43.70
-(q) Clean-OCR
-Rank Model Score
-  1 XComposer-VL 73.94
-2 Blip2 66.98
-3 InstructBlip 66.77
-4 Otter 63.82
-5 Qwen-VL-Chat 61.80
-6 Shikra 57.21
-7 Shikra-VQA 56.89
-8 LLAVA-1.5-13B 55.68
-9 LLAVA-1.5-7B 49.05
-10 MiniGPT-4 41.47
-(r) Clean-Style
-Rank Model Score
-  1 XComposer-VL 73.27
-2 Shikra-VQA 68.47
-3 InstructBlip 67.84
-4 Shikra 66.95
-5 Blip2 66.17
-6 Otter 63.23
-7 Qwen-VL-Chat 60.51
-8 LLAVA-1.5-13B 59.00
-9 LLAVA-1.5-7B 47.64
-10 MiniGPT-4 41.02
-(s) Clean-Background
-Rank Model Score
-  1 XComposer-VL 88.50
-2 InstructBlip 88.24
-3 Blip2 86.78
-4 Shikra-VQA 74.14
-5 Shikra 73.37
-6 Qwen-VL-Chat 69.72
-7 LLAVA-1.5-13B 58.36
-8 Otter 53.12
-9 LLAVA-1.5-7B 46.91
-10 MiniGPT-4 42.90
-(t) Clean-Color
-Table 6: Leaderboards of all Perceptual tasks under clean scenario.
-Rank Model Score
-  1 XComposer-VL 78.20
-2 InstructBlip 71.30
-3 Blip2 68.69
-4 Shikra-VQA 64.19
-5 Shikra 63.47
-6 Otter 61.39
-7 Qwen-VL 59.27
-8 LLAVA-1.5-13B 58.67
-9 LLAVA-1.5-7B 52.36
-10 MiniGPT-4 43.02
-(a) Corru.-Movie
-Rank Model Score
-  1 XComposer-VL 95.86
-2 InstructBlip 95.09
-3 Blip2 93.87
-4 Shikra-VQA 88.72
-5 Shikra 77.77
-6 Qwen-VL 75.09
-7 Otter 70.09
-8 LLAVA-1.5-13B 68.34
-9 LLAVA-1.5-7B 63.00
-10 MiniGPT-4 47.22
-(b) Corru.-Action
-Rank Model Score
-  1 XComposer-VL 72.86
-2 InstructBlip 62.42
-3 Otter 60.71
-4 Shikra 60.09
-5 Shikra-VQA 59.58
-6 Blip2 58.16
-7 Qwen-VL 57.95
-8 LLAVA-1.5-13B 55.18
-9 LLAVA-1.5-7B 52.34
-10 MiniGPT-4 38.70
-(c) Corru.-TV Show
-Rank Model Score
-  1 XComposer-VL 78.48
-2 InstructBlip 77.84
-3 Blip2 75.77
-4 Shikra 73.72
-5 Shikra-VQA 73.38
-6 Qwen-VL 64.97
-7 Otter 62.71
-8 LLAVA-1.5-13B 62.23
-9 LLAVA-1.5-7B 56.95
-10 MiniGPT-4 43.22
-(d) Corru.-Profession
-Rank Model Score
-  1 InstructBlip 96.33
-2 Blip2 96.28
-3 XComposer-VL 95.97
-4 Shikra-VQA 83.01
-5 Shikra 82.38
-6 LLAVA-1.5-13B 74.63
-7 Qwen-VL 73.55
-8 LLAVA-1.5-7B 68.78
-9 Otter 61.65
-10 MiniGPT-4 49.55
-(e) Corru.-Landmark
-Rank Model Score
-  1 XComposer-VL 72.28
-2 InstructBlip 62.79
-3 Qwen-VL 60.44
-4 Blip2 58.67
-5 Otter 57.24
-6 LLAVA-1.5-13B 53.67
-7 Shikra-VQA 52.28
-8 Shikra 52.00
-9 LLAVA-1.5-7B 47.87
-10 MiniGPT-4 40.91
-(f) Corru.-Anime
-Rank Model Score
-  1 XComposer-VL 84.87
-2 InstructBlip 82.84
-3 Blip2 76.69
-4 Shikra-VQA 70.59
-5 Shikra 67.05
-6 Qwen-VL 64.26
-7 LLAVA-1.5-13B 60.88
-8 Otter 59.87
-9 LLAVA-1.5-7B 48.12
-10 MiniGPT-4 41.38
-(g) Corru.-Clothes
-Rank Model Score
-  1 XComposer-VL 87.25
-2 InstructBlip 79.23
-3 Blip2 78.36
-4 Qwen-VL 64.89
-5 Shikra-VQA 62.55
-6 Shikra 61.68
-7 LLAVA-1.5-13B 58.64
-8 LLAVA-1.5-7B 56.36
-9 Otter 53.59
-10 MiniGPT-4 38.80
-(h) Corru.-Celebrity
-Rank Model Score
-  1 XComposer-VL 90.61
-2 InstructBlip 90.08
-3 Blip2 89.19
-4 Shikra-VQA 80.55
-5 Shikra 75.63
-6 Qwen-VL 70.71
-7 LLAVA-1.5-13B 67.46
-8 Otter 65.06
-9 LLAVA-1.5-7B 52.94
-10 MiniGPT-4 47.76
-(i) Corru.-Food
-Rank Model Score
-  1 XComposer-VL 92.45
-2 InstructBlip 92.23
-3 Blip2 91.15
-4 Shikra-VQA 80.80
-5 Shikra 77.69
-6 Otter 73.42
-7 Qwen-VL 72.01
-8 LLAVA-1.5-13B 67.22
-9 LLAVA-1.5-7B 48.84
-10 MiniGPT-4 48.04
-(j) Corru.-Plant
-Rank Model Score
-  1 XComposer-VL 73.04
-2 InstructBlip 61.59
-3 Shikra-VQA 61.09
-4 Shikra 60.40
-5 Blip2 60.03
-6 Qwen-VL 55.26
-7 LLAVA-1.5-13B 53.67
-8 LLAVA-1.5-7B 46.76
-9 Otter 45.09
-10 MiniGPT-4 40.84
-(k) Corru.-Age
-Rank Model Score
-  1 XComposer-VL 98.53
-2 Blip2 97.76
-3 InstructBlip 93.75
-4 LLAVA-1.5-13B 91.09
-5 Shikra-VQA 90.17
-6 Shikra 85.91
-7 Otter 80.22
-8 Qwen-VL 67.00
-9 MiniGPT-4 52.75
-10 LLAVA-1.5-7B 52.27
-(l) Corru.-Gender
-Rank Model Score
-  1 XComposer-VL 86.24
-2 InstructBlip 84.95
-3 Shikra-VQA 83.56
-4 Blip2 81.56
-5 Shikra 80.56
-6 Qwen-VL 70.55
-7 Otter 66.64
-8 LLAVA-1.5-13B 64.94
-9 LLAVA-1.5-7B 56.16
-10 MiniGPT-4 46.55
-(m) Corru.-Expression
-Rank Model Score
-  1 XComposer-VL 78.80
-2 InstructBlip 76.19
-3 Blip2 71.72
-4 LLAVA-1.5-13B 65.33
-5 Shikra-VQA 64.73
-6 Shikra 63.94
-7 Qwen-VL 63.55
-8 Otter 49.73
-9 LLAVA-1.5-7B 47.92
-10 MiniGPT-4 39.65
-(n) Corru.-Race
-Rank Model Score
-  1 XComposer-VL 97.06
-2 InstructBlip 95.86
-3 Blip2 95.69
-4 Shikra-VQA 83.03
-5 Otter 82.84
-6 Shikra 79.44
-7 LLAVA-1.5-13B 71.47
-8 Qwen-VL 71.43
-9 LLAVA-1.5-7B 50.67
-10 MiniGPT-4 48.66
-(o) Corru.-Animal
-Rank Model Score
-  1 XComposer-VL 90.81
-2 InstructBlip 89.76
-3 Blip2 88.05
-4 Shikra-VQA 82.76
-5 Shikra 71.49
-6 Qwen-VL 67.78
-7 LLAVA-1.5-13B 67.23
-8 Otter 66.18
-9 LLAVA-1.5-7B 54.09
-10 MiniGPT-4 51.02
-(p) Corru.-Object
-Rank Model Score
-  1 Blip2 94.06
-2 InstructBlip 90.54
-3 XComposer-VL 87.19
-4 Shikra-VQA 80.80
-5 Otter 78.70
-6 Shikra 77.17
-7 Qwen-VL 69.66
-8 LLAVA-1.5-13B 68.58
-9 LLAVA-1.5-7B 49.10
-10 MiniGPT-4 44.63
-(q) Corru.-OCR
-Rank Model Score
-  1 XComposer-VL 72.34
-2 InstructBlip 65.06
-3 Blip2 64.86
-4 Otter 61.14
-5 Qwen-VL 60.34
-6 Shikra 56.61
-7 Shikra-VQA 55.81
-8 LLAVA-1.5-13B 53.65
-9 LLAVA-1.5-7B 47.67
-10 MiniGPT-4 40.81
-(r) Courruption-Style
-Rank Model Score
-  1 XComposer-VL 72.53
-2 Shikra-VQA 68.25
-3 InstructBlip 67.48
-4 Shikra 67.19
-5 Blip2 64.69
-6 Otter 62.99
-7 Qwen-VL 60.50
-8 LLAVA-1.5-13B 58.50
-9 LLAVA-1.5-7B 47.12
-10 MiniGPT-4 40.53
-(s) Corru.-Background
-Rank Model Score
-  1 XComposer-VL 88.24
-2 InstructBlip 88.09
-3 Blip2 86.25
-4 Shikra-VQA 74.14
-5 Shikra 73.07
-6 Qwen-VL 70.78
-7 LLAVA-1.5-13B 57.70
-8 Otter 56.14
-9 LLAVA-1.5-7B 46.76
-10 MiniGPT-4 42.86
-(t) Corru.-Color
-Table 7: Leaderboards of all Perceptual tasks under corruption scenario.
-Rank Model Score
-  1 XComposer-VL 67.09
-2 Blip2 52.55
-3 LLAVA-1.5-13B 46.74
-4 LLAVA-1.5-7B 45.24
-5 Shikra 44.84
-6 Shikra-VQA 44.36
-7 Qwen-VL 44.30
-8 InstructBlip 43.06
-9 MiniGPT-4 40.26
-10 Otter 37.97
-(a) Pr.Att.-Movie
-Rank Model Score
-  1 XComposer-VL 86.24
-2 Blip2 79.13
-3 InstructBlip 75.94
-4 Shikra-VQA 71.36
-5 Qwen-VL 65.26
-6 Shikra 58.82
-7 LLAVA-1.5-13B 57.66
-8 LLAVA-1.5-7B 51.62
-9 MiniGPT-4 47.07
-10 Otter 39.77
-(b) Pr.Att.-Action
-Rank Model Score
-  1 XComposer-VL 55.08
-2 LLAVA-1.5-7B 43.55
-3 Qwen-VL 40.72
-4 LLAVA-1.5-13B 40.09
-5 Blip2 39.71
-6 Shikra 39.01
-7 MiniGPT-4 37.40
-8 Shikra-VQA 36.50
-9 InstructBlip 32.17
-10 Otter 31.48
-(c) Pr.Att.-TV Show
-Rank Model Score
-  1 XComposer-VL 67.14
-2 Blip2 59.62
-3 Shikra 55.70
-4 InstructBlip 54.52
-5 Shikra-VQA 54.25
-6 Qwen-VL 48.66
-7 LLAVA-1.5-13B 47.49
-8 LLAVA-1.5-7B 42.86
-9 MiniGPT-4 42.22
-10 Otter 38.12
-(d) Pr.Att.-Profession
-Rank Model Score
-  1 XComposer-VL 89.81
-2 Blip2 84.85
-3 InstructBlip 76.78
-4 Shikra-VQA 68.28
-5 LLAVA-1.5-13B 65.98
-6 Shikra 64.47
-7 Qwen-VL 64.29
-8 LLAVA-1.5-7B 56.39
-9 MiniGPT-4 49.58
-10 Otter 43.89
-(e) Pr.Att.-Landmark
-Rank Model Score
-  1 XComposer-VL 59.72
-2 Blip2 45.05
-3 Qwen-VL 44.95
-4 LLAVA-1.5-7B 42.23
-5 LLAVA-1.5-13B 40.95
-6 InstructBlip 38.26
-7 Shikra-VQA 38.25
-8 MiniGPT-4 37.27
-9 Shikra 37.02
-10 Otter 36.16
-(f) Pr.Att.-Anime
-Rank Model Score
-  1 XComposer-VL 76.43
-2 Blip2 66.53
-3 InstructBlip 62.14
-4 Qwen-VL 51.56
-5 LLAVA-1.5-13B 49.63
-6 Shikra-VQA 49.02
-7 Shikra 47.23
-8 LLAVA-1.5-7B 42.22
-9 MiniGPT-4 39.66
-10 Otter 34.67
-(g) Pr.Att.-Clothes
-Rank Model Score
-  1 XComposer-VL 73.48
-2 Blip2 57.30
-3 Qwen-VL 52.34
-4 LLAVA-1.5-7B 45.11
-5 InstructBlip 43.36
-6 Shikra 43.13
-7 Shikra-VQA 42.01
-8 LLAVA-1.5-13B 41.59
-9 MiniGPT-4 38.64
-10 Otter 33.95
-(h) Pr.Att.-Celebrity
-Rank Model Score
-  1 XComposer-VL 83.60
-2 Blip2 81.67
-3 InstructBlip 75.94
-4 Qwen-VL 60.43
-5 Shikra-VQA 58.61
-6 LLAVA-1.5-13B 56.20
-7 Shikra 56.08
-8 MiniGPT-4 46.70
-9 LLAVA-1.5-7B 45.20
-10 Otter 38.39
-(i) Pr.Att.-Food
-Rank Model Score
-  1 XComposer-VL 82.18
-2 Blip2 82.16
-3 InstructBlip 74.20
-4 Qwen-VL 60.29
-5 Shikra 54.03
-6 LLAVA-1.5-13B 53.77
-7 Shikra-VQA 52.86
-8 MiniGPT-4 46.45
-9 LLAVA-1.5-7B 42.07
-10 Otter 39.52
-(j) Pr.Att.-Plant
-Rank Model Score
-  1 XComposer-VL 60.37
-2 LLAVA-1.5-7B 42.53
-3 Blip2 42.20
-4 LLAVA-1.5-13B 41.30
-5 MiniGPT-4 39.18
-6 Shikra 38.98
-7 Qwen-VL 37.56
-8 Shikra-VQA 37.47
-9 InstructBlip 35.75
-10 Otter 32.18
-(k) Pr.Att.-Age
-Rank Model Score
-  1 Blip2 93.97
-2 InstructBlip 80.44
-3 XComposer-VL 79.56
-4 LLAVA-1.5-13B 74.03
-5 Shikra-VQA 63.62
-6 Shikra 62.06
-7 Qwen-VL 56.05
-8 MiniGPT-4 51.87
-9 LLAVA-1.5-7B 51.33
-10 Otter 39.31
-(l) Pr.Att.-Gender
-Rank Model Score
-  1 XComposer-VL 78.59
-2 Blip2 67.94
-3 InstructBlip 67.72
-4 Shikra-VQA 65.00
-5 Shikra 61.51
-6 Qwen-VL 59.16
-7 LLAVA-1.5-13B 58.33
-8 LLAVA-1.5-7B 53.00
-9 MiniGPT-4 45.37
-10 Otter 39.91
-(m) Pr.Att.-Expression
-Rank Model Score
-  1 XComposer-VL 62.21
-2 Blip2 52.98
-3 InstructBlip 48.42
-4 LLAVA-1.5-13B 47.80
-5 Qwen-VL 44.81
-6 Shikra-VQA 42.09
-7 LLAVA-1.5-7B 41.75
-8 Shikra 41.07
-9 MiniGPT-4 38.03
-10 Otter 28.77
-(n) Pr.Att.-Race
-Rank Model Score
-  1 Blip2 90.07
-2 XComposer-VL 87.80
-3 InstructBlip 82.43
-4 Qwen-VL 65.20
-5 LLAVA-1.5-13B 60.13
-6 Shikra-VQA 55.53
-7 Shikra 53.98
-8 MiniGPT-4 48.02
-9 Otter 46.92
-10 LLAVA-1.5-7B 44.73
-(o) Pr.Att.-Animal
-Rank Model Score
-  1 Blip2 83.55
-2 XComposer-VL 81.03
-3 InstructBlip 76.06
-4 Qwen-VL 61.06
-5 LLAVA-1.5-13B 56.84
-6 Shikra-VQA 55.02
-7 MiniGPT-4 49.61
-8 Shikra 46.96
-9 LLAVA-1.5-7B 44.26
-10 Otter 38.32
-(p) Pr.Att.-Object
-Rank Model Score
-  1 XComposer-VL 60.60
-2 Blip2 52.84
-3 LLAVA-1.5-13B 45.26
-4 Qwen-VL 43.55
-5 LLAVA-1.5-7B 42.23
-6 Shikra 41.06
-7 MiniGPT-4 40.32
-8 Shikra-VQA 40.12
-9 InstructBlip 40.03
-10 Otter 33.86
-(q) Pr.Att.-Style
-Rank Model Score
-  1 XComposer-VL 64.99
-2 Shikra-VQA 55.50
-3 Blip2 54.94
-4 Shikra 54.16
-5 LLAVA-1.5-13B 51.16
-6 InstructBlip 49.98
-7 Qwen-VL 49.17
-8 LLAVA-1.5-7B 44.34
-9 MiniGPT-4 39.23
-10 Otter 37.94
-(r) Pr.Att.-Background
-Rank Model Score
-  1 XComposer-VL 78.66
-2 Blip2 74.47
-3 InstructBlip 63.59
-4 Qwen-VL 58.40
-5 Shikra 51.75
-6 Shikra-VQA 51.10
-7 LLAVA-1.5-13B 49.48
-8 LLAVA-1.5-7B 43.40
-9 MiniGPT-4 41.70
-10 Otter 35.56
-(s) Pr.Att.-Color
-Table 8: Leaderboards of all Perceptual tasks under print attacking scenario.
-Rank Model Score
-  1 Qwen-VL 58.17
-2 Shikra 57.41
-3 Shikra-VQA 57.26
-4 LLAVA-1.5-13B 57.17
-5 Otter 56.61
-6 LLAVA-1.5-7B 51.91
-7 MiniGPT-4 36.95
-8 Blip2 33.33
-9 InstructBlip 31.97
-10 XComposer-VL 31.75
-(a) Ad.Att.-Movie
-Rank Model Score
-  1 Shikra-VQA 85.84
-2 Qwen-VL 79.11
-3 Shikra 77.17
-4 LLAVA-1.5-13B 66.37
-5 Otter 65.81
-6 LLAVA-1.5-7B 60.84
-7 MiniGPT-4 37.59
-8 Blip2 34.64
-9 XComposer-VL 33.88
-10 InstructBlip 32.85
-(b) Ad.Att.-Action
-Rank Model Score
-  1 Qwen-VL 55.79
-2 Shikra-VQA 53.57
-3 Otter 53.05
-4 LLAVA-1.5-13B 53.03
-5 Shikra 52.61
-6 LLAVA-1.5-7B 48.36
-7 MiniGPT-4 38.22
-8 XComposer-VL 34.45
-9 Blip2 32.04
-10 InstructBlip 30.93
-(c) Ad.Att.-TV Show
-Rank Model Score
-  1 Shikra-VQA 71.47
-2 Shikra 67.66
-3 Qwen-VL 64.80
-4 Otter 61.44
-5 LLAVA-1.5-13B 59.83
-6 LLAVA-1.5-7B 56.15
-7 MiniGPT-4 36.95
-8 Blip2 36.14
-9 InstructBlip 35.02
-10 XComposer-VL 30.86
-(d) Ad.Att.-Profession
-Rank Model Score
-  1 Shikra-VQA 78.40
-2 Shikra 77.85
-3 Qwen-VL 76.34
-4 LLAVA-1.5-13B 71.92
-5 LLAVA-1.5-7B 68.62
-6 Otter 56.87
-7 MiniGPT-4 35.73
-8 XComposer-VL 35.33
-9 Blip2 34.53
-10 InstructBlip 32.03
-(e) Ad.Att.-Landmark
-Rank Model Score
-  1 Qwen-VL 60.07
-2 Otter 51.86
-3 LLAVA-1.5-13B 51.81
-4 Shikra-VQA 48.45
-5 Shikra 47.75
-6 LLAVA-1.5-7B 47.32
-7 MiniGPT-4 38.38
-8 Blip2 37.95
-9 InstructBlip 34.29
-10 XComposer-VL 30.89
-(f) Ad.Att.-Anime
-Rank Model Score
-  1 Qwen-VL 69.92
-2 Shikra-VQA 68.32
-3 Shikra 65.73
-4 LLAVA-1.5-13B 58.83
-5 Otter 57.58
-6 LLAVA-1.5-7B 47.05
-7 MiniGPT-4 38.05
-8 InstructBlip 36.65
-9 XComposer-VL 36.41
-10 Blip2 36.18
-(g) Ad.Att.-Clothes
-Rank Model Score
-  1 Qwen-VL 64.77
-2 Shikra 60.02
-3 Shikra-VQA 59.70
-4 LLAVA-1.5-13B 55.53
-5 LLAVA-1.5-7B 54.95
-6 Otter 52.26
-7 MiniGPT-4 37.48
-8 Blip2 34.72
-9 InstructBlip 33.52
-10 XComposer-VL 33.09
-(h) Ad.Att.-Celebrity
-Rank Model Score
-  1 Shikra-VQA 75.26
-2 Qwen-VL 73.72
-3 Shikra 71.07
-4 LLAVA-1.5-13B 64.95
-5 Otter 64.39
-6 LLAVA-1.5-7B 52.30
-7 MiniGPT-4 36.94
-8 XComposer-VL 33.16
-9 Blip2 32.48
-10 InstructBlip 32.13
-(i) Ad.Att.-Food
-Rank Model Score
-  1 Qwen-VL 79.59
-2 Shikra-VQA 76.06
-3 Shikra 73.53
-4 Otter 67.16
-5 LLAVA-1.5-13B 64.44
-6 LLAVA-1.5-7B 47.41
-7 MiniGPT-4 37.74
-8 Blip2 35.84
-9 InstructBlip 34.02
-10 XComposer-VL 34.00
-(j) Ad.Att.-Plant
-Rank Model Score
-  1 Shikra 56.95
-2 Shikra-VQA 56.80
-3 LLAVA-1.5-13B 51.43
-4 Qwen-VL 48.11
-5 LLAVA-1.5-7B 45.89
-6 Otter 42.82
-7 MiniGPT-4 37.14
-8 Blip2 34.16
-9 InstructBlip 33.14
-10 XComposer-VL 30.78
-(k) Ad.Att.-Age
-Rank Model Score
-  1 LLAVA-1.5-13B 89.35
-2 Shikra-VQA 89.06
-3 Shikra 85.34
-4 Otter 82.25
-5 Qwen-VL 79.18
-6 LLAVA-1.5-7B 51.95
-7 MiniGPT-4 48.38
-8 Blip2 39.48
-9 InstructBlip 38.51
-10 XComposer-VL 38.09
-(l) Ad.Att.-Gender
-Rank Model Score
-  1 Shikra-VQA 82.04
-2 Shikra 79.86
-3 Qwen-VL 73.91
-4 LLAVA-1.5-13B 63.13
-5 Otter 60.44
-6 LLAVA-1.5-7B 55.70
-7 MiniGPT-4 41.07
-8 Blip2 35.95
-9 InstructBlip 33.52
-10 XComposer-VL 32.80
-(m) Ad.Att.-Expression
-Rank Model Score
-  1 LLAVA-1.5-13B 63.06
-2 Shikra-VQA 61.19
-3 Shikra 60.24
-4 Qwen-VL 59.50
-5 LLAVA-1.5-7B 45.97
-6 Otter 45.50
-7 MiniGPT-4 37.47
-8 Blip2 36.81
-9 InstructBlip 36.36
-10 XComposer-VL 34.77
-(n) Ad.Att.-Race
-Rank Model Score
-  1 Qwen-VL 81.78
-2 Shikra-VQA 80.44
-3 Otter 78.00
-4 Shikra 77.66
-5 LLAVA-1.5-13B 70.94
-6 LLAVA-1.5-7B 49.30
-7 MiniGPT-4 37.61
-8 Blip2 35.23
-9 XComposer-VL 34.38
-10 InstructBlip 34.01
-(o) Ad.Att.-Animal
-Rank Model Score
-  1 Shikra-VQA 79.05
-2 Qwen-VL 77.26
-3 Shikra 69.98
-4 LLAVA-1.5-13B 65.33
-5 Otter 62.69
-6 LLAVA-1.5-7B 52.62
-7 MiniGPT-4 37.84
-8 Blip2 37.62
-9 XComposer-VL 36.97
-10 InstructBlip 36.88
-(p) Ad.Att.-Object
-Rank Model Score
-  1 Shikra-VQA 79.12
-2 Shikra 77.62
-3 Qwen-VL 76.50
-4 Otter 74.91
-5 LLAVA-1.5-13B 66.43
-6 XComposer-VL 53.21
-7 InstructBlip 50.03
-8 Blip2 47.80
-9 LLAVA-1.5-7B 46.78
-10 MiniGPT-4 36.78
-(q) Ad.Att.-OCR
-Rank Model Score
-  1 Qwen-VL 56.80
-2 Otter 53.47
-3 Shikra 51.05
-4 Shikra-VQA 50.30
-5 LLAVA-1.5-13B 50.01
-6 LLAVA-1.5-7B 44.85
-7 MiniGPT-4 37.98
-8 XComposer-VL 37.20
-9 InstructBlip 36.20
-10 Blip2 35.87
-(r) Ad.Att.-Style
-Rank Model Score
-  1 Shikra-VQA 65.41
-2 Shikra 64.50
-3 Otter 60.30
-4 Qwen-VL 60.09
-5 LLAVA-1.5-13B 56.28
-6 LLAVA-1.5-7B 46.53
-7 MiniGPT-4 36.91
-8 Blip2 35.05
-9 XComposer-VL 34.71
-10 InstructBlip 34.45
-(s) Ad.Att.-Background
-Rank Model Score
-  1 Qwen-VL 69.63
-2 Shikra 69.57
-3 Shikra-VQA 69.05
-4 LLAVA-1.5-13B 54.28
-5 Otter 53.12
-6 LLAVA-1.5-7B 46.08
-7 MiniGPT-4 37.54
-8 XComposer-VL 37.23
-9 Blip2 36.83
-10 InstructBlip 35.81
-(t) Ad.Att.-Color
-Table 9: Leaderboards of all Perceptual tasks under adversarial attacking scenario.
-Rank Model Score
-  1 XComposer-VL 72.37
-2 InstructBlip 67.41
-3 Blip2 66.13
-4 Qwen-VL-Chat 62.68
-5 Shikra 62.52
-6 LLAVA-1.5-13B 59.28
-7 Shikra-VQA 58.64
-8 Otter 55.49
-9 LLAVA-1.5-7B 51.75
-10 MiniGPT-4 42.64
-(a) Clean
-Rank Model Score
-  1 XComposer-VL 71.63
-2 InstructBlip 67.62
-3 Blip2 65.49
-4 Shikra 62.37
-5 Qwen-VL-Chat 60.58
-6 LLAVA-1.5-13B 58.73
-7 Shikra-VQA 58.47
-8 Otter 55.42
-9 LLAVA-1.5-7B 46.44
-10 MiniGPT-4 42.87
-(b) Corru.
-Rank Model Score
-  1 XComposer-VL 63.75
-2 Blip2 56.20
-3 Qwen-VL-Chat 51.57
-4 InstructBlip 51.23
-5 LLAVA-1.5-13B 50.70
-6 Shikra 49.08
-7 LLAVA-1.5-7B 46.47
-8 Shikra-VQA 43.72
-9 MiniGPT-4 42.87
-10 Otter 47.10
-(c) Print Attack
-Rank Model Score
-  1 Qwen-VL-Chat 60.39
-2 Shikra 59.31
-3 LLAVA-1.5-13B 56.84
-4 Shikra-VQA 55.16
-5 Otter 51.94
-6 LLAVA-1.5-7B 50.07
-7 InstructBlip 33.13
-8 MiniGPT-4 32.95
-9 Blip2 32.65
-10 XComposer-VL 30.52
-(d) Adversarial Attack
-Table 10: Leaderboards of comprehensive performance in clean, corruption, print attack and adversarial attack scenario.
-The model performance leaderboards for each subtask under each scenario are shown in Tab.
+The model performance results for each subtask under clean scenario are shown in Tab.
+5, Tab.
 6, Tab.
-7, Tab.
-8, and Tab.
-9.
-We compute the average of the multi-choice and true-or-false evaluation results under each subtask.
+7, and Tab.
+8.
 Since the free-form question allows to assess the model’s perception abilities across multiple subtasks at the same time, the results of free-form are not taken into account.
-The overall performances on each scenario are displayed in Tab.
-10.
-We calculate the average of the scores of the three questioning types (i.e., multi-choice, true-or-false and free-form) as the final score.
+Table 5: Evaluation results on the 5 perceptual subtasks.
+The top two results on each subtask are bolded and underlined, respectively.
+“MC” and “TF” indicate the accuracy (%) of “Multi-choices” and “True-or-false”, respectively.
+Model LLM Visual Encoder Movie Action Tv Show Profession Landmark
+MC TF MC TF MC TF MC TF MC TF
+MiniGPT-4 Vicuna-7B EVA-CLIP ViT-G 29.53 51.03 45.06 50.67 30.43 47.09 35.85 46.88 52.37 51.15
+MiniGPT-4 Vicuna-13B EVA-CLIP ViT-G 49.76 49.20 68.99 50.16 38.24 39.47 51.85 49.02 68.67 51.69
+MiniGPT-4 LLaMA2-7B EVA-CLIP ViT-G 61.35 52.94 91.38 57.72 47.06 55.26 81.48 49.02 87.95 58.43
+MiniGPT-2 LLaMA2-7B EVA-CLIP ViT-G 72.95 50.27 95.31 62.22 73.53 52.63 79.63 62.75 87.95 50.56
+BLIP2 Flan-T5-XL EVA-CLIP ViT-G 72.45 67.72 97.16 93.02 57.97 57.00 81.60 75.45 98.11 93.44
+BLIP2 OPT-3B EVA-CLIP ViT-G 31.88 41.18 34.49 48.23 32.35 47.37 25.93 54.90 32.53 43.82
+BLIP2 OPT-7B EVA-CLIP ViT-G 19.90 42.78 24.05 47.91 23.53 47.37 18.52 50.98 24.39 39.33
+InstructBLIP Vicuna-7B EVA-CLIP ViT-G 83.57 56.68 97.88 71.38 88.24 55.26 81.48 58.82 97.59 71.91
+InstructBLIP Vicuna-13B EVA-CLIP ViT-G 83.57 59.36 98.79 64.63 88.24 63.16 77.78 60.78 97.59 66.29
+InstructBLIP Flan-T5-XL EVA-CLIP ViT-G 77.93 67.24 97.60 93.98 68.12 56.04 82.08 75.00 97.16 95.08
+InstructBLIP Flan-T5-XXL EVA-CLIP ViT-G 75.80 62.83 97.75 93.54 66.67 61.84 84.91 78.57 96.85 90.16
+LLava-1.5 Vicuna-7B CLIP ViT-L 55.71 53.54 74.10 50.89 53.14 57.49 57.08 58.93 83.28 55.08
+LLava-1.5 Vicuna-13B CLIP ViT-L 65.45 55.91 79.04 59.58 58.45 58.45 67.45 62.95 94.01 57.05
+Otter LLaMA-7B CLIP ViT-L 66.01 59.53 66.62 72.44 70.05 57.00 68.87 55.36 59.31 66.56
+Shikra LLaMA-7B CLIP ViT-L 68.34 61.26 78.44 78.01 60.87 57.97 82.08 68.30 88.96 70.82
+Xcomposer-VL InternLM-7B EVA-CLIP ViT-G 80.82 77.64 97.01 94.80 78.26 70.53 84.43 76.34 97.16 95.41
+Xcomposer2-VL InternLM2-7B CLIP ViT-L 91.79 88.77 98.18 95.34 85.29 84.21 88.89 90.20 98.80 100.00
+Qwen-VL-Chat Qwen-7B OpenClip ViT-bigG 71.08 49.61 95.96 63.52 68.12 42.51 80.66 51.34 95.90 55.41
+Emu2-Chat LLaMA-33B EVA2-CLIP-E 91.30 54.55 98.79 55.14 85.29 50.00 90.74 41.18 100.00 69.66
+GLM-4V GLM-4-9B-Chat EVA2-CLIP-E 93.72 91.98 99.55 98.39 97.06 89.47 88.89 94.12 100.00 100.00
+MiniCPM-V2.5 Llama3-Instruct 8B SigLIP SoViT-400m 93.24 89.84 98.18 92.77 94.12 89.47 90.74 88.24 100.00 98.88
+Yi-VL Yi-6B-Chat OpenClip ViT-H 89.37 82.89 97.88 85.85 88.24 89.47 94.44 74.51 97.59 93.26
+mPLUG-Owl-2 LLaMA2-7B CLIP ViT-L 90.34 76.47 96.82 85.37 82.35 81.58 87.04 82.35 98.80 86.52
+Phi-3-Vision Phi-3 CLIP ViT-L 87.44 70.05 97.13 90.03 85.29 65.79 88.89 84.31 100.00 96.63
+[…]
+Gemini-1.5-Pro / / 92.72 63.64 99.24 89.55 93.55 68.42 90.74 84.31 100.00 94.38
+Table 6: Evaluation results on the 5 perceptual subtasks.
+The top two results on each subtask are bolded and underlined, respectively.
+“MC” and “TF” indicate the accuracy (%) of “Multi-choices” and “True-or-false”, respectively.
+Model LLM Visual Encoder Anime Clothes Celebrity Food Plant
+MC TF MC TF MC TF MC TF MC TF
+MiniGPT-4 Vicuna-7B EVA-CLIP ViT-G 27.44 47.27 30.61 49.49 27.26 47.87 43.96 50.00 47.07 49.94
+MiniGPT-4 Vicuna-13B EVA-CLIP ViT-G 44.25 45.14 47.76 53.57 31.06 51.05 73.66 51.21 68.35 49.94
+MiniGPT-4 LLaMA2-7B EVA-CLIP ViT-G 40.71 53.47 66.94 51.43 54.49 54.46 81.61 52.90 84.97 58.14
+MiniGPT-2 LLaMA2-7B EVA-CLIP ViT-G 50.44 47.92 65.71 47.50 57.81 53.00 86.58 56.76 89.76 59.38
+BLIP2 Flan-T5-XL EVA-CLIP ViT-G 56.91 61.77 84.08 75.26 80.78 76.45 93.24 90.56 92.53 90.15
+BLIP2 OPT-3B EVA-CLIP ViT-G 26.55 45.14 28.57 45.00 31.89 48.70 32.42 48.31 35.11 48.82
+BLIP2 OPT-7B EVA-CLIP ViT-G 24.78 46.53 29.39 45.00 23.29 51.86 26.96 49.15 26.73 50.25
+InstructBLIP Vicuna-7B EVA-CLIP ViT-G 83.19 70.14 89.80 58.21 80.40 58.67 93.66 68.12 96.28 73.29
+InstructBLIP Vicuna-13B EVA-CLIP ViT-G 76.11 62.50 93.06 60.71 77.91 54.13 95.03 65.46 95.61 72.30
+InstructBLIP Flan-T5-XL EVA-CLIP ViT-G 60.77 63.65 88.27 82.72 83.11 75.75 93.96 91.36 92.98 91.36
+InstructBLIP Flan-T5-XXL EVA-CLIP ViT-G 62.80 67.42 87.24 87.32 82.11 82.93 94.28 91.91 94.01 93.90
+LLava-1.5 Vicuna-7B CLIP ViT-L 48.37 48.02 47.86 50.92 56.38 54.56 56.68 50.83 50.80 50.06
+LLava-1.5 Vicuna-13B CLIP ViT-L 58.54 49.34 65.41 59.00 61.04 57.14 82.51 56.15 79.59 55.91
+Otter LLaMA-7B CLIP ViT-L 62.80 54.99 47.24 71.37 44.15 63.76 44.44 79.46 69.29 80.56
+Shikra LLaMA-7B CLIP ViT-L 47.15 58.19 76.63 59.71 63.90 62.65 84.26 67.72 88.73 66.39
+Xcomposer-VL InternLM-7B EVA-CLIP ViT-G 74.80 74.01 86.84 87.22 88.23 87.25 93.72 90.96 92.34 91.55
+Xcomposer2-VL InternLM2-7B CLIP ViT-L 88.50 84.03 95.10 92.50 84.72 88.82 96.89 94.08 97.74 93.42
+Qwen-VL-Chat Qwen-7B OpenClip ViT-bigG 73.37 54.61 80.71 62.17 87.17 50.03 92.05 54.24 90.99 59.21
+Emu2-Chat LLaMA-33B EVA2-CLIP-E 86.73 66.67 94.69 49.64 94.02 47.16 97.76 49.15 98.54 54.66
+GLM-4V GLM-4-9B-Chat EVA2-CLIP-E 92.92 96.53 95.51 87.86 98.67 96.11 98.26 92.63 99.07 96.89
+MiniCPM-V2.5 Llama3-Instruct 8B SigLIP SoViT-400m 82.30 89.58 91.84 92.50 88.04 80.71 96.02 93.48 98.14 95.03
+Yi-VL Yi-6B-Chat OpenClip ViT-H 77.88 81.94 92.24 76.07 87.21 86.71 95.28 79.59 96.14 79.63
+mPLUG-Owl-2 LLaMA2-7B CLIP ViT-L 83.19 70.14 89.80 81.79 86.05 72.61 93.79 84.18 94.95 84.47
+Phi-3-Vision Phi-3 CLIP ViT-L 76.11 72.22 88.57 90.00 83.89 62.07 94.16 91.79 96.01 93.66
+[…]
+Gemini-1.5-Pro / / 91.82 76.39 97.55 93.57 94.85 91.57 98.01 92.75 97.87 93.54
+Table 7: Evaluation results on the 5 perceptual subtasks.
+The top two results on each subtask are bolded and underlined, respectively.
+“MC” and “TF” indicate the accuracy (%) of “Multi-choices” and “True-or-false”, respectively.
+Model LLM Visual Encoder Age Gender Expression Race Animal
+MC TF MC TF MC TF MC TF MC TF
+MiniGPT-4 Vicuna-7B EVA-CLIP ViT-G 30.02 48.28 52.00 46.75 42.29 49.82 25.10 48.42 45.54 49.39
+MiniGPT-4 Vicuna-13B EVA-CLIP ViT-G 29.32 50.06 58.30 48.60 59.51 53.58 33.08 51.51 60.39 52.37
+MiniGPT-4 LLaMA2-7B EVA-CLIP ViT-G 49.48 49.74 84.34 56.14 57.26 57.74 28.35 52.96 89.84 54.49
+MiniGPT-2 LLaMA2-7B EVA-CLIP ViT-G 41.62 52.84 86.13 61.60 65.69 63.85 51.34 53.56 92.53 60.20
+BLIP2 Flan-T5-XL EVA-CLIP ViT-G 62.72 58.71 99.32 94.12 88.41 75.30 75.76 71.87 96.62 95.59
+BLIP2 OPT-3B EVA-CLIP ViT-G 24.61 43.28 61.41 48.60 34.53 53.84 31.55 51.51 36.92 48.78
+BLIP2 OPT-7B EVA-CLIP ViT-G 29.84 42.76 53.13 47.81 30.93 53.78 26.55 49.82 28.40 49.43
+InstructBLIP Vicuna-7B EVA-CLIP ViT-G 70.16 57.36 99.66 74.00 83.13 67.23 84.16 67.55 97.46 69.17
+InstructBLIP Vicuna-13B EVA-CLIP ViT-G 67.28 63.44 99.33 65.49 86.28 74.38 81.23 56.45 97.01 71.45
+InstructBLIP Flan-T5-XL EVA-CLIP ViT-G 64.06 58.48 99.58 86.91 91.11 78.06 78.28 73.98 97.08 95.52
+InstructBLIP Flan-T5-XXL EVA-CLIP ViT-G 67.84 81.16 99.58 98.50 82.56 80.11 82.46 83.55 97.32 95.71
+LLava-1.5 Vicuna-7B CLIP ViT-L 37.74 55.40 53.62 49.89 63.49 51.34 42.55 51.34 49.45 52.18
+LLava-1.5 Vicuna-13B CLIP ViT-L 49.29 59.78 98.60 83.59 70.12 58.87 70.81 63.45 86.12 58.74
+Otter LLaMA-7B CLIP ViT-L 36.78 50.92 78.99 77.84 72.50 61.02 42.38 58.62 81.57 83.57
+Shikra LLaMA-7B CLIP ViT-L 65.07 54.96 98.14 73.29 90.08 70.52 75.20 54.19 90.20 70.32
+Xcomposer-VL InternLM-7B EVA-CLIP ViT-G 67.09 79.02 99.53 97.57 90.25 81.34 79.47 76.26 97.43 96.32
+Xcomposer2-VL InternLM2-7B CLIP ViT-L 86.65 91.73 99.11 99.03 96.74 94.41 88.38 89.02 98.06 89.56
+Qwen-VL-Chat Qwen-7B OpenClip ViT-bigG 53.19 48.71 97.50 59.18 85.55 64.04 74.43 55.95 95.45 63.89
+Emu2-Chat LLaMA-33B EVA2-CLIP-E 84.95 49.35 99.89 50.91 92.58 53.06 91.95 51.39 98.21 53.51
+GLM-4V GLM-4-9B-Chat EVA2-CLIP-E 86.13 93.80 100.00 100.00 98.09 94.28 94.51 94.69 99.40 99.02
+MiniCPM-V2.5 Llama3-Instruct 8B SigLIP SoViT-400m 77.23 68.86 99.55 99.76 95.05 92.07 81.61 88.66 98.65 94.62
+Yi-VL Yi-6B-Chat OpenClip ViT-H 71.86 86.69 98.88 98.66 91.11 88.30 84.55 74.31 97.31 83.85
+mPLUG-Owl-2 LLaMA2-7B CLIP ViT-L 73.43 81.91 98.21 96.60 88.30 84.27 80.84 84.44 95.52 83.85
+Phi-3-Vision Phi-3 CLIP ViT-L 86.65 70.03 98.99 99.51 96.40 81.66 91.19 89.51 96.86 88.42
+[…]
+Gemini-1.5-Pro / / 82.07 77.26 99.55 98.91 96.40 87.65 31.42 63.57 97.91 89.72
+Table 8: Evaluation results on the 5 perceptual subtasks.
+The top two results on each subtask are bolded and underlined, respectively.
+“MC” and “TF” indicate the accuracy (%) of “Multi-choices” and “True-or-false”, respectively.
+Model LLM Visual Encoder Object Text Style Background Color
+MC TF MC TF MC TF MC TF MC TF
+MiniGPT-4 Vicuna-7B EVA-CLIP ViT-G 52.54 50.66 29.68 50.25 35.68 18.89 31.20 48.98 35.55 48.86
+MiniGPT-4 Vicuna-13B EVA-CLIP ViT-G 61.27 51.89 36.55 51.79 56.00 44.16 48.15 50.99 54.25 50.75
+MiniGPT-4 LLaMA2-7B EVA-CLIP ViT-G 92.33 55.07 49.67 57.84 73.19 48.73 58.10 49.50 69.79 54.14
+MiniGPT-2 LLaMA2-7B EVA-CLIP ViT-G 96.67 60.97 43.60 57.54 79.35 48.22 56.19 50.20 69.02 59.79
+BLIP2 Flan-T5-XL EVA-CLIP ViT-G 90.26 89.86 74.33 62.32 83.87 35.45 76.23 72.02 88.58 86.64
+BLIP2 OPT-3B EVA-CLIP ViT-G 37.83 44.63 26.68 53.08 27.90 44.67 29.21 45.03 32.34 46.33
+BLIP2 OPT-7B EVA-CLIP ViT-G 26.83 48.33 26.25 51.98 29.35 47.21 25.21 46.42 23.00 47.46
+InstructBLIP Vicuna-7B EVA-CLIP ViT-G 97.67 67.32 78.74 62.60 99.28 47.21 85.82 62.23 94.21 68.08
+InstructBLIP Vicuna-13B EVA-CLIP ViT-G 98.17 71.10 81.34 57.04 99.82 48.73 90.16 56.46 96.33 59.13
+InstructBLIP Flan-T5-XL EVA-CLIP ViT-G 90.70 91.10 75.14 60.76 83.94 33.05 77.08 73.41 90.26 86.45
+InstructBLIP Flan-T5-XXL EVA-CLIP ViT-G 90.26 92.13 78.61 60.09 83.65 38.62 78.77 80.23 88.81 88.76
+LLava-1.5 Vicuna-7B CLIP ViT-L 58.36 48.52 50.36 50.73 38.44 19.45 46.62 51.57 42.03 52.71
+LLava-1.5 Vicuna-13B CLIP ViT-L 80.16 56.66 64.65 54.02 72.17 19.73 70.05 56.85 66.01 53.37
+Otter LLaMA-7B CLIP ViT-L 48.62 83.92 67.20 62.46 82.27 23.54 71.49 67.59 49.89 57.76
+Shikra LLaMA-7B CLIP ViT-L 74.35 70.26 64.99 53.78 79.65 22.97 79.58 64.28 83.06 60.57
+Xcomposer-VL InternLM-7B EVA-CLIP ViT-G 90.19 91.24 72.46 79.01 79.65 29.18 81.35 81.01 86.60 89.37
+Xcomposer2-VL InternLM2-7B CLIP ViT-L 98.83 97.13 89.26 88.79 100.00 94.42 92.06 92.94 94.11 95.95
+Qwen-VL-Chat Qwen-7B OpenClip ViT-bigG 90.12 65.15 73.82 52.18 82.63 21.07 78.73 51.57 86.37 53.25
+Emu2-Chat LLaMA-33B EVA2-CLIP-E 98.67 51.44 87.09 53.67 100.00 47.21 91.85 48.51 95.46 52.07
+GLM-4V GLM-4-9B-Chat EVA2-CLIP-E 99.00 98.49 93.60 87.20 100.00 94.92 94.39 95.33 94.69 94.54
+MiniCPM-V2.5 Llama3-Instruct 8B SigLIP SoViT-400m 98.50 96.97 88.39 84.13 100.00 98.98 90.69 88.87 96.33 95.10
+Yi-VL Yi-6B-Chat OpenClip ViT-H 98.17 75.04 84.49 75.60 100.00 84.26 90.05 75.94 94.79 92.09
+mPLUG-Owl-2 LLaMA2-7B CLIP ViT-L 96.33 88.80 81.45 65.77 99.28 78.17 85.82 74.65 91.80 88.23
+Phi-3-Vision Phi-3 CLIP ViT-L 98.17 97.73 32.32 28.87 100.00 95.94 90.14 78.42 88.63 89.75
+[…]
+Gemini-1.5-Pro / / 98.83 95.76 89.37 80.85 98.91 92.39 91.96 88.17 87.15 88.42
+It can be observed that certain models exhibit inconsistency when faced with different forms of questions.
+We have noted this phenomenon across several state-of-the-art models (e.g., Emu2 (Sun et al., 2023b) and Qwen-VL-Chat (Bai et al., 2023a)).
+We also provide a visualization of the scores for the top six models across 20 subtasks in different scenarios, as shown in Fig.
+8.
+It can be observed that the same model exhibits performance variations across different dimensions.
+For instance, age perception is the most significant weakness for the Yi-VL (AI et al., 2024).
+Additionally, by comparing the radar charts of different scenarios, we can see that corruption scenario has the least impact on the models, while print attacking is almost catastrophic.
+This highlights the need for future work to focus on improving model robustness against print attacking.
+((a)) Comparison of 6 top LVLMs on 20 subtasks under clean scenario. The full score of each subtask is 100.
+((b)) Comparison of 6 top LVLMs on 20 subtasks under corruption scenario. The full score of each subtask is 100.
+((c)) Comparison of 6 top LVLMs on 20 subtasks under adversarial attacking scenario. The full score of each subtask is 100.
+((d)) Comparison of 6 top LVLMs on 19 subtasks under print attacking scenario. The full score of each subtask is 100.
+Figure 8: Radar chart for the top 6 LVLMs in each scenario,
 Appendix B Discussion
@@
 Synthesis Data for Training / Fine-tuning.
-The use of synthetic data for model training / fine-tuning has been adopted in the field of Natural Language Processing (NLP) [65].
+The use of synthetic data for model training / fine-tuning has been adopted in the field of Natural Language Processing (NLP) (Meta, 2024).
 In this work, we do not explore the possibility of utilizing our benchmark for model training.
@@
 Nevertheless, considering that Dysca has the capability to synthesize high-resolution and unlimited amounts of annotated multimodal data, we believe that Dysca also holds potential as a training data synthesis tool for LVLMs.
-Reproducibility and Licence.
+Reproducibility and License.
 All the experiments are built on 8 * RTX 4090.
-All the data and the code for generation and evaluation are released in https://github.com/Benchmark-Dysca/Dysca.
-The licence of Dysca is "CreativeML Open RAIL++-M", which follows the licence set by the Stable Diffusion XL.
+All the data and the code for generation and evaluation are released at https://github.com/Robin-WZQ/Dysca.
+The license of Dysca is “CreativeML Open RAIL++-M”, which follows the license set by the Stable Diffusion XL.
 Ethical Concerns.
-Our Dysca leverages the Stable Diffusion XL [36] to generate images.
+Our Dysca leverages the Stable Diffusion XL (Podell et al., 2023) to generate images.
 In order to prevent the model generating unsafe images, e.g., NSFW and offensive images, lots of efforts have been made.
-First, we use the safety checker [66] to post filter the unsafe images.
+First, we use the safety checker (Rando et al., 2022) to post filter the unsafe images.
+The safety checker is a post-processor deployed by model developers (Podell et al., 2023) to prevent the generation of NSFW images.
 With the unsafe image that is recognized by the safety checker, the model’s output will be a blank image.
 Besides, we manually exclude the specific styles or the word that may trigger the unsafe images generation from the Metadata $M$.
-We believe that our Dysca involves fewer ethical concerns.
+To validate the safety of Dysca, we utilize the NudeNet detector (Bedapudi, 2022) with with a threshold of 0.8 to identify images that contain NSFW content.
+NudeNet is the widely used automatic tool for detecting NSFW images which is adopted by (Gandikota et al., 2023; Lu et al., 2024b).
+With 77847 images (clean scenario) in our Dysca, only 5 images are classified as NSFW (0.006%).
+We check the 5 images and find they are all false positive samples.
+This indicates that our Dysca contains a minimal amount of such unsafety images.
 B.2 The Stability of Dysca
-Figure 6: The tendency of model’s overall performance under clean scenario with different scale of evaluation data.
+Figure 9: The tendency of 10 model’s overall performance under clean scenario with different scale of evaluation data.
 In this section, we focus on examining the stability of Dysca.
 We partition Dysca into 11 different scales: 1%, 10%, 20%, 30%, 40%, 50%, 60%, 70%, 80%, 90% and 100%.
-We compute the evaluation scores using each of these data scales.
+We randomly select 10 models and compute their evaluation scores using each of these data scales.
 The score is calculated as the sum of scores obtained from multiple-choice, true-or-false and free-form questions.
 As can be seen in Fig.
-6, when the evaluation data scale is less than 30% of Dysca (i.e., less than 46.8K samples), the evaluation score show significant fluctuations.
+9, when the evaluation data scale is less than 30% of Dysca (i.e., less than 46.8K samples), the evaluation score show significant fluctuations.
 When the data scale exceeds 40%, we obtain the stable results, reflecting current scale of Dysca achieves the stable and reliable evaluation results.
 Although 40% evaluation scale of Dysca has achieved stable scores, Dysca aims to provide more than just stable rankings, but also draws on massive amounts of data to provide in-depth feedback across different image styles and perceptual subtasks.
-Appendix C The Metadata ($M$) of Dysca
+Appendix C The Correctness of Dysca
+Since our generative pipeline relies on text-to-image models, the quality of the Dysca dataset is inherently influenced by the generation performance of the model.
+To address this issue, we have employed extensive data cleaning techniques to ensure the correctness of the dataset.
+Furthermore, to validate this, we randomly sample 7.7K images (10% of our data) and manually select images that contain wrong content or potentially lead to incorrect responses.
+In the end, a total of 167 images (2.2%) are filtered.
+Compared to previous but wide spread benchmarks, the incorrect data ratio of Dysca is less (Northcutt et al., 2021).
+For example, the ImageNet dataset (Deng et al., 2009) has at least 2,916 errors in its validation set, with an error rate of 6%; the QuickDraw dataset (Ha & Eck, 2018) has at least 5 million errors, with an error rate of approximately 10%.
+Subsequently, we compute the scores with and without these incorrect images, and the results are presented below.
+Table 9: The scores with and without these incorrect images.
+Model Before After
+minigpt4-vicuna-7b 41.38 41.24
+minigpt4-vicuna-13b 50.17 50.78
+minigpt4-llama2 56.61 56.82
+minigpt-v2 58.46 58.43
+blip2-flan-t5-xl 65.30 65.32
+blip2-opt-3b 39.54 39.80
+blip2-opt-7b 39.55 39.58
+instructblip-vicuna-7b 67.54 67.89
+instructblip-vicuna-13b 64.89 65.02
+instructblip-flan-t5-xl 66.54 66.73
+instructblip-flan-t5-xxl 68.65 68.74
+llava-1.5-7b 51.27 51.44
+llava-1.5-13b 59.23 59.46
+otter 54.90 54.75
+shikra-7b 62.24 62.35
+internlm-xcomposer-vl-7b 71.40 71.55
+internlm-xcomposer2-vl-7b 79.13 79.23
+qwen-vl-chat 62.18 62.26
+emu2-chat 63.64 63.67
+glm-4v-9b 82.09 82.13
+minicpm-llama2-v2.5 78.75 79.02
+yi-vl 75.71 76.32
+mplug-owl2 74.09 74.43
+phi-3-vision 73.23 73.57
+[…]
+Gemini-1.5-Pro-Vision-latest 77.79 77.99
+As observed, the score difference is negligible, making the automatic evaluation result reliable.
+We believe that with the development of deep generative models, Dysca will serve as an evolving benchmark with improving quality.
+Appendix D Blind Experiment
+To demonstrate that Dysca has minimal data leakage issues, we conduct the "Blind" experiment where only textual questions are provided to LVLMs.
+We present the evaluation results of LLaVa-1.5-7B  (Liu et al., 2023b) on Dysca and the six benchmarks, including MMMU  (Yue et al., 2024), MMB  (Liu et al., 2023c), ScienceQA  (Lu et al., 2022), AI2D  (Kembhavi et al., 2016), Seed  (Li et al., 2023c) and MathVista  (Lu et al., 2024a).
+Table 10: Scores of LLaVa under the blind setting across 7 benchmarks.
+Model MMMU MMB ScienceQA AI2D SEED MathVista Dysca
+Random Choice 22.1 0 24.2 23.8 24.3 17.9 37.5
+LLaVA-1.5-7B 29.9 19.5 64.1 48.7 37.5 20.3 38.7
+As shown in Tab.
+10, LLAVA significantly outperforms random selection when only text questions are provided in the other 6 benchmarks.
+In contrast, Dysca achieves results closest to random selection, indicating that our work has a limited data leakage issue.
+Appendix E The Metadata ($M$) of Dysca
 Metadata ($M$) is the core of Dysca, which is randomly assembled from our collected source material and contains all the information needed to generate prompt ($P$), image ($I$), and question-answer pairs($Q$).
@@
 On the one hand, some of the content is polysemous, which can easily be misinterpreted by the model’s when generating images.
-On the other hand, there are backgrounds or artistic styles (e.g., “Pokemon Style", “architectural style", etc.) that negatively affect the quality of the image and do not accurately generate the desired content.
+On the other hand, there are backgrounds or artistic styles (e.g., “Pokemon Style”, “architectural style”, etc.) that negatively affect the quality of the image and do not accurately generate the desired content.
 In order to test the usability of these source materials, we went through several small-scale pre-generations covering all the source materials.
@@
 Clothes We selected 16 common types of clothing that are highly distinguishable from each other. 16
-Object We took the annotations from the MSCOCO [50] dataset after removing people, animals, and plants, and added some additional common objects. 80
+Object We took the annotations from the MSCOCO (Lin et al., 2014) dataset after removing people, animals, and plants, and added some additional common objects. 80
 Landmark We chose 23 characteristic landmarks from around the globe and they can be generated by the Stable Diffusion model accurately. 23
@@
 Color We selected 8 easily distinguishable colors: red, orange, yellow, green, blue, purple, white, black. 8
-Appendix D Scenarios Details
-D.1 Print Attack Scenario
-Followed by the settings in [35], we add the attack text on the images.
-Consider that the image resolution in our Dysca is much higher than the one in [35], we increase more font form in terms of font position and font orientation.
+Appendix F Scenarios Details
+F.1 Print Attack Scenario
+Followed by the settings in (Cheng et al., 2024), we add the attack text on the images.
+Consider that the image resolution in our Dysca is much higher than the one in (Cheng et al., 2024), we increase more font form in terms of font position and font orientation.
 Fig.
@@
 ’R1C1’ means the typo is located in row 1, column 1, which is the top left corner of the image.
-D.2 Corrupted Scenario
+F.2 Corruption Scenario
 Examples of the 11 image corruptions are shown in Fig 15.
 Figure 15: Examples of 11 image corruptions applied to a single clean image.
-Appendix E More Examples of Dysca
+Appendix G Hard Samples
+In this section, we provide hard samples that LVLMs are likely to answer incorrectly, as shown in Fig. 16 to Fig. 18.
+Figure 16: Hard samples.
+Figure 17: Hard samples.
+Figure 18: Hard samples.
+Appendix H More Examples of Dysca
 For each subject we collected in Metadata ($M$), we display one example of their prompt ($P$), generated image ($I$) and corresponding question-answer pairs ($Q$).
@@
 Figure 32: OCR
-Appendix F Data Sheet
-We follow the documentation frameworks provided by Gebru et al. [67].
-F.1 Motivation
+Appendix I JSON structure for evaluation
+Here, We present two example JSON structures generated by Dysca for evaluation in Listing 1.
+Listing 1: Example JSON structure for evaluation
+"id": 1,
+"images": [
+"images/1.png"
+"prompt": "face portrait of a 25 years old female, Indian race, calm and placid, on Mars, retro arcade style, 8-bit, pixelated, vibrant, classic video game, old school gaming, reminiscent of 80s and 90s arcade games",
+"negative_prompt": "modern, ultra-high resolution, photorealistic, 3D",
+[…]
+"question": "What is the background in the picture?",
+"answer": "on Mars",
+"question_type": "multi choice",
+"task": "recognition",
+"options": [
+"on the lawn",
+"under the water background",
+"sunset background",
+"on Mars"
+"question_majority": "background",
+"granularity": "coarse"
+"id": 2,
+"images": [
+"images/2.png"
+"prompt": "1 white color cat, on Mars, thick layered papercut art of, deep 3D, volumetric, dimensional, depth, thick paper, high stack, heavy texture, tangible layers",
+"negative_prompt": "2D, flat, thin paper, low stack, smooth texture, painting, drawing, photo, deformed",
+[…]
+"question": "Is there a cat in the picture?",
+"answer": "True",
+"question_type": "true or false",
+"task": "recognition",
+"options": [
+"False",
+"True"
+"question_majority": "animal",
+"granularity": "fine"
+Appendix J Data Sheet
+We follow the documentation frameworks provided by Gebru et al. Gebru et al. (2021).
+J.1 Motivation
 For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
@@
 Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grant or and the grant name and number.
-  • It is funded by the Key Laboratory of AI Safety of CAS, Institute of Computing Technology.
-F.2 Composition
+  • This work is partially supported by Strategic Priority Research Program of the Chinese Academy of Sciences (No.
+XDB0680202), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No.
+SYG202325), and Youth Innovation Promotion Association CAS.
+J.2 Composition
 What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?
@@
   • No, our benchmark does not contain any sensitive data.
-F.3 Collection Process
+J.3 Collection Process
 How was the data associated with each instance acquired?
@@
   • No.
-F.4 Preprocessing/cleaning/labeling
+J.4 Preprocessing/cleaning/labeling
 Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
@@
   • Yes.
-We leverage the off-the-shelf models, i.e., PP-OCRv3 [56] and CLIP-L-14 [16], to clean the data.
-PP-OCRv3 [56] is leveraged as the filter to exclude the failure image that TextDiffusion2 [58] generates the wrong text on the image.
-For the other images, we use CLIP-L-14 [16] to filter out the images with low text-image consistency.
+We leverage the off-the-shelf models, i.e., PP-OCRv3 Li et al.
+(2022) and CLIP-L-14 Radford et al.
+(2021), to clean the data.
+PP-OCRv3 Li et al.
+(2022) is leveraged as the filter to exclude the failure image that TextDiffusion2 Chen et al.
+(2023a) generates the wrong text on the image.
+For the other images, we use CLIP-L-14 Radford et al.
+(2021) to filter out the images with low text-image consistency.
 Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
@@
 Is the software that was used to preprocess/clean/label the data available? If so, please provide a link or other access point.
-F.5 Uses
+J.5 Uses
 Has the dataset been used for any tasks already? If so, please provide a description.
@@
   • The proposed dataset should not be used to generate offensive data.
-F.6 Distribution
+J.6 Distribution
 Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
@@
   • Not yet.
-F.7 Maintenance
+J.7 Maintenance
 Who will be supporting/hosting/maintaining the dataset?

### v4

@@
 1, we provide an end-to-end process from image to Vision-QA generation.
-The approach significantly reduces annotation costs compared to manually labeling images (e.g., MME (Fu et al., 2023)) while achieving the correctness for evaluating LVLMs.
+This approach significantly reduces annotation costs compared to manually labeling images (e.g., MME (Fu et al., 2023)) while achieving the correctness for evaluating LVLMs.
 It also avoids the risk of hallucinate annotations that may occur when using ChatGPT for labeling based on image prompts (e.g., JourneyDB (Sun et al., 2023a)).
@@
 At the forefront of these developments, BLIP-2 (Li et al., 2023d) introduces a lightweight Q-Former (Li et al., 2023d) that facilitates alignment between textual and visual representations through a cross-attention mechanism (Li et al., 2023d).
-InstructBLIP (Dai et al., 2023) takes a step further by incorporating textual instructions into the Q-Former, which significantly improves performance.
-LLAVA (Liu et al., 2023b) employs GPT-4 (OpenAI, 2023) to transform data into multimodal instruction-following data and uses CLIP (Radford et al., 2021) and LLAMA (Touvron et al., 2023) for fine-tuning instructions, achieving advanced performance.
+InstructBLIP (Dai et al., 2023) takes a step further by incorporating textual instructions into the Q-Former, which significantly improves the zero-shot performance.
+LLAVA (Liu et al., 2023b) employs GPT-4 (OpenAI, 2023) to transform data into multimodal instruction-following data and uses CLIP (Radford et al., 2021) and LLAMA (Touvron et al., 2023) for fine-tuning instructions, achieving advanced multimodal chat abilities.
 LLAVA-1.5 (Liu et al., 2023a) extends this paradigm by integrating MLP projection and introducing academic task-specific Vision-language QA data.
@@
 These benchmarks evaluate models through the objective answer types and most of them leverage the automatic annotation and evaluation manner for revealing the fine-grained drawbacks of current LVLMs.
-However, the previous benchmarks primarily concentrate on evaluating LVLMs using realistic images and clean scenario, leaving multi-stylized images and noisy scenarios unexplored.
+However, the previous benchmarks primarily concentrate on evaluating LVLMs using realistic images under clean scenario, leaving multi-stylized images and noisy scenarios unexplored.
 Moreover, many of them conduct QA by selecting images from publicly available datasets (e.g., (Lin et al., 2014; Russakovsky et al., 2014)).
@@
 One possible way to solve data leakage is using novel but synthesis images, where JourneyDB (Sun et al., 2023a) is the first work aiming to leverage synthesis images to evaluate current LVLMs.
-The prompts and the corresponding images are downloaded from Midjourney (mid, ) and ChatGPT (OpenAI, 2022) is leveraged to label the images.
+The prompts together with the corresponding images are downloaded from Midjourney (mid, ) and ChatGPT (OpenAI, 2022) is leveraged to label the images.
 However, JourneyDB is a top-down framework where the number of images is fixed.
 Besides, the ChatGPT labeling may cause hallucinate annotations, leading to the unreliable evaluation results.
-In contrast, our Dysca serves as the bottom-up framework, allowing for dynamic and scalable generation for both images and evaluation questions and involving more evaluation tasks.
+In contrast, Dysca serves as a bottom-up framework which enables dynamic and scalable generation of both images and evaluation questions, while also supporting varied evaluation tasks.
 The rule-based question generation method also makes the annotations more accuracy.
@@
 Perception is one of the most fundamental capabilities of LVLMs and previous works (Fu et al., 2023) have shown that the lack of perceptual ability may result in hallucination (Li et al., 2023e).
-In order to comprehensively evaluate LVLMs’ perception capability, we collect and organize existing sub-dimensions from current evaluation datasets, resulting in a total of 20 assessment dimensions where we show all the subtasks and the corresponding amount of their annotation in the Fig.
+In order to comprehensively evaluate LVLMs’ perception capability, we collect and organize existing sub-dimensions from current benchmarks, resulting in 20 assessment dimensions where we show all the subtasks and the corresponding amount of their annotation in the Fig.
 2.
@@
 The metadata M is the core of the whole Dysca, containing all the information for generating $P$, $I$ and $Q$.
-The metadata $M$ consists of foreground, attribute, background and style, and these information guide the generation of the prompt ($P$) through pre-designed templates.
+The metadata $M$ consists of foreground, attribute, background and style, which guides the generation of the prompt ($P$) through pre-designed templates.
 Then, we utilize the T2I diffusion model to generate the corresponding image using the prompt $P$.
-For generating the image with specific text on it for the OCR subtask, we leverage TextDiffusion2 (Chen et al., 2023a), which is the state-of-the-art text rendering method.
+For generating the image with a specific text on it for the OCR subtask, we leverage TextDiffusion2 (Chen et al., 2023a), which is the state-of-the-art text rendering method.
 For the rest of images, we leverage Stable Diffusion XL (Podell et al., 2023).
@@
 Besides, for the free-form question type, since it allows to assess the model’s perception abilities across multiple subtasks at the same time, we reduce the number of free-form questions for achieving a balanced data distribution.
-Figure 3: The process of generating the prompt (P), image (I) and question-answer pairs (Q) from the metadata (M).
+Figure 3: The process of generating the prompt (P), image (I) and QA pairs (Q) from metadata (M).
 3.4 Evaluation Strategy
@@
 Figure 4: The failure cases for the noisy scenarios. From left to right are: corruption scenario, adversarial attacking scenario, and print attacking scenario.
+Figure 5: Models exhibit different performance when facing the same image but different question types.
 […]
 Besides, thanks to generative evaluation framework of Dysca, we are able to effectively control variables and conduct a detailed analysis of the model’s fine-grained capabilities.
-Specifically, we introduce two novel metrics to measure the sensitivity of LVLMs on question types and covariate shift.
+Specifically, we introduce two novel metrics to measure the sensitivity of LVLMs on question types and covariate shift (i.e., image style).
 The sensitivity to question type aims to evaluate whether LVLMs exhibit inconsistent performance when facing different types of question types (i.e., multi-choice vs.
@@
 where $SQ_{Avg}=\frac{TFSQ+MCSQ}{2}$.
-The sensitivity to covariate shift aims to evaluate whether LVLMs exhibit inconsistent performance when facing the same content and question format, but with variations in image covariates (i.e., image style).
+The sensitivity to covariate shift aims to evaluate whether LVLMs exhibit inconsistent performance when facing the same content and question format, but with variations in image covariates.
 It is defined by:
 $SC=\frac{\sum_{i=1}^{N}(S_{i}-SC_{Avg})^{2}}{N},$ (2)
-where $SC_{Avg}=\frac{\sum_{i=1}^{N}S}{N}$ and $S_{i}$ denotes the score of the models in each style. Since there are 51 image styles in Dysca, we set $N=51$.
+where $SC_{Avg}=\frac{\sum_{i=1}^{N}S_{i}}{N}$ and $S_{i}$ denotes the score of the models in each style. Since there are 51 image styles in Dysca, we set $N=51$.
 4 Results and Analysis
@@
 (1) For LVLMs, the capacity of the language model plays a crucial role.
-When using the same visual encoder, models that adopt a language model with larger parameter sizes (e.g., Vicuna-13B models generally outperform Vicuna-7B models by 8%) or models with a different but stronger architecture (e.g., GLM-4V (GLM et al., 2024) with GLM-4-9B-Chat shows a 20% performance increase compared to Emu2 (Sun et al., 2024) with LLaMA-33B) tend to achieve better performance.
+When using the same visual encoder, models that utilize a language model with a larger parameter sizes (e.g., MiniGPT-4 achieves an improvement score of 8.79 in the clean scenario when using Vicuna-13B compared to Vicuna-7B) or with a stronger capability (e.g., GLM-4V that uses the GLM-4-9B-Chat language model shows an improvement score of 18.45 in the clean scenario compared to Emu2, which uses the LLaMA-33B language model) tend to achieve a better performance.
 (2) Each model shows robustness in the corruption scenario, but experiences significant degradation in both attack scenarios.
-In the image corruption scenario, all models demonstrate minimal score variations (less than 1%), indicating consistent performance under non-targeted disruptions.
+In the image corruption scenario, all models demonstrate minimal score variations, i.e., less than 1%.
 However, under print attacks, performance drops are notable.
-For example, two closed-source models experience substantial declines: Gemini-1.5-pro (Team et al., 2024) drops by 21.5% (from 77.79 to 61.05), and GPT-4o (OpenAI, 2023) suffers a 25.8% decrease (from 75.69 to 56.10).
-Among the advanced open-source models, GLM-4V’s (GLM et al., 2024) performance decreases sharply by 36.5% (from 82.09 to 52.09), and Phi-3-Vision (Abdin et al., 2024) records a 21.39% drop (from 73.23 to 57.78).
-Notably, mPLUG-Owl-2 (Ye et al., 2023) shows the highest robustness, with only a 1.7% reduction (from 74.09 to 72.84), while the XComposer-VL (Zhang et al., 2023; Dong et al., 2024) series also exhibits strong resistance to print attacks.
-In the adversarial attack scenario, where the attack algorithm directly targets the image encoder, LVLMs with shared encoder architecture (e.g., Blip2 (Li et al., 2023d), InstructBLIP (Dai et al., 2023), and XComposer-VL (Zhang et al., 2023), all using EVA-CLIP (Fang et al., 2022) as their image encoder) suffer significant performance drops, with some models even performing worse than random chance.
-For instance, XComposer-VL (Zhang et al., 2023) experiences a 57.6% drop (from 71.40 to 30.28).
-Models with alternative image encoders also experience degradation, ranging from 1% to 5%.
-The greater impact of adversarial noise compared to corruption noise suggests that adversarial attacks exhibit a certain degree of transferability across different model architectures.
+For instance, two closed-source models exhibit significant performance degradation: Gemini-1.5-pro declines by 21.5%, resulting in a score reduction from 77.79 to 61.05, while GPT-4o shows a 25.8% decrease, dropping its score from 75.69 to 56.10.
+Among the leading open-source models, GLM-4V experiences a sharp 36.5% performance drop, lowering its score from 82.09 to 52.09, and Phi-3-Vision records a 21.39% decline, reducing its score from 73.23 to 57.78.
+Notably, mPLUG-Owl-2 demonstrates the highest robustness, with only a 1.7% reduction.
+The XComposer-VL series also exhibits strong resilience against print-based attacks.
+In the adversarial attack scenario, where the attack algorithm directly targets the image encoder, LVLMs employing a shared encoder architecture (e.g., Blip2, InstructBLIP, and XComposer-VL, all of which utilize EVA-CLIP (Fang et al., 2022) as their image encoder) suffer substantial performance declines, with some models performing even worse than random chance.
+For example, XComposer-VL experiences a 57.6% drop, reducing its score from 71.40 to 30.28.
+Models with different image encoders also experience performance degradation ranging from 1% to 5%, showing a higher impact than that of corruption noise.
 More detailed results are available in Appendix F.
-Figure 5: The Visualization of hierarchical clustering results of 20 subtasks.
-Figure 6: The Visualization of hierarchical clustering results of 51 animal categories.
 (3) Models exhibit varying sensitivity to different question types and covariate shifts.
-For the Sensitivity to Question Types, Xcomposer2-VL (Dong et al., 2024) achieves the best result with a score of 20.9.
+For the sensitivity to question types, we present two examples in Fig.
+5.
+As can be seen, the XComposer-VL (Zhang et al., 2023) recognizes the sparrow in the image under a multiple-choice setting but fails to identify the sparrow in the same image under a true-or-false setting.
+The quantity results are shown in Tab.
+3.
+Xcomposer2-VL achieves the best result with a score of 20.9.
 However, we observe that the perception ability of LVLM does not show a positive correlation with sensitivity to question types.
-For instance, while the GLM-4V (GLM et al., 2024) achieves the highest performance in evaluation tasks, it exhibits higher sensitivity to question types than Xcomposer2-VL (Dong et al., 2024).
+For instance, while the GLM-4V achieves the highest performance in evaluation tasks, it exhibits higher sensitivity to question types than Xcomposer2-VL.
 One of the factors influencing the sensitivity to question types may be the inherent biases of the language model.
 Using the same base language model may result in similar outcomes for this metric.
-it is also noted that Gemini (Team et al., 2024) performs bad in this metric, revealing its preference for certain question types.
-For the Sensitivity to Covariate Shifts, GLM-4V (Dong et al., 2024) achieves the best result with a score of 0.8.
+it is also noted that Gemini-1.5-pro performs bad in this metric, revealing its preference for certain question types.
+For the sensitivity to covariate shifts, as shown in the last column of Tab.
+3, GLM-4V achieves the best result with a score of 0.8.
 However, we also observe that the perception ability of LVLM does not show a positive correlation with covariate shifts.
-For example, InstructBLIP-Flan-T5-XXL (Dai et al., 2023) outperforms InstructBLIP-Flan-T5-XL in terms of performance but shows higher sensitivity to covariate shifts.
+For example, InstructBLIP-Flan-T5-XXL outperforms InstructBLIP-Flan-T5-XL in terms of performance but shows higher sensitivity to covariate shifts.
 4.3 Analysis on inter-task and intra-task
 Analysis on Inter-task.
-Here, we perform hierarchical clustering based on the Euclidean distance of the scores across 20 evaluation dimensions for 26 models.
-As shown in the Fig.
-6, based on the consistency of model performance across the 20 evaluation dimensions, the models can be categorized into four tiers.
-We also observe that LVLMs tend to perform better in dimensions that involve well-defined image perception, such as landmark recognition and object recognition.
+In order to investigate the inter-relationships across evaluation dimensions, we conduct hierarchical clustering based on the Euclidean distance of the scores across 20 dimensions for 26 models.
+We observe that models show varied consistency of performance across the dimensions.
+LVLMs tend to perform better in dimensions that involve well-defined image perception, such as landmark recognition and object recognition.
 However, dimensions likes style recognition and movie recognition exhibit greater challenges for LVLMs, which may be attributed to the limited training resources in these specific domains.
@@
 Taking animal categories as an example, we analyze the performance of LVLMs on different animals.
-Here, we perform same hierarchical clustering methods across 51 animal categories.
-As shown in the Fig.
-6, LVLMs performance vary significantly across these categories.
-Models tend to perform poorly when applied to marine life.
-This could be caused by the challenges in collecting ocean-related data or possibly to the fact that human understanding of land environments is better than that of the ocean.
+Here, we perform the same hierarchical clustering methods across 51 animal categories.
+LVLMs performance vary significantly across these categories.
+We find that models tend to perform poorly when applied to marine life.
+This could be caused by the challenges in collecting ocean-related data.
 This observation highlights the need to direct the model’s focus to oceanic domains.
@@
 4.
-For the MMbench (Liu et al., 2023c) and OCRBench (Liu et al., 2024), both metrics show the high correlation, with $\rho$ and $\tau$ higher than 0.6.
-However, the correlation for SeedBench-2 (Li et al., 2023b) is not as strong.
+For the MMbench and OCRBench, both metrics show the high correlation, with $\rho$ and $\tau$ higher than 0.6.
+However, the correlation for SeedBench-2 is not as strong.
 Considering that SeedBench-2 only contains realistic images, we conduct additional experiments using the evaluation ranks on our realistic style images only.
@@
 The results are shown in Fig.
-7, with darker colors indicating larger distances.
+6.
 Each row represents the value of KMMD between two benchmarks.
 The value of last row denotes the average value of KMMD.
-As can be seen, the distribution distance between Dysca and real-image benchmarks ranks in the middle compared to all other benchmarks.
-This demonstrates that the distribution between our benchmark and real benchmarks is similar, indicating that the evaluation results can effectively reflect the model’s performance in real-world scenarios.
+As can be seen, the distribution distance between Dysca and real-image benchmarks ranks in the middle compared to all other benchmarks, indicating that the evaluation results can effectively reflect the model’s performance in real-world scenarios.
 5 Conclusion
@@
 We also provide a visualization of the scores for the top six models across 20 subtasks in different scenarios, as shown in Fig.
-8.
+7.
 It can be observed that the same model exhibits performance variations across different dimensions.
@@
 As can be seen in Fig.
-9, when the evaluation data scale is less than 30% of Dysca (i.e., less than 46.8K samples), the evaluation score show significant fluctuations.
+8, when the evaluation data scale is less than 30% of Dysca (i.e., less than 46.8K samples), the evaluation score show significant fluctuations.
 When the data scale exceeds 40%, we obtain the stable results, reflecting current scale of Dysca achieves the stable and reliable evaluation results.